A Complete Bibliography of Publications in Communications in Statistics: Simulation and

Computation: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.02

Title word cross-reference

2 [BG88]. 2 × 2 [Hab87]. 2 × k [KBB84, MP80]. 3 [BSL87]. 4 [Car89, Van87]. 2 [Wil82]. _{B} [BS88]. _{N} [Sch82b]. _{U} [BS88]. ARMA(P, Q) [Hie81]. B [SO87]. c(α) [Wan82a]. χ [Deu84]. D [CWK87, VDD89]. F [DF82, FS84, HW83, Joh86, KH83, Lee88, LS83, LRL^{+}83, Lin88, McA83, NW86, WCT86]. j [UP85]. k [KL85, Mag88, Ode89a, Ode89b, Sza85, Tam82]. k = 3 [KM88]. L [Wel89a, Gon86, KL85, Ode89b]. L_{1} [JS83]. L_{p} [MAGHB82, SHS83]. l_{u} [TJ82]. l_{V c} [Sin82]. m [Aro80, KL85, OTVA80, SS85a, TA80, VOA80]. μ [SS88a]. n [GC89b, Moo81]. p [Ali88, MAGHB82]. p × q [KV87, Ngu85]. q [Wil85a]. ρ_{jk} = α_{jk}α_{jk} [Nel82]. S_{B} [BSS81, BS89b]. S_{U} [BS89b, BS80]. σ [SS88a]. t [BLS86, BB80, Dag88, Deu84, FR80, GS87, HH83, KM87a, KKM86, LS88c, MB88b, Nel81, Nel82, RN80, Som83, Tse84, Wil85b, Wil86a]. v = 9 [KM88]. W [SB87, WCT86].

- [KG84a]. -ages [Moo81]. -Chart [CR88]. -Charts [CP89]. -critical [DF82]. -cumulative [Lee88]. -designs [HH83, KM87a]. -dimensions [Aro80, VOA80]. -distribution [Deu84, LRL^{+}83, Lin88, NW86]. -estimators [SS85a]. -expectation [SO87]. -factor [BG88]. -norm [Gon86]. -optimal [CWK87, VDD89]. -Parameter [Car89, BSL87]. -population [Sza85]. -ratio [LS83]. -ratios [Tse84]. -statistic [Wel89a]. -tail [Som83]. -test [Wil86a, SB87].

1 [EG87].

2SLS [Tse84]. 2x2 [OWJL88].

abscissa [Neu84]. Absolute [RSE89, Cau86, DP88, HN89, LS81, Ode82b, TLF88, WN81, Wes81]. abstract [NNS86]. accuracy [DK83]. adaptive [Hüs87]. additional [STS88]. Additions [Cho81]. adequacy [War82]. adjusted [AW88]. admissibility [Par88]. after [SH81]. against [KG88, Mag88, War84, YB87]. ages [Moo81]. aging [KG88]. agreement [CW85]. Akaike [Kit84]. Alam [BG85a]. algebraic [CBR80, KM87a]. Algorithm [KG89a, KG89b, BW84a, BF85, Bol86, Bol87, BO88, Eng82, Gon86, HH83, Joh86, KM87a, LH88, Lin81, MP80, Mon84a, Mon84b, SW88a, SH83, SA84, TO85, WN84, Wes81]. algorithmic [DM84]. Algorithms [DHL89, KG80, KG82, KG83a, KG83b, KG83c, KG83d, KG84a, KG84b, KG84c, KG84d, KG84e, KG85, Yan88, Gon86, Kem88, PF86a]. alias [Gre81]. Almon [TB85]. almost [Nom88b]. alternative [ELS87, LS84, Mag88, MYT87, Sch87, YB87]. alternatives [Ber82, War84]. among [IC82]. analogue [Eva87]. Analyses [HA89a]. Analysis [Bak89, EW86, LF89, SC89b, SS89b, BR86, BJ85, Bir83, BCL81, CL83, Gna83, HN89, HH86, Kit84, Lev87, NM88, Pai88, Pal83, RSV80, RBL88, SHHR83, SFM81, Sch87, Sho86, Smi84, Sta82, SJ88, Swa88, TF87]. analytic [Hap88]. Analyzing [Chu87]. Andrews [EHA86]. animal [RBL88]. Announcement [Ano86a]. annuity [Moo81]. ANOVA [Ano87d, EW86, Mac81, Ski83, TT85b, WCT86]. application [DP88, Koe81, Smi84]. Applications [SWP89, HF89, Kle81, KL85]. Approach [Sam89, BHP84, CW85, DM84, FM83, FG83, IC82, LB88, MD85, Qui80, SA84, Van87, Wan82b]. approaches [Ski83]. Approximate [AJG86, BM87, CJG86, IS81, LS88a, PN89, BM83, BM85, BSS81, BB86, IBM83]. Approximating [LRW89b, Old85, SC89c, TG82, Wil82]. Approximation [Pep89, Woo89, Bia81, Boi85, DK83, Fra89, GS87, MG84, MM83, MT89, Ump83, Van87, War82, vdLvP87]. Approximations [BB80, Tse84, Wal84]. arbitrary [SD86]. area [AS86, OA86, Som83]. argument [Klo88]. ARL [CP89]. ARMA [OTVA80, PD80b, Qui80, SMF89, dGS88]. assessing [TK88]. associated [BSL87, Zac80]. association [Gro86, OWJL88]. assumed [Tik87]. assuming [Moo81]. asymmetric [KKM86, Lev81]. asymmetry [FB87]. Asymptotic [Har84, Hen88, SS85a, FK86, Koe81, Mar87b, NK84, Nor85, SQ83b, Wel89a, Win80]. asymptotically [KG88, Mar87a]. Attributes [SK89]. Autocorrelated [Fir89, AR86, GP85, Ind87, WF82]. autocorrelation [AD88, JP86, Qui80]. autocorrelations [Puk84]. autocovariance [AD88, NT82]. Automatic [JP86]. autoregression [PT84]. Autoregressive [HA89b, Sha89a, TA80, BA84, LH88, Sim86]. auxiliary [HI88]. Average [HA89b, Sha89a, LH88, NT82, VOA80].

B [Smi82]. B-spline [Smi82]. backgrounds [SMF89]. balanced [BHP84, Fei82, HD88, LG86, SF85]. Based [Som89, TBA89, BF84, BF85, Bol87, BB86, Cau85, Eme81, FS84, KMK86, MM83, Sta82, TC88, WQ85, YF87]. basis [HH83, KM87a, MG82]. Basu [Wel89b]. Bayes [Ali88, MD85]. Bayesian [BZ87, FM83, Kit84, NI85, RSV80, RBL88, Sha89a, SC89b, Smi84, SG84, Tan80]. be [Oht87]. Bechhofer [BG85b, BG86, BG87, BG88, BG89, SG89]. behavior [Kle81, MG84]. behaviour [Hie81, Qui80]. Bernoulli [Cau85, Gia87, San87, Smi84]. best [AB86, KL85, MB88b, SD86, SS89a, ZH84]. Beta [Car89, Klo88, UP85]. better [CCS85, Kle86]. Between [Far89, Som89, ATAGD89, AP83, CW85, DF82, Nor85, OTVA80, Puk84, TF87]. Bias [Ind87, AA88, BF81, Mcl82, OH88, SD88]. bias-reducing [SD88]. biased [LB88]. BIB [KM88]. Bingham [Woo87]. Binomial [GC89b, Ali88, BS87, CCS85, LfH85, LS83, LK86, Wal84]. bioassay [FM83]. birthday [MTNY82]. Bivariate [Fuj89, LS89b, BBK84, Joh86, LK86, MLH81, McA83, MT89, Nel87a, Ode82a, RN80, Win80]. block [AHS84, All86, Ber82, BO88, FS84, Fei82, Gro87, HS86, IBM83, SF85]. blocks [YF87]. BLU [Bug89]. BLUE [BL88a]. Board [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano87a, Ano87b, Ano87c, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b]. Bootstrap [SC89a, SS88b, AA88, CC83, FB87, PT86, RD81, SC89c]. Bootstrapped [SK88]. Bootstrapping [Ale88, DP88, Old85]. bordering [KMK86]. Borges [Kle86]. bounded [Bia81, LfH85, MT89]. Bounds [Bia81, Sch82a, Som83, BS81a, KRS86, MB88b, NNS86, SS88b]. bridge [Kli80].

calculating [Joh86]. Calculation [BHW83, Sim84]. calculations [Cau85, EB89]. calibration [CyD88, THT82]. cancer [SS88b]. capture [RBL88]. Carlo [ATAGD89, BZ87, EG87, FS84, GP85, Gro87, Ham80, HW85, JRW82, Ket82, KG88, Koe81, Krä84, MMW85, OWJL88, OT85, Old85, SFM81, Sch82b, Sta82, SG84, Swa88, SS89b, TG82, WR87, WCT86, YF87]. Case [Hol89, Bar85, HD84, Sin82, Wil82, dWN88]. cases [SC89c]. categorical [Chu87]. Cauchy [Win83]. cell [Che88]. Censored [KM89, LS89a, GL88, HL86, HSV87, MMW85, Pal83, SAU83, Slu83, vdLvP87]. censoring [EG87, Eme81, Hom89, PT86]. Censorship [Bug89, Mag88]. centers [OA86]. central [BLS86, NW86, Nel81, RN80, TG82]. centrality [NW86]. Certain [Haw89, Ale88, AP83, MK82]. Chain [Hol80]. Change [LS89b, HGF86, SW88a, SW85, TO85]. change-points [SW88a]. changepoints [SW85]. characteristic [MK82, Qui80]. Characteristics [BG89, Rub85, SG89, WC81]. Characterizations [Fro88]. Characterizing [BB87]. Chart [CR88]. Charts [CP89]. Chebychev [SA84]. Chebyshev [HS80]. Chi [Woo89, BB80, Cau86, DF82, NK81, War82]. chi-square [BB80, Cau86, DF82, NK81]. Chi-squared [Woo89, War82]. choice [MAGHB82]. choosing [ATAGD89]. circle [Qui80]. Clarification [Que82]. Class [DHL89, Cha88, Par88]. classical [BDH89, Lon81a]. Classification [TBA89, SD80, SK88]. classifier [MYT87]. classify [BBK84]. clinical [EB89, Smi84]. Closed [Che88, PN89, RBL88, Lyo80]. clustering [CL83, FG83, RF83]. Coeditor [KG85]. Coeditors [KG83b, KG83c, KG83d, KG84b, KG84c, KG84d, KG84e, KG89b, KG85]. Coefficient [LS89b, AP83, Hen88, MM83, Neu82, NK84, Ode82a, TB86, Tik87, Ump83, War82, Win80]. Coefficients [CDY89, BS87, DD83, HHF83a, HHF83b, Joh86, KPR89, SS86, Tan86]. collinearity [MH87]. color [BS86]. Combination [Woo89, Wan88]. combined [LHH80, Que82]. combining [AJ83, Fei82, Lem87]. Comment [HA89a, Ump83, Wil82]. Comments [AA84, Kle86, KV87, MR89, BLS86]. Common [Ped89, KS81, Oht87]. communication [Ano82g]. Communications [KG84a, KG83a]. Comparative [Bug89, SHHR83, Tan80, ASD85, AS87, BDH89, TT85a]. Comparing [HD84, Mar87a, CD88, Gan81, KDG87, TT86]. Comparison [ABS87, Hom89, MST89, NM85, SD88, TM89, TK89, BI84, Cam88, CMC88, Cha84, Die87, Hab87, HS83, HHF83a, HHF83b, KW87, LS81, LHS81, LS84, MS84, PD80a, SCJ86, SW80, SL84, SC89c, Sta82, SJ88, SG84, ZH84, vdLvP87]. Comparisons [FP85, RD82, Ano87d, Ski83]. compartment [FY80]. compensation [HG85]. competing [KM87b]. Competitions [Ano86a]. Complete [HL86, AHS84, Ang88, YF87]. completely [BG81a]. complex [LHH80, Sin82]. component [FK86]. Components [Amm89, Bre89, Cha84, Fel87, KDG87, LB88, LG86]. Computation [BM83, BM84, BM85, Bow84, HW83, Huh86, HSV80, IBM83, MLH81, NW86, SHA80, CB85, KG84a, Kli80, Lee88, Wan82a, KG83a]. Computational [PF86a, AJ83]. computations [Lon81b]. Computer [Dag88, Gna83, LK86, WQ85]. Computing [CP89, EMR89, BF85, Eng82, LRL^{+}83, Mar86, McA83, SW88a, SH83, Sun88, Wel89a]. concentrated [McA87]. concentration [BF81]. concerning [BZ87]. concordant [KV87, Ngu85]. condition [Lon81b]. Conditional [CyD88, Lev87, Hab87, HH86]. conditioned [Gon86]. Confidence [LG86, MB88b, Som89, AJG86, BI84, CJG86, Eme81, KRS86, LKK81, LRL^{+}83, OL88, TC88, Thu85, Wan88, Wil86b]. connection [AP83]. consequences [NNS86]. Considerations [HA89a, LF89, BR86, OTVA80]. consistency [Dex86]. consistent [McC86]. constraints [Lee84, TB85]. Contain [Ode89a, Ode89b]. Containing [LS89a, KL85]. contaminated [TSF83]. contamination [AO87, BCL81]. contemporaneous [Bol87]. context [Mat82]. contingency [Gro86, Hab87, KW87, KV87, MP80, Ngu85, Van87]. contingency-table [Van87]. Continuous [TBA89, Bia81, HHF83a, SHHR83]. contrast [SH81]. contrasts [Wil85a, Wil86b]. Control [SWP89, Swa88, SS89b, CD88, CCS85, HD84, HD88, KP86, SCJ86]. controlling [Gil84, Käm88]. Controls [LRW89a]. convergence [Hen88, Jef80, Neu84, SW88b]. convex [BCC80]. convolution [Bax81, Joh86, MP81]. CORFIS [Dav80a]. Cornish [Deu84, GS87, MM89]. Cornish-Fisher [MM89]. correct [MD85]. correction [Fom87, Fra89]. Corrections [Ano81h]. correlated [GPS85, ID82, Ode82b, SW80, Wil85b, Wil86a]. Correlation [CDY89, Hol89, LS89b, MST89, AP83, Bec88, Fom87, Hen88, IC82, ID82, KPR89, MM83, Nel87a, NK84, Ode82a, Sca85, TB86, Tik87, TK88, Win80]. correlationa [DK83]. correlations [Cam81]. Cost [Old85]. countries [Ano86a]. covariance [AW88, Bir83, CWK87, Eng82, HH86, Hie81, KDG87, SJ88, Sza85, Wan82b]. covariance-adjusted [AW88]. covariances [BL88a, Bol87]. covariate [BS89a, Chu87]. covariates [LT88]. coverage [Kab80]. coverages [OYO87]. covered [OA86]. Cox [WR87]. Cramer [NNS86]. Cramer-rao [NNS86]. Criterion [Sam89, BS89b, Lis82, Sin80, Sin82, WN81]. Critical [Ode82a, PF86b, Wil86a, CF80, DF82, Har84, JY88, KH83, Wil85a, vdLvP87]. criticisms [Sil85]. cross [BLWY87, PG88]. cross-validated [PG88]. crossover [SQ83a]. cubic [Jar88, PG88]. Cumulants [MM89, Deu84, Wit89]. cumulative [CCW83, HP84, Kem88, Lee88, LS88c]. curtailment [EB89]. curvatures [BHW83]. curve [RD81, SFM81, SHA80, SA84]. Curves [Don89, Lis85, RM88]. cusum [AA84].

Data [BC89, Fuj89, KM89, LS89a, MST89, AJ83, AW88, BSHA85, BK85, CD87, Chu87, EHA86, Eng82, HL86, HSV87, HSV80, IS81, KGC86, Krä84, Lis85, MM88, MG82, MMW85, NM85, OWJL88, Pai88, Pal83, PG88, RM88, SFM81, Slu83, Sta82, Sza85, TF87, Wan82b]. Deal [Oht87]. decision [BZ87]. decomposition [Col87]. definition [Aro80, BG81b]. degree [SCJ86]. degrees [HW83]. denominator [HW83]. densities [MP81]. density [BSL87, ESP88, LKH82, OA86, RSG88, TFH86]. departure [Cau86, Tan86]. Dependent [Fuj89, Fom87, KM87b]. Derivatives [Moo81, Lin81]. deriving [Thu85]. descent [Wes81]. design [Ano87d, BSHA85, Bec88, Ber82, HI88, Kle81, KCV85, VDD89, Wil85a].

Designing [BG81a, Gia87]. Designs [KL89, TM89, AP86, AHS84, All86, Bec88, BO88, BI84, Cha88, CWK87, FS84, Fei82, Gro87, HH83, KM87a, KM88, McA87, SF85, VDD89]. detect [BG81a, EHA86]. Detecting [LS89b, Cam88, CF80, JY88, TSF83, Zac80]. detection [FP85, Sim84]. determinants [Hie81]. Determination [PG89, LHS81, Nor82, Zac80, dGS88]. Determining [DM89, Rhi89, DT80, Sch87]. developing [Ano86a]. deviate [CF80, JY88]. deviation [Bar83, LS81]. diagnostic [MH87]. Dichotomous [TBA89]. Difference [Som89, Bar83, KM87b, SS85b]. differences [Cam88, Con86, DS80, Puk84, Sch82a]. Different [DM89, BZ87, HD88]. digits [KR81, UP85]. dilution [SD88]. dimension [Van87]. dimensional [SMF89]. Dimensionality [Sam89]. dimensions [Aro80, OTVA80, Sun88, TA80, VOA80]. Direct [MK82]. directions [Thu85]. Discrete [Nel87a, BZ87, Gre81, HHF83a, SHHR83, SA84, Zac80]. discriminant [BBK84, BR86, BCL81, HHF83a, KRS86, MYT87, MG82, SHHR83, SM80]. disease [LHS81, Lev87, RF83]. disks [OA86]. dispersion [Gbu81, HSV80]. Distance [SWP89, MYT87]. distinguishing [TF87]. distributed [Mar87b, OA86]. Distribution [Bug89, Car89, De 89, DHL89, GC89b, Ode89a, Ode89b, SC89b, Woo89, Ada89, ABS87, BS81a, BS81b, BL88b, Bia81, Boi85, BS83, BS85, BS82, CM86, CCW83, CCW84, CMC88, CCS85, CG86, CW82, CS85, CS86, Dav80a, Deu84, DS80, ELS87, EMR89, FK86, Fra89, FP85, Gil84, GPS85, Hen88, HP84, HF89, HS86, Hüs87, IC82, Joh86, Kap87, KR81, KK82a, KSK87, kL80, LfH85, LRL^{+}83, Lin88, Mac81, MK83, MLH81, McA83, McC86, MD87, MM83, NW86, Nel81, Neu84, NI85, NK84, Nor85, Ode82a, Ode82b, ÖK88, PM89, Par81, SAU83, SB87, SO86, SO87, Sin80, Sin82, SC89c, SS89a, TG82, UP85, USM80, Wal84, WS88, Wel89a, WC81, Wil82, Win83, Wit89, WYMD86, YM88, YF87]. distribution-free [Hüs87, IC82, Mac81]. distributional [TJ85]. Distributions [Dan89, Fuj89, GS81, AHA84, BB87, BS89b, Dag88, Gre81, Gue82, HW83, HSV87, Lev81, MT85, MP81, Nel87a, Nel82, Neu82, Nor85, PF86b, Puk84, Swa88, TJ82, Woo87, Zac80]. disturbances [Ind87]. diversity [Jai86]. divisive [CL83]. Dixon [PF86b]. does [KM87b]. Dose [PN89, KK82b, SS88b]. Dose-Response [PN89, SS88b]. doses [SS88b]. double [SO86, SO87, SS85b]. doubly [MD87]. drawn [Rub85]. dual [SA84]. Duality [Far89]. due [BLS86]. Dunnett [HD88]. Durbin [Eva87]. dynamic [Bol88, Bol86, HG85, Ind87, Nel87b, Tse84].

Early [Far89]. Easily [Dag89]. easy [Mac81]. ED50 [Ham80]. Edgeworth [Fra89, MM89, SC89c, Tse84]. Editorial [Ano81a, Ano81f, Owe85, Qwe81, Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84a, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85a, Ano85b, Ano85c, Ano85d, Ano87a, Ano87b, Ano87c, Ano88a, Ano88b, Ano88c, Ano88d, Ano89a, Ano89b]. Effect [HA89b, KMH87, AW88, LHH80, WR87]. Effective [HA89a, LF89]. Effects [Bir80, BB83, Eme81, BHW83, BG81a, FS84, GWR88, Hol80, Mac81, Pal83, SF85]. Efficiency [CD87, AO87, Fom87, GL88, MT85, SR83, SQ83b, SS85a, SHS83]. efficient [Bol86, Bol87, Eng82, Mar87a]. elliptical [BB87]. emphasis [PD85]. empirical [Ali88, BG81a, BF84, BF85, Hap88, KH89, MD85]. endogenous [WF82]. endpoint [TB85]. epochs [Zac80]. equal [PM87, SM80]. Equality [CDY89, Fuj89, CJ85, DW81, DD83, Lee84, Mag88, Sat82, SS86]. equally [GPS85, Mar87a, Ode82b]. equation [Qui80, WF82]. equations [Xie89]. equicorrelated [Nel81]. Equioverlapping [KW89]. equiprobable [Wil82]. Error [SC89a, SH81, TBA89, ARG85, Arm85, Bia81, CL83, DRM83, KH89, KMK86, KMH87, Lis82, MT89, OM87, PD84, PD85, SK88, SG84, Swa88, TC89]. error-bounded [MT89]. error-free [KMK86]. Errors [Fir89, AR86, BCL81, GP85, HN89, Hap88, Ket82, Krä84, Mar87b, OP85, PM87, WF82, WN81]. errors-in-variables [Ket82, Krä84]. estimate [Ali88, LHH80]. estimated [ATAGD89, KRS86, KCV85, Noz84, Puk84, SHA80]. estimates [Bir80, BA84, Coo83, Fei82, HK80, JS83, MG84, Mar87b, OL88, Tse84]. Estimating [BS81b, MD87, Oma80, RD81, SC89a, Sca85, SQ83b, Win83, BG81b, CD87, HGF86, HHF83b, LS81, LHH80, LS84, McA87, MD85, Pep89, SD88]. Estimation [BS81a, Cam81, Car89, CC83, DRM83, DS80, DHL89, GC89a, GL88, GC89b, HA89b, KW89, Lee86, LS89a, Lis85, Liu84, Lju89, PM89, Win80, ARG85, ABS87, AA88, Bol87, BSL87, BHP84, BCC80, BK85, BI84, CMC88, EG87, Fel87, FY80, GG87, Gon86, GP85, GWR88, HS80, HL86, Hom89, HI88, Kap87, KS81, Kou81, Krä84, KSK87, LfH85, LS88a, LS88b, LS84, Mat82, MAGHB82, PT86, PD85, PD80a, RSV80, SAU83, SD86, SW85, STS88, SL84, Sza85, Tan80, TFH86, TO85, TC89, THT82]. Estimator [Haw89, Hut89, Rhi89, AO87, BS85, Fom87, HS80, Ind87, KK82a, KYR^{+}86, Maz83, Nom88b, Oht87, Old85, PM84, SS85b]. Estimators [Bre89, Bug89, Fir89, Hor89, KM89, KGC86, Pat89, Ped89, WF82, Bar85, BSHA85, BF81, BB83, BS83, BS82, BB86, CCW84, Cha84, CW82, CB85, Die87, ELS87, HC83, HSV87, HHF83a, HS86, Ket82, LB88, Lis82, McC86, Mcl82, Nom88a, OWJL88, Par88, PD84, RD81, RSG88, SW88a, SCJ86, SS88a, SD80, SW88b, SS85a, SK88, Spa87, SHS83, SG84]. Euclidean [MYT87]. Evaluation [BS80, OH88, BSHA85, Lin81, McA87, Nel81, Sta82]. even [HW83]. event [BG85a, BG85b, BG86]. evidence [ATAGD89, Krä84]. Exact [Cau85, Sin82, Som89, BM84, BM85, EMR89, Hab87, KV87, LH88, MP80, Ngu85, Nom88a, Sin80, YF87, vdLvP87]. exactly [Old85]. examination [Sil85]. exchangeable [Pep89]. executive [HG85]. expansion [Deu84, SC89c, Wel89a]. Expansions [MM89, NK84, Win80]. expectation [SO87]. expectation1 [Kle86]. Expected [Cau86, Sch82a]. experiment [SQ83a]. experimental [BSHA85, Bec88, HD84, HD88, Kle81, KCV85]. experiments [BG88, BGJ89, BG81a, Gia87, Hol80, Nel87b, RBL88]. Explicit [BSS81]. Exponential [LS89a, BLS86, Eme81, HL86, KSK87, kL80, MT85, NI85, SAU83, SO86, SO87, YB87]. exponential-based [Eme81]. exponentiality [KG88]. Exposure [Dan89]. expressions [Lyo80]. Extended [CF80, JY88, WS88, BG86, Wil85a]. Extension [Pat89, HD88]. Extensive [vdLvP87]. extrapolating [BS81a]. extrapolation [KK82b]. Extreme [Bug89, CF80, FP85, HS86, JY88, MK82, ÖK88, PF86b]. extreme-value [FP85, HS86].

F [MLH81, SH81, Woo89]. facilitate [TJ82]. factor [AJG86, BG88, BGJ89, CJG86, Mac81, PF86a]. Factorial [Bre89, Hol80]. factors [EMR89, HW84]. failures [San87]. families [Slu83]. Family [MST89, Pat89, LS83, Nor85, Woo87]. Faults [SH83]. Finding [RSE89]. Finite [Bol88, DM89, BM83, CyD88, IBM83, SR83, VDD89]. firms [HG85]. first [BA84, Fom87, Gon86, KR81, Lin81, NT82, UP85]. first-order [Gon86]. Fisher [KV87, BF81, Deu84, DK83, GS87, KBB84, MM89, Ngu85, Que82, Woo87]. Fisher-Bingham [Woo87]. Fisher-Cornish [GS87]. Fisher-expansion [Deu84]. Fit [Wel89b, HKL84, WYMD86, YM88]. fitted [BS89b]. Fitting [PG88, MM88, SA84, TJ82]. fixed [Hol80]. Flack [MR89]. fold [TC89]. Forecasting [HA89b, Bol86]. Forecasts [Sha89a]. Form [PN89]. forms [DC81]. formulas [Gue82, Har84]. formulation [Gre81]. Fortran [Sun88, TFH86]. four [OWJL88, SHHR83, WC81]. four-parameter [WC81]. Fourier [TFH86]. Fourth [DH85]. Fractional [Fed87, LB88, And88]. fractions [BS80]. free [BO88, Hüs87, IC82, KMK86, Mac81]. freedom [HW83]. full [Col87]. full-period [Col87]. Function [GC89a, Haw89, KM89, BS81a, BS81b, Boi85, BCHJ88, Cam81, CCW83, HHF83a, JP86, KRS86, Lin81, Liu84, MYT87, MG82, Moo81, NT82, OA86, Qui80, SM80]. functional [Krä84]. functionals [And88]. Functions [De 89, BBK84, Bow84, HP84]. Further [BK85, Dav80b, USM80, HG85, Que82]. Future [Ode89a, Ode89b, ASD85].

gains [Fom87]. Gamma [And88, DHL89, Ada89, BS83, BSL87, BS82, CW82, Kim88, LKH82, Lin81, Pep89, Sim86, SS89a, WS88, WC81, YB87]. Gauss [Bai88, BW84a]. Gaussian [AHA84, BM87, CCW83, CCW84, Dag89, ESP88, Haw89, Kro89, NT82, PM89]. GCVPACK [BLWY87]. General [OTVA80, BSHA85, BS89b, LB88]. Generalised [Dag89, Arm85]. generalization [KV87, Neu86, Ngu85, RN80]. Generalizations [CD88]. Generalized [Noz84, SC89b, SWP89, BL88a, BL88b, BLWY87, Bol88, BS85, CS85, Con86, CS86, CB85, Dav80b, HC83, KMK86, Nom88a, SHA80, WC81]. generate [KM87a]. Generating [GC89a, Jon85, Rub84, BO88, HH83, MK82]. Generation [VDD89, CWK87, Dag88, Dav80a, LK86]. Generator [Dag89]. geometric [ASD85, Coo83]. Gini [Bar83, Jai86]. given [Nel87a]. Givens [BW84b].

GLM [Yan88]. global [Win83]. Gompertz [Moo81]. good [Nor82]. Goodness [Wel89b, HKL84, WYMD86, YM88]. Goodness-of-Fit [Wel89b, WYMD86, YM88]. gradient [Gon86]. Gram [Lon81a]. graphical [KDG87, KP86, TF87]. Graybill [Oht87]. group [HD88, LHS81, OP85, PM87, SM80]. grouped [CD87, GL88, IS81, SS88a]. grouping [CL83, WR87]. groups [CW85, HD84, HD88]. growth [Lis85, SFM81, SHA80]. Grubbs [BB86]. guidelines [WQ85].

H. [BLS86]. half [MT85]. half-normal [MT85]. Halperin [vdLvP87]. Harrison [Bol86]. Hatanaka [Fom87]. having [De 89, MP81]. hazard [CD87, Jar88]. heterogeneity [ZH84]. heteroscedastic [MM88, OT85, SG84, TT85a]. heteroscedasticity [BB83, TT85b, Tan86, TT86]. heteroskedasticity [Ken85]. hierarchical [Pep89]. Higher [NK84]. homogeneity [BB86, KW87]. homogeneous [And88, DM84]. homoscedastic [DT80]. Hotelling [Dav80b]. Hyper [DHL89]. Hyper-Gamma [DHL89]. hypergeometric [Lyo80]. hypotheses [IS81, JM83, Slu83]. Hypothesis [BC89, Smi82, DRM83, DP88].

Identification [TJ85, JP86]. identify [Lee89]. ii [BG88, BG86, Sam89, dWN88]. III [BG89]. ILM0001 [CP89]. imperfect [TH85]. implementation [TFH86]. Implemented [Dag89]. Implementing [Haw89]. Improved [Kap87, Rhi89, Wal84, Wil86b, BG86, KP86, Lam88, Mon84a, Mon84b, Wil85a]. improvement [Koe81]. incomplete [All86, Fei82, HSV80, Klo88, Lin81, Lis85, SF85, SFM81]. Independence [PD80b, AS87, KH89]. Independent [KM87b, Fri88, KL85]. index [Gbu81]. inducing [IC82]. inequality [Lee84]. Inference [Ham80, CyD88]. Inferences [Ang88, Sha89a, Tan86]. Influence [Hut89, RD81]. Information [AHA84, Sam89, AJ83, All86, HI88]. Inner [Gil84]. input [IC82, ID82]. insect [SCJ86]. Inspection [SK89]. inspections [TH85]. integer [CS86, FG83]. Integral [Ano81g, RN80, Xie89]. integrals [Ano81h, Cho81, MP81, Owe80]. integrated [NT82, PD84]. intensity [Lee86]. inter [Fei82]. inter-block [Fei82]. interaction [BG88, LG86, Mar86]. interactions [BSH87]. interblock [All86]. internal [Dex86]. Interpenetrating [GG87]. interrelated [Smi84]. interrelationships [OTVA80]. intersection [Sca85]. Interval [Hor89, KS81, KS80, TC88]. Intervals [Ode89a, Ode89b, Som89, AJG86, ASD85, CJG86, CR88, Eme81, LKK81, LG86, LT88, LRL^{+}83, OYO87, OL88, SO86, SO87, Wan88, Wil86b]. intra [Fei82]. intra- [Fei82]. Intraclass [MST89, DK83, KPR89]. intrinsic [BHW83]. invariant [SS89a]. Inverse [Dag89, AHA84, CCW83, CCW84, Che88, ESP88, Jon85, PM89, THT82, Wal84]. inversion [KMK86]. invertibility [PD80c]. inverting [Wan82b]. investigation [EHA86, Ham80, Hap88]. involve [HW84]. involving [LRL^{+}83, Lin88, MP81, TO85]. IR [SMF89]. isotonic [OH88]. isotropic [FK86]. iterated [Bir80, Jef80, MG84]. iterative [Kou81, MM88]. IV [Swa88].

jackknife [RD85, RD81]. Jackknifing [PM84, GG87]. James [Mat82]. Jensen [Joh86, MLH81, McA83]. jn [IF86]. Johnson [BS88, JRW82, KKM86, Sha89b]. Joint [De 89, McA83]. judged [Mar87a]. judges [CW85].

K38 [Ano81h]. Kaplan [AO87]. Karber [Ham80]. Kiefer [BG85b, BG86, BG87, BG88, BG89, SG89]. known [CWK87, SD86]. Kolmogorov [ESP88, Har84, HKL84, Wil89]. Kruskal [Lev81]. kurtosis [MB88a, MK83, SHS83].

l [Klo88, Sch82b]. l-x [Klo88]. L1 [Hom89]. L2 [Hom89]. Labeling [RSE89]. Lachenbruch [MR89]. Lack [RF83]. Lad [Osb89]. lag [Puk84]. lagged [Fom87, WF82]. Laplace [Neu84, YM88]. Laplacian [Hut89]. Lar [RSE89]. Large [ASD85, TLF88, DK83, HG85, Ket82]. larger [CS85]. Largest [BG89, BG85a, BG85b, BG86, BG87, BG88, Che88, SG89]. latent [Ale88]. laws [Kou81, PD84, PD85]. Least [KS80, Lon81b, Ode89b, RSE89, Bir80, DP88, HS80, HGF86, HS83, Hel88, HK80, Ind87, IF86, Jef80, JM83, KMK86, Kle81, KCV85, Krä84, LS81, MG84, Noz84, PM84, San87, Wes81, KS80]. least-failures [San87]. Left [LS89a]. Lemma [Wel89b]. Level [LRW89b, Dav80a, Hol80]. life [HSV87, NI85]. lifetime [Eme81, GL88]. Likelihood [Car89, KW89, BF81, BS83, CCW84, CW82, ELS87, EG87, Fel87, HHF83a, Kit84, Koe81, Lev87, LH88, LS83, MM83, SS88a, Sin80, Tan80]. likelihood-ratio-based [MM83]. Limitations [BSH87]. Limiting [Qui80]. limits [EMR89, Gil84]. Line [RSE89]. Linear [BC89, De 89, Woo89, ARG85, Ali88, AA88, AD88, Arm85, AB86, BSHA85, Bol88, Bol86, BO88, Die87, Ind87, JM83, KMK86, KM87a, LS88a, LS84, MYT87, MAGHB82, Nor82, OT85, Raz87, SD86, STS88, SM80, SH83, SA84, TC89, Wan88, WN81, Wil85a, Wil86b]. Lines [Far89, MM88, Sca85, TT85a]. linkage [Tan80]. LM [OT85]. Locally [SS89a, McA87]. Location [Bug89, Don89, BCL81, Jef80, KMH87, MG84, ND81, RD81, Win83, WYMD86]. log [KK82a]. log-zero-Poisson [KK82a]. Logistic [BS89a, Sam89, BL88a, BL88b, BCHJ88, HHF83a, HHF83b, LS88b, ME86, WYMD86]. lognormal [Ang88, ELS87]. lognormality [BM87]. logrank [CG86, MMW85]. long [EB89]. long-term [EB89]. look [HG85]. Low [KK82b, PD80c]. lower [BS81a, Kim88]. LR [OT85].

M [Hor89]. M.D.I. [BCC80]. M28 [Qui80]. M29 [Lyo80]. M30 [HK80]. M31 [KS80]. M32 [KR81]. M34 [BG81b]. M42 [Sat82]. M498 [FM83]. MA [OTVA80]. main [Mac81]. make [KM87b]. Makeham [Moo81]. management [HG85]. Mandel [Mar86]. Marginal [Fuj89]. marginals [Nel87a]. Markov [Bai88, DM84, KS86]. matched [BM85, Lev87, MTNY82, MMW85]. matches [MTNY82]. Mathisen [CD88]. Matrices [Hol89, Ale88, Eng82, Fel87, Jon85, KDG87, MK82]. Matrix [Hut89, AHA84, CWK87, HSV80, KMK86, Lev81, Lon81b, Sza85, Wan82b]. max [AB86, KH83]. max-test [KH83]. maximization [Kab80]. Maximum [BS83, Car89, EG87, KW89, BF81, Boi85, CCW84, CW82, ELS87, GPS85, HHF83a, Ode82b, SS88a, Tan80, USM80]. may [Oht87]. McKay [Ump83]. Mean [BG89, Haw89, Ped89, Ang88, ASD85, Bar83, BG87, BG88, CS85, FK86, HGF86, HL86, HSV87, HSV80, KS81, KKM86, LKK81, Lis82, OL88, PM89, SG89, SR83, Sza85]. Means [Bak89, BL88a, Ode89a, Ode89b, CL83, DW81, Gan81, KGC86, Oht87, Pep89, SM80, TT86]. measure [Dex86, Jai86, Lev81]. Measurement [Arm85, BCL81, Pai88]. Measures [OM87, TM89, BF84, BF85, Fro88, Wil85a]. measuring [Ken85]. Media [Dan89]. median [Cau86, CD88, Eme81, Lem87, MS84, Sch82b, SR83, THT82]. medians [Mag88]. Meier [AO87]. Method [CP89, RSE89, BJ85, Bol87, CD87, Gre81, HGF86, KMK86, Moo81, Sho86, SC89c, Tan80, Thu85]. methodology [Sil85]. Methods [TM89, ABS87, CL83, CC83, Fel87, HHF83b, LHS81, MK82, Mar86, MS84, NM85, PT86, Pep89, SHHR83, SD88]. microcomputer [EW86]. middle [SAU83]. mild [BB83]. min [AB86]. min-max [AB86]. Mini [Ano82g]. Minicommunication [CB85]. Minicommunications [Ano81i, Ano83g, Ano84g, Ano84h, Cho81]. Minicomputers [Ano82h]. Minimal [GK89b, GK89a]. Minimax [STS88, LS88a, Nom88a]. minimum [HN89, WN81]. Misclassification [BBK84, CC83, RD85]. Mises [BF81]. Missing [BC89, BSHA85, Lju89, Sza85, AP86, BK85, Eng82, Wil86a]. Misspecification [HA89b]. misspecified [Raz87]. mixed [AJG86, CJG86, FG83, HHF83a]. mixed-integer [FG83]. mixture [AHA84]. Mixtures [MB88a, Neu82, SHHR83]. ML [Bug89, dWN88]. ML-II [dWN88]. MLE’s [PG89]. modal [Kem88]. Model [Bre89, KW89, Ped89, PN89, ARG85, AA88, AJG86, Arm85, Bai88, BSHA85, Bol86, Bol87, BS86, BHP84, CJG86, CyD88, EG87, Fom87, FY80, Hol80, HHF83a, Ind87, Jef80, JP86, JM83, KS86, Ket82, Kit84, Krä84, LG86, LT88, Mar87b, MAGHB82, OT85, OTVA80, PF86a, Pep89, PD80a, SD86, SHA80, SW88b, SA84, SS88b, Spa87, SMF89, Tan86, TJ85, TC89, Tse84, WF82, WR87]. Models [BC89, HA89b, Sam89, Sha89a, SS89b, ATAGD89, BHW83, Bol88, Cam81, Cha84, Chu87, GL88, GWR88, IS81, KK82b, LS81, Lee84, LH88, Lis85, Liu84, OTVA80, PD80b, PD80c, Raz87, RD82, SH83, SG84, Swa88, TA80, VOA80, dGS88]. models-time [VOA80]. modification [GS87]. Modified [CCW84, CW82, HKL84, Lon81a, MM89, PD85, WYMD86, YM88, ESP88, Har84, KKM86, PD84, San87]. modulus [USM80]. Moment [GC89a, BS85, CCW84, CW82, Lee89, Ode82a, TLF88]. Moments [Ada89, DC81, SR83, BS86, CS86, CB85, Koe81, Wal84, WS88]. monotone [Chu87]. monotone-scores [Chu87]. Monte [ATAGD89, BZ87, EG87, FS84, GP85, Gro87, Ham80, HW85, JRW82, Ket82, KG88, Koe81, Krä84, MMW85, OWJL88, OT85, Old85, SFM81, Sch82b, Sta82, SG84, Swa88, SS89b, TG82, WR87, WCT86, YF87]. mortality [Moo81]. most [Lin88]. Moving [HA89b, Sha89a, VOA80, LH88, NT82]. MRBP [IBM83]. MRPP [BM84, BM83, TK89]. Mulholland [BLS86]. Multi [SWP89, BW84a, BGJ89, BM83, BM84, Bol86, DD83, IBM83]. multi-factor [BGJ89]. multi-process [Bol86]. multi-response [BW84a, BM83, BM84, IBM83]. Multi-Sample [SWP89]. multicollinear [NM85]. Multidimensional [Lee84, And88, WQ85, Yan88]. Multinomial [De 89, Sam89, BG85a, BG85b, BG86, BGJ89, Che88, Coo83]. Multiple [CDY89, SW85, Sha89a, SK89, CF80, MM83, Ski83, WN81]. multiplicativity [BGJ89]. multiresponse [CWK87]. multistage [SS88b]. Multivariate [BC89, De 89, Kro89, Nel82, PT84, BB87, Bol86, DT80, HD84, HD88, HSV80, JY88, KP86, MK83, Nel81, NM88, Sch82b, SS85a, Sza85, TK88, Van87, dWN88]. multivariate-quality [KP86]. Multiway [Osb89]. myths [MB88a].

nearest [BJ85]. nearly [BO88]. Negative [Ped89, BW84b, Cha84, CS86, LS83, LK86]. neighbour [BJ85]. nested [TC89]. network [Käm88, MP80]. Newton [BW84a]. Neyman [SS83]. nine [Sun88]. no [BG88, MM89, PM87]. noise [Cam81, KRS86]. noisy [PG88]. nomparametric [vdLvP87]. Non [BW84b, Ped89, AR86, BB83, BLS86, CM86, Cha84, DM84, FK86, LHH80, Mar87b, MM83, NW86, Nel81, RN80, SS85b, TG82]. non-central [BLS86, NW86, Nel81, RN80, TG82]. non-centrality [NW86]. non-homogeneous [DM84]. non-isotropic [FK86]. Non-Negative [Ped89, BW84b, Cha84]. non-normal [AR86]. non-normality [BB83, LHH80]. non-normally [Mar87b]. non-null [CM86, MM83]. non-response [SS85b]. noncentral [Deu84, HW83, Lyo80]. noncentrality [Deu84]. nonhomogeneous [DRM83]. Nonlinear [KS86, LRW89a, SS89b, ATAGD89, BHW83, Gon86, HS83, JM83, MH87, Swa88]. nonmetric [RM88]. nonnormal [Gue82, GS81, Pai88, Swa88]. nonnormality [TT85b, TT86]. Nonparametric [GWR88, Ski83, TM89, Cam88, CW85, KH89, LS83, PT86, Par81, RSV80, SW88a, SFM81, Sch82b, Sho86, SJ88, TC88]. nonparmetric [SW85]. nonsquare [Bec88]. nonstationary [KS86]. Norm [JS83, Gon86, MAGHB82]. Normal [BG89, De 89, Gue82, LS89b, Ode89a, Ode89b, Pat89, AR86, Ano81h, BBK84, BG87, BG88, BHP84, Cho81, CJ85, Dag88, DD83, DT80, DC81, EMR89, Gil84, GPS85, KH89, KL85, MT85, MMW85, MD87, MM83, MT89, Nel82, Ode82a, Ode82b, OYO87, Oht87, Owe80, SS88a, SG89, Sun88, Sza85, TSF83, Van87, Win80, ZH84, dWN88, vdLvP87]. normal-theory [KH89]. Normal/Independent [Pat89]. Normality [SWP89, BDH89, BB83, LHH80, Tan86, TSF83, TK88, War84]. normally [Mar87b]. normals [Thu85]. Note [Lju89, DK83, Fei82, Fra89, HK80, KH83, Mar86, SM80, TB85, Thu85]. null [CM86, MK83, MM83, Wel89a]. Number [MTNY82, SW80, BZ87]. Numbers [BG89, CG86, Gre81, SG89]. Numerical [Bai88, Bre89, Gon86, Kli80, MP81, Nel81, Zac80, Bax81, Gro86, Hie81, Win83].

o [Van87]. Objects [DM89]. observation [Fom87, JY88]. Observations [LS89a, AP86, BK85, HW84, OP85, PM87, SW80, Wil86a]. observed [BS89a]. obtained [Bir80]. obtaining [KRS86]. odds [OWJL88]. once [NT82]. One [SO86, Wan88, AA84, FY80, GWR88, HI88, KK82a, LT88, Ode89a, Ski83, SC89c, TC89]. one-fold [TC89]. One-sided [SO86, Wan88, AA84, Ode89a]. one-way [GWR88, LT88, Ski83]. open [FY80]. optimal [Bec88, CWK87, KG88, SHS83, VDD89, Wan82a]. optimization [Win83]. Optimizing [De 89, WQ85]. Order [BL88b, BL88a, BBK84, BA84, DH85, Eva87, Gon86, NT82, NK84, PD80c, Sch82a, WS88, dGS88]. ordered [BG81a, Ber82]. Ordering [LRW89b, Fri88]. ordinal [Chu87]. Ordinary [Krä84, Ind87, KCV85, Noz84]. origin [MP81]. orthant [Sun88, Van87]. orthogonal [AV89, Cha88]. other [WC81]. outlier [FP85, Kim88, Sim84]. outliers [CF80, EHA86, Hom89, KMH87, Kit84, OM87, PF86b, SS89a]. Overview [SS89b].

P [BLS86]. Page [Wel89a]. paired [KGC86, Sta82]. paired-comparison [Sta82]. pairs [BM85, Kim88]. pairwise [Ano87d]. Papadakis [BJ85]. Paper [MR89, AA84]. parallelism [TT85a]. Parameter [Car89, DHL89, GC89b, HA89b, KW89, LS89a, PG89, Ali88, BHW83, BS85, BSL87, CCW84, CMC88, CW82, Deu84, ELS87, HL86, KK82b, LfH85, Lin81, Mar87b, NW86, NI85, Nor82, SO86, SD88, TC89, WC81]. parameter-effects [BHW83]. parameterization [ELS87]. Parameters [Bug89, ABS87, AB86, BB87, BM83, BF81, BG81b, BS80, BA84, CW82, ESP88, FY80, Hom89, HS86, IBM83, JM83, KK82a, Kou81, LS81, McC86, MD87, NNS86, PD80a, SAU83, SS88a, SHA80, SL84, Wil86b, WYMD86]. Parametric [TM89, Cam88]. Pareto [CMC88]. Part [Gon86, Sam89]. partial [AP83, Hel88, Lin81, Puk84]. partially [BS89a]. PASCAL [Pan88]. patterned [Sza85]. Pearson [Dav80a, Fra89, Que82]. Independent [Pat89]. people [MTNY82]. percent [BF84]. Percentage [Wil85b, Wil89, BB80, Dav80a, Huh86, Ode82b, Sin82]. Percentiles [WC81, BS89b, CG86, CJ85, GS87, HP84, OL88]. Performance [BG89, KYR^{+}86, Old85, San87, SHHR83, SG89]. period [Col87, EHA86, SQ83a]. periodic [LH88]. permutation [BM83, BM84, BM85, IBM83, KMH87, KBB84]. perspective [Nel87b]. Perturbation [HN89]. PH [Neu84]. PH-distribution [Neu84]. pharmaceutical [Chu87]. pharmokinetic [FY80]. pitfalls [TF87]. Plans [SK89, NI85]. plot [Bec88]. plots [EHA86, ID82]. plotting [HW85, Hom89]. point [OL88, Sca85]. Points [Far89, BF84, BB80, Dav80a, Huh86, Ode82b, Rub84, SW88a, SW85, Sin82, Wil85b, Wil89]. Poisson [And88, Bar85, BS85, CS85, Con86, CS86, Gbu81, KW89, Kem88, KK82a, Lee86, Par88, SC89b, TO85]. pollutant [McA87]. polygamma [Bow84]. polynomial [MT89]. polynomials [CBR80]. polytomous [Lev87]. polytone [RM88]. polytope [Rub84]. pooling [Hol80]. Population [BG89, DM89, BG87, BG88, BM83, Bol88, CyD88, DW81, IBM83, KKM86, Oht87, SS88a, SG89, Sza85, Tik87, ZH84]. populations [BBK84, DD83, DT80, Gia87, GS81, KL85, LS83, MB88b, Par88, RBL88, Sch82b, SR83]. portable [Gre81]. positions [HW85]. Positive [Gro86, DC81, KG88, Wal84, Wan88]. Positively [Fuj89]. possible [Hom89]. Posterior [SD80]. Power [Don89, HH86, TSF83, Bir83, EB89, KH89, KMH87, KCV85, KBB84, MMW85, TLF88, War84, WQ85]. powerful [Lin88, SF85]. powers [BS87]. Practical [Haw89]. precision [OYO87]. Predicting [GK89b, GK89a]. Prediction [AR86, Hor89, Ode89a, Ode89b, SC89a, TH85, ASD85, Bol88, Mat82, NM85]. predictive [Kit84, ME86]. preliminary [Gan81, KK82a]. presence [DRM83, TH85]. preserving [Bar85]. Principal [Amm89, FK86, KDG87, LB88]. prior [AJ83, CCS85, PM87, dWN88]. priority [Que82]. Probabilities [LRW89b, BBK84, CC83, Kli80, Lee88, LS88c, Lyo80, MLH81, McA83, ME86, MT89, PM87, RD85, Sun88, Van87]. probability [BG85a, BG85b, BG86, BM83, BM84, BM85, BF84, BF85, CCW83, Che88, Fri88, IBM83, Kab80, Kem88, MD85, NW86, Neu82, OA86, RN80, SD80]. problem [Ano86b, BF84, BCL81, CW85, FG83, JS83, MTNY82, ND81, Raz87, SW85, Sza85, Wes81]. problem-nonparmetric [SW85]. Problems [Haw89, Aro80, BSL87, Gon86, GP85, Kab80, THT82]. Procedure [BG89, Sha89b, TBA89, AJ83, BG85a, BG85b, BG86, BG87, BG88, BGJ89, Che88, DT80, Ham80, JM83, KDG87, KL85, Kou81, KP86, Lam88, MM88, Nor82, RD85, San87, Sch87, SG89, TT86, TF87, Tik87, TK88, dWN88]. Procedures [MST89, AA84, BM83, BM84, FP85, IBM83, MMW85, PD85, PD80a, SW85, SFM81, TT85b, Zac80, ZH84]. Process [GK89b, SWP89, And88, BM87, Bol86, BA84, GK89a, Lee86, Lon81a, NT82, Qui80, Sim86]. processes [AD88, DM84, Smi84]. product [Ode82a, Wil85b]. product-moment [Ode82a]. program [Pan88]. programming [BCC80, Coo83, FG83, KS86, SA84]. Properties [BS82, AHS84, AW88, ELS87, HS83, HC83, Jef80, Ket82, LKH82, Mar87b, Nom88a, PD84, SS88a]. proportion [Mcl82, Oma80]. Proportions [Som89, Coo83, Gil84, Sat82, SW80]. Proschan [Kle86]. prospects [Aro80].

Quadratic [PN89, DC81]. quality [KP86]. quantile [Kap87, PT86]. quantiles [WN84]. quasi [Ada89, LKH82]. quasi-range [Ada89, LKH82]. queue [Cam88]. quick [Mac81].

R [HF89]. R-distribution [HF89]. Ramey [BG85a]. Random [De 89, DM89, Hol89, TBA89, Ale88, BS81b, BZ87, Bia81, BG81b, CBR80, Dag88, Fri88, Gre81, GWR88, GPS85, HP84, LT88, LK86, MK82, Ode82b, Pan88, Rub84, Rub85, SS85b, UP85, Zac80]. randomization [SQ83a]. randomized [BG81a, Ber82, FS84, IBM83]. Randomly [BC89, KM89, HSV87, OA86]. randomness [HW84]. Range [Bar85, Rhi89, Ada89, LKH82]. Range-preserving [Bar85]. ranges [GS81]. Rank [AHS84, ID82, TK89, BSH87, Gro87, Hen88, IC82, Lem87, LS83, ND81, SF85, Sch82b]. rankings [SQ83b, YF87]. Ranks [Bak89, CM86, FS84]. rao [NNS86]. rare [LHS81]. Rate [SC89a, CD87, KH89, SH81, SK88]. Rates [SW88b, TBA89, CL83, Jar88]. ratio [BI84, DC81, GS81, KRS86, KGC86, Koe81, LHH80, LS83, Maz83, MM83, OWJL88, Sin80]. rational [BS80, Liu84]. ratios [Thu85, Tse84]. real [NNS86]. realized [KRS86]. recapture [RBL88]. reciprocal [PM89]. Recognition [AAA^{+}89]. Recovery [All86]. recursions [BS86]. recursive [Bol87]. reducing [KM87a, SD88]. Reduction [LRW89a, Sam89, Nel87b]. Referee [AAA^{+}89]. region [BHP84]. regions [BI84, PD80c]. Regression [Fir89, RSE89, Sam89, AA88, AR86, AV89, AB86, BHW83, BW84b, BB83, BS89a, Bol87, CD87, DRM83, DP88, Die87, EG87, Fom87, GG87, GP85, GL88, HN89, Hel88, HHF83a, HHF83b, Ind87, IS81, IF86, JS83, LS88a, LS81, LB88, LS88b, LS84, MH87, Mar87b, Mat82, ME86, MAGHB82, Nom88b, Nom88a, Nor82, OM87, OH88, SL84, Sil85, Smi82, Spa87, Tan86, TC89, WR87, WN81, WN84, Wes81, Wil86b, YB87]. regressions [TO85]. Reinforcement [KL89]. Rejoinder [Ano89c]. related [ARG85, BK85, UP85]. relations [KV87, Ngu85]. relationship [DF82, Nor85]. Relative [AO87, LKK81, KYR^{+}86, SQ83b, SS85a]. relevance [Ket82]. Reliability [KSK87]. Remarks [kL80, Bax81, Que82]. renewal [Xie89]. renewal-type [Xie89]. Repair [GK89b, GK89a]. reparameterizing [SH83]. reparametrization [SD86]. Repeated [NK81, Pai88, TM89, Wil85a]. Replacement [DM89]. replicate [Cha88]. replication [Fed87]. report [FM83]. required [SW80]. requirements [OYO87]. residual [Gon86, Nor85]. residuals [AP83, PD80b]. respect [Lin81, Tan86]. Response [AP86, KL89, PN89, ARG85, BW84a, BM83, BM84, IBM83, KK82b, RBL88, SS85b, SS88b, Sta82]. response-related [ARG85]. responses [GL88]. restricted [Bai88, Coo83, Raz87]. restriction [Kha87]. restrictions [KS80, STS88]. result [Neu86]. Resulting [BG89, SG89]. results [AD88, BK85, THT82, WCT86]. retrospective [Lev87]. Revisited [RSE89, AS86, BS83, KKM86]. reweighted [Jef80]. Rho [Fra89, TC88]. Ridge [Fir89, GP85, LS88b, CB85, HC83, Nom88b, Nom88a, Nor82, Sil85, Tam82]. Right [LS89a, Mag88, Slu83]. right-censored [Slu83]. risk [Ali88, BCHJ88]. risks [KM87b]. robusl [Wan82a]. Robust [Amm89, HL86, NM88, Pat89, TBA89, AP86, AV89, Bir80, Bir83, Die87, KMH87, MG84, RD81, SF85, SL84, TT85b, TT86, TB86, Tik87]. robustly [ATAGD89]. Robustness [KM89, BCHJ88, Eme81, LS83, RF83, Tan86, WCT86]. Rodrigues [Kle86]. root [Boi85]. roots [Ale88, CBR80, MK82, Qui80]. rotatability [DH85]. rotations [BW84b]. routines [BLWY87]. row [Ano87d]. Rule [SC89a, BZ87, Bir80]. rules [Hom89]. run [Rub85]. runs [CR88, WQ85].

S [BS88]. safe [SS88b]. same [Puk84]. Sample [Bre89, DM89, Rhi89, SWP89, AW88, ASD85, BM87, Bar83, BF84, BF85, CS85, DK83, DD83, ELS87, Fom87, FR80, Gan81, Gue82, HS83, Hie81, HD88, HSV87, HI88, Hüs87, JP86, Ket82, Kle81, KYR^{+}86, kL80, LHS81, LS84, MG84, McA87, MS84, NT82, ND81, NK84, Nom88a, Ode82a, OYO87, Par81, San87, SS88a, Sch87, SHS83, SC89c, TLF88, WQ85, vdLvP87]. Sampled [AD88]. Samples [Bug89, KW89, TK89, Ang88, GS81, HL86, Hom89, Kim88, LKH82, Pai88, Rub85, SAU83, TSF83, vdLvP87]. Sampling [Hap88, Lev81, MT85, SK89, Ano86b, BLS86, Che88, CyD88, CR88, kL80, NI85, Oma80, San87, SS85b, SG84]. Scale [Bug89, Hom89, TSF83, WYMD86]. scaling [Lee84]. Scheffé [Thu85]. Schmidt [Lon81a]. scores [Chu87, MMW85]. screening [OP85, PM87]. second [BBK84]. second-order [BBK84]. Section [KG89b, KG80, KG82, KG83b, KG83c, KG83d, KG84b, KG84c, KG84d, KG84e, KG85, KG89a]. seemingly [Bol87, Spa87]. select [dWN88]. Selected [BF84, Dav80a, KH89]. Selecting [BG89, CCS85, Spa87, BG85a, BG85b, BG86, BG87, BG88, Che88, KL85, SG89, SHS83, ZH84]. Selection [Sha89b, BGJ89, Gia87, Hol80, Kha87, Lam88, MD85, MB88b, San87, WN81, ZH84]. self [MM88]. self-weighting [MM88]. semi [BM87, DM84]. semi-Markov [DM84]. semi-variogram [BM87]. semiparametric [SW88b, Yan88]. separate [Slu83]. Sequence [LS89b]. sequences [Col87]. Sequential [BG89, BG85a, BG85b, BG86, BG87, BG88, Cau85, CWK87, LKK81, SG89, Zac80]. serial [Fom87]. Series [BLS86, Kro89, Lju89, Aro80, BK85, Deu84, EHA86, Klo88, OTVA80, RM88, SD88, TA80, TFH86, VOA80]. set [Chu87]. Sets [LS89a, KH89, Thu85]. Several [KM89, Ber82, CD88, DW81, HD84, KDG87, OWJL88, Qui80, TT85a, TT86]. shape [And88, Hom89]. shift [Zac80]. Short [Rub85]. Short-run [Rub85]. shortest [LRL^{+}83]. Shrinkage [RSV80, Lis82]. Sided [Ode89b, AA84, EMR89, Ode89a, SO86, Wan88]. Sieve [Haw89]. Sigma [Rhi89]. sign [Lem87]. signal [KRS86]. signed [Lem87]. signed-rank [Lem87]. Significance [MST89, Que82]. significant [SH81]. Simple [CP89, Kem88, McC86, RSE89, Die87, HGF86, HS86, JM83, JS83, Kle86, Mag88, SA84, TC89]. Simplified [SAU83]. Simulated [Don89, BSHA85]. Simulating [Dan89, Pan88]. Simulation [BSL87, Fir89, HA89a, KW89, KG83a, Kro89, LF89, RSG88, Sim86, TM89, AA88, BJ85, EB89, Fed87, IF86, Käm88, KG84a, KL85, LS88a, Nel87b, RSV80, RD82, STS88, SS86, Sil85, SMF89, Tam82, THT82, WQ85, Woo87, YF87]. Simulations [Sam89, PT86, SD80]. Simultaneous [LT88, Ode89a, Ode89b, WF82, Wil86b]. single [BGJ89, Cha88, JY88, KK82b]. single-stage [BGJ89]. Singularities [MM89, MP81]. situations [Hom89]. Size [DM89, Gue82, KM87a, KL85, LHS81, OYO87, Sch87]. Sizes [Rhi89, Hie81, HD88, KM88]. slightly [Wil85a]. slope [LS84, TC88, TC89]. Small [AW88, Bre89, Bug89, ELS87, Fom87, HSV87, Kle81, KYR^{+}86, MG84, Par81, San87, SS88a, CG86, Gon86, HS83, HL86, Hom89, LKH82, Nom88a, Pai88]. Small-sample [AW88, Kle81, San87]. Smirnov [ESP88, HKL84, Wil89]. Smooth [PT86, SC89a, KYR^{+}86]. smoothed [SK88]. Smoothing [Hut89, Jar88, RM88, PG88]. Sobel [BG85b, BG86, BG87, BG88, BG89, SG89]. Solution [PN89, Xie89]. Solutions [BS88, BSS81, KS86, KS80]. solving [Gon86]. Some [Bax81, Far89, Haw89, HC83, Jef80, PD84, Sam89, TT85a, TT85b, TK89, AHS84, Cha84, Don89, Gue82, Hab87, Kab80, KH89, KG88, Krä84, Mcl82, RD82, SW88b, SG84, War84, ATAGD89]. space [PD80b, RF83, VDD89]. space-time [PD80b]. Sparse [Fel87, KW87, OWJL88]. spatial [Bec88]. Spearman [Fra89, Ham80, Hen88, TC88]. special [PD85]. specification [HH86, dGS88]. spherical [Woo87]. sphericity [PD80b]. spline [KYR^{+}86, Smi82]. Splines [Hut89, Jar88, PG88]. spurious [JY88]. square [Bec88, BB80, Cau86, DF82, Lis82, NK81]. squared [PD84, PD85, War82, Woo89]. squares [Bir80, HS80, HGF86, HS83, Hel88, HK80, Ind87, IF86, Jef80, JM83, KMK86, Kle81, KCV85, KS80, Krä84, LS81, Lon81b, Mar86, MG84, Noz84, PM84]. Stability [CBR80]. stable [HP84, Kou81, McC86, Pan88, PD84, PD85]. stage [Ano86b, BGJ89, HS83, Lam88, OP85, PM84]. Standard [AV89, TK89, Bar83]. standardized [CCW83, HP84]. star [PD80a]. STARMA [PD80c]. starship [Owe88, OL88]. starting [Bir80]. states [Lev87]. Stationarity [PD80c]. stationary [Sim86]. Statistic [Som89, BSH87, CG86, FK86, Har84, Koe81, Sch82b, SHA80, SS83, Wel89a, WQ85, Wil89, Wit89]. Statistical [FY80, Gna83, LKH82, SS89b, Dex86, Swa88]. statisticians [Ano86a]. Statistics [KG83a, BL88a, BL88b, CF80, Hap88, KG84a, PF86b, Sch82a, Sim84, WS88, WCT86]. Stein [Mat82]. Step [PM87]. Step-wise [PM87]. Stevens [Bol86]. Stieltjes [Neu84]. Stochastic [EB89, Hut89, MG84, RD82, Tam82]. stopping [Bir80, Zac80]. straight [MM88, TT85a]. strategies [LS84]. strategy [dGS88]. stratified [Pal83]. stratum [HI88, Lev87]. stratum-matched [Lev87]. stress [Lev81]. String [Col87]. structural [TO85]. structure [Hel88, TC89]. structures [Bec88]. Student [BLS86, BB80, GS87]. Studentised [Wit89]. Studentized [Boi85, CF80, GPS85, USM80]. Studentizing [MS84]. Studies [Pat89, TC89, Fed87, IF86, JRW82, Lev87, TT85a, Tan80]. Study [Bre89, Bug89, Fir89, HA89a, KW89, LF89, TM89, AA88, ASD85, AS87, BDH89, BG81a, Ber82, BZ87, BJ85, EG87, FS84, GP85, Gro87, Gro86, HSV87, KH89, Ket82, KG88, LS88a, MMW85, OWJL88, OT85, RSV80, RSG88, STS88, SFM81, Sch82b, SS86, Tam82, TG82, WR87]. Studying [Ali88]. subpopulation [Oma80]. subroutine [Dav80a, Sun88]. subsampling [GG87]. subset [AB86, KL85, PT84, San87]. success [ME86]. suggestion [Ken85]. sum [DS80, HN89, LS83, Sch82b, WN81]. Sums [BS87, Mar86]. superiority [Raz87]. Supplementary [BG89, SG89]. support [KM87a, KM88]. supports [Bia81]. Surface [KL89, AP86, Sta82]. Surrogate [Coo83]. surveys [LHH80, LS84]. Survival [Fuj89, CD87, HSV87, KYR^{+}86, KGC86, RSV80]. Survivorship [KM89, AW88]. symmetry [FB87]. System [Sha89b, Gna83, JRW82, McA83].

table [Ano81h, Cho81, Gro86, ME86, MP80, Owe80, Van87, Wil85a]. Tables [HD88, HP84, KRS86, LRL^{+}83, Ode82b, Osb89, TJ82, BG86, BHP84, CCW83, Hab87, KW87, KBB84, KV87, Ngu85, OWJL88, USM80, WS88, vdLvP87]. tabulation [Dav80b]. Tail [AS86, Dag88, Gil84, KMH87, Som83]. taken [DM89]. Tausworthe [Col87]. Technical [FM83]. technique [Koe81]. Techniques [SC89a, CMC88, SL84, SJ88]. term [EB89]. Test [Fuj89, SWP89, AHS84, BM85, BF84, BF85, BB86, CD88, CF80, DD83, ESP88, Eva87, FK86, FS84, Gan81, Har84, HD88, Hüs87, JY88, KH83, KK82a, Kle86, KKM86, KBB84, KV87, Lee89, Lem87, LS83, Lis82, Mac81, Mag88, Mon84a, Mon84b, Ngu85, NI85, ÖK88, Par81, PF86b, SF85, Sat82, SS86, SB87, SQ83a, SH81, SS89a, TB86, Wan82a, Wil86a, YM88, vdLvP87]. Testing [BC89, CJ85, DW81, KPR89, Kim88, KKM86, MST89, Slu83, YB87, DRM83, DP88, FB87, JM83, NK81, PF86b, Sin80, Ski83, SH81, Smi82, SS83, Sza85, TT85a, TB86, Tik87, Wil85a]. Tests [CDY89, CP89, Don89, HW84, TK89, Wel89b, YF87, AHS84, AW88, AS87, BDH89, Ber82, Bir83, BSH87, Cam88, Cau85, FS84, Gro87, Hab87, HKL84, HH86, IS81, KH89, KMH87, KG88, KW87, Lem87, LS83, Lin88, MT85, Mar87a, OWJL88, OT85, PD80b, Que82, RD82, Sch82b, TSF83, TLF88, War84, WYMD86]. Theoretic [Sam89]. theory [Gue82, KH89, SG84]. there [Wil86a]. Three [OP85, PG89, AJG86, CJG86, CCW84, CMC88, CW82, ELS87, FS84, HS83, HHF83b, OWJL88, SQ83a]. three-factor [AJG86, CJG86]. Three-Parameter [PG89, CCW84, CW82, ELS87]. three-period [SQ83a]. Three-stage [OP85, HS83]. three-treatment [SQ83a]. ties [AHS84]. Time [Aro80, Kro89, Lju89, TA80, AD88, BK85, Cam88, EHA86, Nor85, OTVA80, PD80b, RM88, RF83, VOA80]. time-series [RM88]. time-space [RF83]. Times [GK89b, GK89a, HSV87]. Toeplitz [Hol89]. tolerance [EMR89, Gil84, LT88, OYO87, SO86, SO87]. Trace [Hut89]. trades [HH83, KM87a]. transfer [Cam81, Liu84]. transform [BSH87, ND81, Neu84]. Transformation [BR86, SD86]. Transformations [LRW89a, Sha89b, DT80]. transformed [BS81b]. translation [JRW82]. trap [RBL88]. treatment [AW88, Bai88, BG81a, FS84, MP81, MP80, SF85, SQ83a]. treatments [CD88, CCS85]. tree [Mag88]. trend [BO88, Par81]. trend-free [BO88]. trended [Krä84]. Trenkler [Neu86]. trial [LHS81]. trials [Cau85, EB89, Smi84]. trimmed [Ham80, SR83]. triples [Mon84a, Mon84b]. truncated [MD87]. Truncation [BG85b, BG86, BG87, BG88, BG89, SG89]. Tukey [Mar86]. tumors [CG86]. twelfth [Eva87]. Two [Bre89, CDY89, Ode89b, Ped89, Som89, TK89, ATAGD89, AHA84, Ano86b, Ano87d, BF84, BF85, BS86, Cha84, Con86, CW85, CJ85, DC81, EMR89, EW86, FR80, Gan81, Gro86, GS81, HS83, Hol80, HW84, Hüs87, KH89, Lam88, LG86, LHS81, Lin81, Mac81, Maz83, ND81, NI85, Oht87, PM84, Sat82, Sca85, SW80, SFM81, Sch82b, SO86, SC89c, SMF89, SG84, TF87, Wan88, WQ85, Wil85b, vdLvP87]. two-color [BS86]. two-factor [Mac81]. two-level [Hol80]. two-parameter [SO86]. two-sample [BF84, BF85, FR80, Hüs87, ND81, WQ85, vdLvP87]. Two-Sided [Ode89b, EMR89]. two-stage [Ano86b, HS83, Lam88]. Two-Way [Bre89, Ano87d, EW86, Gro86]. Type [Dav80a, NNS86, PF86b, SS89a, TK88, Xie89, BL88a, BL88b, Bug89, CL83, EG87, KH89, KMH87]. Type-1 [EG87]. Type-I [KMH87].

unbalanced [EW86, Wan82b]. unbiased [Bar85, Lin88, Nom88b, SD86]. unbiasedness [BG81b]. unclassified [MG82]. unconditional [Hab87]. unconstrained [BCC80]. underlying [LHS81]. Unequal [TK89, DW81, KCV85, Noz84]. uniform [Fri88]. uniformity [Lee89, SS83]. uniformly [Lin88]. Unimodal [LRW89b]. unit [HI88, Qui80]. univariate [TFH86]. Unknown [MM89, ESP88, WYMD86]. unrelated [Bol87, Spa87]. Updating [MG82]. upper [BS81a, CG86, Kim88]. urn [BS86]. usage [Gna83]. Use [Gan81, IF86, ID82, Oht87, SC89a, Wel89a]. used [BBK84, Kle86]. useful [Ber82]. Using [Bak89, FB87, HS80, LRW89a, Bol86, Chu87, Ham80, HSV87, HI88, JP86, JM83, Koe81, KV87, Ngu85, Pep89, PG88, SA84, Slu83, SHS83, WN81, Win80].

validated [PG88]. validation [BLWY87, RD82]. Value [Bug89, Bir80, DP88, FP85, HS86, Ode82b, ÖK88, PF86b, Tik87, Wes81]. Values [Lju89, BM83, BM84, BM85, BZ87, CF80, DF82, Eng82, Har84, IBM83, JY88, KH83, Ode82a, PF86b, Wil85a, Wil86a, vdLvP87]. variability [Fro88, NM88]. Variable [WN81, BZ87, CR88, Fom87]. variable-serial [Fom87]. Variables [De 89, LS89b, TBA89, Woo89, BG81b, DC81, Fri88, GPS85, HHF83a, HP84, IC82, ID82, KH89, Ket82, Krä84, LK86, Ode82b, SHHR83, Spa87, UP85, WF82, Zac80]. Variance [Bre89, HI88, LRW89a, AA88, BB86, CL83, Cha84, CS85, CWK87, DK83, Fel87, kL80, LG86, LHH80, LS84, Mcl82, NT82, Nel87b, SD86, Sch87, Sho86, SC89c, ZH84]. variance-covariance [CWK87]. Variances [Bre89, BL88a, DRM83, DW81, KCV85, Maz83, Noz84, RD81, TT85a, Wan88]. Variates [SS89b, And88, Con86, Dag88, Käm88, Pan88, Swa88, Wil85b]. variation [DD83, Neu82, SS86, Ump83, War82]. variogram [BM87]. various [Hie81, KM88]. VARSTAB [DT80]. Vector [HA89b, BS81b, HSV80, JP86]. vectors [Bia81]. version [Wan82a]. versus [FB87, Old85]. VI [Dav80a]. via [BCC80, EB89, Kit84]. vs [Lon81a].

waiting [Cam88, Nor85]. Wald [OT85, SHA80]. walks [Rub85]. Watson [Eva87]. Way [Bre89, Ano87d, EW86, GWR88, Gro86, LG86, LT88, Ski83]. Weibull [ABS87, EG87, FP85, Hom89, LS81, PG89, PF86b, SB87, SL84, Sim86, War84]. weight [KMH87]. weighted [Bir80, IF86, JM83, Kle81, KCV85, MG84, PD85, SQ83b, WN81, Wil89, YF87]. weighting [MM88]. weights [Fei82]. well [Gon86]. well-conditioned [Gon86]. where [LHS81]. which [BG85a, BG85b, BG86, BG87, BG88, BG89, BG81b, SG89]. whose [OA86]. width [LKK81]. Wilcoxon [MMW85]. wildlife [RSG88]. Wilks [Sin82]. Winsorized [FR80]. wise [PM87]. Wishart [Jon85, TG82]. Within [Ano87d]. without [GG87]. Work [Far89].

x [Klo88, Wil82].

young [Ano86a].

Z [Som89]. Zacks [AA84]. zero [KK82a, Par88]. Zone [CP89].

References

Abel:1984:CZR

[AA84] Volker Abel and Rudolf Avenhaus. Comments on Zacks’ recent paper on one-sided cusum procedures. Communications in Statistics: Simulation and Computation, 13(2):289–292, 1984. CODEN CSSCDB. ISSN 0361-0918.

Alkuzweny:1988:SSB

[AA88] Baha M. D. Alkuzweny and Donald A. Anderson. A simulation study of bias in estimation of variance by bootstrap linear regression model. Communications in Statistics: Simulation and Computation, 17(3):871–886, 1988. CODEN CSSCDB. ISSN 0361-0918.

Abraham:1989:RR

[AAA^{+}89] Bovas Abraham, A. Adatia, James Albert, Saeid B. Amini, Graig F. Ansley, N. Balakrishnan, Charles K. Bayne, Kenneth N. Berk, Kenneth J. Berry, C. Champ, John P. Chandler, Yogendra P. Chaubey, Peyton Cook, Ram C. Dahiya, Dipak Dey, A. H. Feiveson, William H. Fellner, Thomas B. Fomby, Y. Fujikoshi, D. V. Gokhale, Joel B. Greenhouse, Paul A. Games, Ramesh C. Gupta, Irwin Guttman, D. L. Hawkins, W. W. Hauck, James J. Higgins, Paul Hinz, C. T. Isaki, S. Jeyaratnam, Michael P. Jones, Robert Keener, R. M. Korwar, H. K. Lam, Kwan Lee, Huaixiang Li, B. D. Macpherson, Henrick J. Malik, B. F. J. Manly, Edward R. Mansfield, Nathan Mantel, R. D. Martin, John I. McCool, Robert Mee, H. Moors, N. Mukhopadhyay, P. B. Nagarsenker, Peter R. Nelson, M. R. Nenno, S. K. Perng, Poduri S. R. S. Rao, J. C. W. Rayner, Roch Roy, R. T. Rust, M. Anthony Schork, Noriaki Seto, L. R. Shenton, G. R. Shorack, J. J. Shuster, D. G. Simpson, Eric V. Slud, L. Strijbosch, William H. Swallow, Georgia Thompson, M. L. Tiku, Steven Vardeman, Y. Vardi, William Warde, Sanford Weisberg, Thomas K. Witt, Dennis A. Wolf, Robert A. Wolfe, W. A. Woodward, and Linda J. Young. Referee recognition. Communications in Statistics: Simulation and Computation, 18(4):1603–1605, 1989. CODEN CSSCDB. ISSN 0361-0918.

Armstrong:1986:BSP

[AB86] Ronald D. Armstrong and Philip O. Beck. The best subset of parameters in min-max linear regression. Communications in Statistics: Simulation and Computation, 15(1):263–270, 1986. CODEN CSSCDB. ISSN 0361-0918.

Al-Baidhani:1987:CME

[ABS87] P. A. Al-Baidhani and C. D. Sinclair. Comparison of methods of estimation of parameters of the Weibull distribution. Communications in Statistics: Simulation and Computation, 16(2):373–384, 1987. CODEN CSSCDB. ISSN 0361-0918.

Anderson:1988:SAA

[AD88] Oliver D. Anderson and Jan G. De Gooijer. Sampled autocovariance and autocorrelation results for linear time processes. Communications in Statistics: Simulation and Computation, 17(2):489– 513, 1988. CODEN CSSCDB. ISSN 0361-0918.

Adatia:1989:MQR

[Ada89] A. Adatia. Moments of quasi-range from a gamma distribution. Communications in Statistics: Simulation and Computation, 18 (1):263–279, 1989. CODEN CSSCDB. ISSN 0361-0918.

Al-Hussaini:1984:IMM

[AHA84] E. K. Al-Hussaini and K. E. Ahmad. Information matrix for a mixture of two inverse Gaussian distributions. Communications in Statistics: Simulation and Computation, 13(6):785–800, 1984. CODEN CSSCDB. ISSN 0361-0918.

Alle:1984:RTT

[AHS84] M. Alle, R. Haux, and M. Schumacher. Rank tests and ties — some properties of a rank test for complete block designs. Communications in Statistics: Simulation and Computation, 13(2):183–212, 1984. CODEN CSSCDB. ISSN 0361-0918.

Allen:1983:CPC

[AJ83] David M. Allen and David C. Jordan. A computational procedure for combining data and prior information. Communications in Statistics: Simulation and Computation, 12(4):389–398, 1983. CODEN CSSCDB. ISSN 0361-0918.

Alvin:1986:ACI

[AJG86] James A. Alvin, S. Jeyaratnam, and Franklin A. Graybill. Approximate confidence intervals for the three-factor mixed model. Communications in Statistics: Simulation and Computation, 15 (4):893–903, 1986. CODEN CSSCDB. ISSN 0361-0918.

Alemayehu:1988:BLR

[Ale88] Demissie Alemayehu. Bootstrapping the latent roots of certain random matrices. Communications in Statistics: Simulation and Computation, 17(3):857–869, 1988. CODEN CSSCDB. ISSN 0361-0918.

AliMousa:1988:SRL

[Ali88] M. A. M. Ali Mousa. Studying the risk of the linear empirical Bayes estimate of the binomial parameter p. Communications in Statistics: Simulation and Computation, 17(1):137–152, 1988. CODEN CSSCDB. ISSN 0361-0918.

Allen:1986:RII

[All86] O. B. Allen. Recovery of interblock information for incomplete block designs. Communications in Statistics: Simulation and Computation, 15(1):207–214, 1986. CODEN CSSCDB. ISSN 0361-0918.

Ammann:1989:RPC

[Amm89] Larry P. Ammann. Robust principal components. Communications in Statistics: Simulation and Computation, 18(3):857–874, 1989. CODEN CSSCDB. ISSN 0361-0918.

Anderson:1988:GVF

[And88] Charles Anderson. Gamma variates of fractional shape as functionals of a homogeneous multidimensional Poisson process. Communications in Statistics: Simulation and Computation, 17(3): 781–787, 1988. CODEN CSSCDB. ISSN 0361-0918.

Angus:1988:ILM

[Ang88] John E. Angus. Inferences on the lognormal mean for complete samples. Communications in Statistics: Simulation and Computation, 17(4):1307–1331, 1988. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBa

[Ano80a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(1):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBb

[Ano80b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(2):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBc

[Ano80c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(3):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBd

[Ano80d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(4):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBe

[Ano80e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(5):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1980:EBf

[Ano80f] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 9(6):ebi–ebii, 1980. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:E

[Ano81a] Anonymous. Editorial. Communications in Statistics: Simulation and Computation, 10(3):iii, 1981. CODEN CSSCDB. ISSN 0361-

Anonymous:1981:EBa

[Ano81b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 10(1):ebi–ebii, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:EBb

[Ano81c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 10(4):ebi–ebii, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:EBc

[Ano81d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 10(5):ebi–ebii, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:EBd

[Ano81e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 10(6):ebi–ebii, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:EE

[Ano81f] Anonymous. Editorial editorial. Communications in Statistics: Simulation and Computation, 10(2):3–4, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:I

[Ano81g] Anonymous. Integral. Communications in Statistics: Simulation and Computation, 10(5):539, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:KCT

[Ano81h] Anonymous. K38. Corrections to the table of normal integrals. Communications in Statistics: Simulation and Computation, 10 (5):541, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1981:MM

[Ano81i] Anonymous. Minicommunications minicommunications. Communications in Statistics: Simulation and Computation, 10(5):529, 1981. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBa

[Ano82a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(1):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBb

[Ano82b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(2):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBc

[Ano82c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(3):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBd

[Ano82d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(4):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBe

[Ano82e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(5):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:EBf

[Ano82f] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 11(6):ebi–ebii, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:MC

[Ano82g] Anonymous. Mini communication. Communications in Statistics: Simulation and Computation, 11(2):227, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1982:M

[Ano82h] Anonymous. Minicomputers. Communications in Statistics: Simulation and Computation, 11(5):635, 1982. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBa

[Ano83a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(1):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBb

[Ano83b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(2):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBc

[Ano83c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(3):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBd

[Ano83d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(4):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBe

[Ano83e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(5):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:EBf

[Ano83f] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 12(6):ebi–ebii, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1983:M

[Ano83g] Anonymous. Minicommunications. Communications in Statistics: Simulation and Computation, 12(4):501, 1983. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBa

[Ano84a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(1):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBb

[Ano84b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(2):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBc

[Ano84c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(3):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBd

[Ano84d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(4):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBe

[Ano84e] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(5):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:EBf

[Ano84f] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 13(6):ebi–ebii, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:Ma

[Ano84g] Anonymous. Minicommunications. Communications in Statistics: Simulation and Computation, 13(2):287, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1984:Mb

[Ano84h] Anonymous. Minicommunications. Communications in Statistics: Simulation and Computation, 13(5):717, 1984. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1985:EBa

[Ano85a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 14(1):ebi–ebii, 1985. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1985:EBb

[Ano85b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 14(2):ebi–ebii, 1985. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1985:EBc

[Ano85c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 14(3):ebi–ebii, 1985. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1985:EBd

[Ano85d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 14(4):ebi–ebii, 1985. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1986:ACY

[Ano86a] Anonymous. Announcement: Competitions for young statisticians from developing countries 1987 and 1989. Communications in Statistics: Simulation and Computation, 15(1):289–290, 1986. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1986:PTS

[Ano86b] Anonymous. On a problem in two-stage sampling. Communications in Statistics: Simulation and Computation, 15(3):889–891, 1986. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1987:EBa

[Ano87a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 16(1):ebi–ebii, 1987. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1987:EBb

[Ano87b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 16(3):ebi–ebii, 1987. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1987:EBc

[Ano87c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 16(4):ebi–ebii, 1987. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1987:WRP

[Ano87d] Anonymous. Within row pairwise comparisons in a two-way ANOVA design. Communications in Statistics: Simulation and Computation, 16(4):939–955, 1987. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1988:EBa

[Ano88a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 17(1):ebi–ebii, 1988. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1988:EBb

[Ano88b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 17(2):ebi–ebii, 1988. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1988:EBc

[Ano88c] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 17(3):ebi–ebii, 1988. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1988:EBd

[Ano88d] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 17(4):ebi–ebii, 1988. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1989:EBa

[Ano89a] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 18(1):ebi–ebii, 1989. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1989:EBb

[Ano89b] Anonymous. Editorial Board. Communications in Statistics: Simulation and Computation, 18(4):ebi–ebii, 1989. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:1989:R

[Ano89c] Anonymous. Rejoinder. Communications in Statistics: Simulation and Computation, 18(4):1249–1250, 1989. CODEN CSSCDB. ISSN 0361-0918.

Aranda-Ordaz:1987:REK

[AO87] Francisco J. Aranda-Ordaz. Relative efficiency of the Kaplan– Meier estimator under contamination. Communications in Statistics: Simulation and Computation, 16(4):987–997, 1987. CODEN CSSCDB. ISSN 0361-0918.

Antti:1983:CBP

[AP83] Kanto Antti and Simo Puntanen. A connection between the partial correlation coefficient and the correlation coefficient of certain residuals. Communications in Statistics: Simulation and Computation, 12(5):639–641, 1983. CODEN CSSCDB. ISSN 0361-0918.

Akhtar:1986:RSD

[AP86] Munir Akhtar and Philip Prescott. Response surface designs robust to missing observations. Communications in Statistics: Simulation and Computation, 15(2):345–363, 1986. CODEN CSSCDB. ISSN 0361-0918.

Amirkhalkhali:1986:PRA

[AR86] Saleh Amirkhalkhali and U. L. Gouranga Rao. Prediction in regression with autocorrelated normal and non-normal errors. Communications in Statistics: Simulation and Computation, 15(1): 185–205, 1986. CODEN CSSCDB. ISSN 0361-0918.

Abd-Rahman:1985:ELM

[ARG85] M. N. Abd-Rahman and T. M. Gerig. An estimation in a linear model with response-related error. Communications in Statistics: Simulation and Computation, 14(1):71–83, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Armstrong:1985:MEG

[Arm85] Ben Armstrong. Measurement error in the generalised linear model. Communications in Statistics: Simulation and Computation, 14(3):529–544, 1985. CODEN CSSCDB. ISSN 0361-0918.

Aroian:1980:TSD

[Aro80] Leo A. Aroian. Time series in m-dimensions definition, problems and prospects. Communications in Statistics: Simulation and Computation, 9(5):453–465, 1980. CODEN CSSCDB. ISSN 0361-0918.

Awad:1986:TAR

[AS86] Adnan M. Awad and Mohammed A. Shayib. Tail area revisited. Communications in Statistics: Simulation and Computation, 15 (4):1215–1234, 1986. CODEN CSSCDB. ISSN 0361-0918.

Awad:1987:CST

[AS87] Adnan M. Awad and Talib H. Sarie. A comparative study of tests of independence. Communications in Statistics: Simulation and Computation, 16(3):645–671, 1987. CODEN CSSCDB. ISSN 0361-0918.

Awad:1985:LSP

[ASD85] Adnan M. Awad, Mohammed A. Shayib, and Ahmad M. Dawagreh. Large sample prediction intervals for future geometric mean. A comparative study. Communications in Statistics: Simulation and Computation, 14(4):983–1006, 1985. CODEN CSSCDB. ISSN 0361-0918.

Aguirre-Torres:1989:CBT

[ATAGD89] Victor Aguirre-Torres, A., R. Gallant, and Jorge Domínguez. On choosing between two nonlinear models estimated robustly. Some Monte Carlo evidence. Communications in Statistics: Simulation and Computation, 18(1):179–200, 1989. CODEN CSSCDB. ISSN 0361-0918.

Ammann:1989:SRO

[AV89] Larry Ammann and John Van Ness. Standard and robust orthogonal regression. Communications in Statistics: Simulation and Computation, 18(1):145–162, 1989. CODEN CSSCDB. ISSN 0361-0918.

Amini:1988:SSP

[AW88] Saeid B. Amini and Robert F. Woolson. Small-sample properties of covariance-adjusted survivorship data tests for treatment effect. Communications in Statistics: Simulation and Computation, 17 (4):1281–1306, 1988. CODEN CSSCDB. ISSN 0361-0918.

Bustos:1984:EPF

[BA84] A. Bustos and M. Ahsanullah. On the estimates of the parameters of the first order autoregressive process. Communications in Statistics: Simulation and Computation, 13(4):489–505, 1984. CODEN CSSCDB. ISSN 0361-0918.

Bai:1988:NTR

[Bai88] Zhaojun Bai. Numerical treatment of restricted Gauss–Markov model. Communications in Statistics: Simulation and Computation, 17(2):569–579, 1988. CODEN CSSCDB. ISSN 0361-0918.

Bakir:1989:AMU

[Bak89] Saad T. Bakir. Analysis of means using ranks. Communications in Statistics: Simulation and Computation, 18(2):757–776, 1989. CODEN CSSCDB. ISSN 0361-0918.

Barker:1983:GMD

[Bar83] Lawrence Barker. On Gini’s mean difference and the sample standard deviation. Communications in Statistics: Simulation and Computation, 12(4):503–505, 1983. CODEN CSSCDB. ISSN 0361-0918.

Barker:1985:RPU

[Bar85] Lawrence E. Barker. Range-preserving unbiased estimators in the Poisson case. Communications in Statistics: Simulation and Computation, 14(4):1029–1030, 1985. CODEN CSSCDB. ISSN 0361-0918.

Baxter:1981:SRN

[Bax81] Laurence A. Baxter. Some remarks on numerical convolution. Communications in Statistics: Simulation and Computation, 10 (3):281–288, 1981. CODEN CSSCDB. ISSN 0361-0918.

Bukac:1980:ASC

[BB80] Josef Bukac and Herman Burstein. Approximations of Student’s t and chi-square percentage points. Communications in Statistics: Simulation and Computation, 9(6):665–672, 1980. CODEN CSSCDB. ISSN 0361-0918.

Birch:1983:ENN

[BB83] Jeffrey B. Birch and Doris A. Binkley. Effects of non-normality and mild heteroscedasticity on estimators in regression. Communications in Statistics: Simulation and Computation, 12(3):331–354, 1983. CODEN CSSCDB. ISSN 0361-0918.

Bradley:1986:ATV

[BB86] Ralph A. Bradley and Dennis A. Brindley. An approximate test of variance homogeneity based on Grubbs’ estimators. Communications in Statistics: Simulation and Computation, 15(1):27–34, 1986. CODEN CSSCDB. ISSN 0361-0918.

Berkane:1987:CPM

[BB87] Maia Berkane and P. M. Bentler. Characterizing parameters of multivariate elliptical distributions. Communications in Statistics: Simulation and Computation, 16(1):193–198, 1987. CODEN CSSCDB. ISSN 0361-0918.

Bayne:1984:MPS

[BBK84] C. K. Bayne, J. J. Beauchamp, and V. E. Kane. Misclassification probabilities for second-order discriminant functions used to classify bivariate normal populations. Communications in Statistics: Simulation and Computation, 13(5):669–682, 1984. CODEN CSSCDB. ISSN 0361-0918.

Barton:1989:HTM

[BC89] Curtis N. Barton and Elliot C. Cramer. Hypothesis testing in multivariate linear models with randomly missing data. Communications in Statistics: Simulation and Computation, 18(3):875–895, 1989. CODEN CSSCDB. ISSN 0361-0918.

Brockett:1980:MDE

[BCC80] P. L. Brockett, A. Charnes, and W. W. Cooper. M.D.I. estimation via unconstrained convex programming. Communications in Statistics: Simulation and Computation, 9(3):223–234, 1980. CO-DEN CSSCDB. ISSN 0361-0918.

Brooks:1988:RLR

[BCHJ88] Camilla A. Brooks, Rhonda R. Clark, Alula Hadgu, and Arthur M. Jones. The robustness of the logistic risk function. Communications in Statistics: Simulation and Computation, 17(1):1–24, 1988. CODEN CSSCDB. ISSN 0361-0918.

Broffitt:1981:MEL

[BCL81] Barbara Broffitt, William R. Clarke, and Peter A. Lachenbruch. Measurement errors — a location contamination problem in discriminant analysis. Communications in Statistics: Simulation and Computation, 10(2):129–141, 1981. CODEN CSSCDB. ISSN 0361-0918.

Baringhaus:1989:RCT

[BDH89] L. Baringhaus, R. Danschke, and N. Henze. Recent and classical tests for normality — a comparative study. Communications in Statistics: Simulation and Computation, 18(1):363–379, 1989. CODEN CSSCDB. ISSN 0361-0918.

Becher:1988:OED

[Bec88] Heiko Becher. On optimal experimental design under spatial correlation structures for square and nonsquare plot designs. Communications in Statistics: Simulation and Computation, 17(3):771–780, 1988. CODEN CSSCDB. ISSN 0361-0918.

Berenson:1982:SSU

[Ber82] Mark L. Berenson. A study of several useful tests for ordered alternatives in the randomized block design. Communications in Statistics: Simulation and Computation, 11(5):563–581, 1982. CODEN CSSCDB. ISSN 0361-0918.

Best:1981:BML

[BF81] D. J. Best and N. I. Fisher. The bias of the maximum likelihood estimators of the von Mises–Fisher concentration parameters. Communications in Statistics: Simulation and Computation, 10(5):493–502, 1981. CODEN CSSCDB. ISSN 0361-0918.

Birch:1984:SPP

[BF84] Jeffrey B. Birch and Robert V. Foutz. Selected percent points for a test for the two-sample problem based on empirical probability measures. Communications in Statistics: Simulation and Computation, 13(3):397–405, 1984. CODEN CSSCDB. ISSN 0361-0918.

Birch:1985:ACT

[BF85] Jeffrey B. Birch and Robert V. Foutz. An algorithm for computing a two-sample test based on empirical probability measures. Communications in Statistics: Simulation and Computation, 14 (1):249–252, 1985. CODEN CSSCDB. ISSN 0361-0918.

Berenson:1981:DCR

[BG81a] Mark L. Berenson and Shulamith T. Gross. Designing completely randomized experiments to detect ordered treatment effects: an empirical study. Communications in Statistics: Simulation and Computation, 10(4):405–431, 1981. CODEN CSSCDB. ISSN 0361-0918.

Bland:1981:MDU

[BG81b] R. P. Bland and Shulamith T. Gross. M34. On the definition of unbiasedness for estimating parameters which are random variables. Communications in Statistics: Simulation and Computation, 10 (4):435–436, 1981. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1985:RAS

[BG85a] Robert E. Bechhofer and David M. Goldsman. On the Ramey– Alam sequential procedure for selecting the multinomial event which has the largest probability. Communications in Statistics: Simulation and Computation, 14(2):263–282, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Bechhofer:1985:TBK

[BG85b] Robert E. Bechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the multinomial event which has the largest probability. Communications in Statistics: Simulation and Computation, 14(2):283–315, 1985. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1986:TBK

[BG86] Robert E. Bechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the multinomial event which has the largest probability (II): extended tables and an improved procedure. Communications in Statistics: Simulation and Computation, 15(3):829–851, 1986. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1987:TBK

[BG87] Robert E. Bechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the normal population which has the largest mean. Communications in Statistics: Simulation and Computation, 16(4):1067–1092, 1987. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1988:TBK

[BG88] Robert E. Bechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the normal population which has the largest mean (ii): 2-factor experiments with no interaction. Communications in Statistics: Simulation and Computation, 17(1):103–128, 1988. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1989:TBK

[BG89] Robert E. Bechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the normal population which has the largest mean (III): Supplementary truncation numbers and resulting performance characteristics. Communications in Statistics: Simulation and Computation, 18(4):1457–1464, 1989. CODEN CSSCDB. ISSN 0361-0918.

Bechhofer:1989:SSS

[BGJ89] Robert E. Bechhofer, David M. Goldsman, and Christopher Jennison. A single-stage selection procedure for multi-factor multinomial experiments with multiplicativity. Communications in Statistics: Simulation and Computation, 18(1):31–61, 1989. CODEN CSSCDB. ISSN 0361-0918.

Bratcher:1984:BAR

[BHP84] T. L. Bratcher, A. Hobbs, and J. Paul. A balanced approach to region estimation with tables for the normal model. Communications in Statistics: Simulation and Computation, 13(6):801–821, 1984. CODEN CSSCDB. ISSN 0361-0918.

Bates:1983:CIP

[BHW83] Douglas M. Bates, David C. Hamilton, and Donald G. Watts. Calculation of intrinsic and parameter-effects curvatures for nonlinear regression models. Communications in Statistics: Simulation and Computation, 12(4):469–477, 1983. CODEN CSSCDB. ISSN 0361-0918.

Buonaccorsi:1984:CCR

[BI84] John P. Buonaccorsi and Hariharan K. Iyer. A comparison of confidence regions and designs in estimation of a ratio. Communications in Statistics: Simulation and Computation, 13(6):723–741, 1984. CODEN CSSCDB. ISSN 0361-0918.

Bialas:1981:BEA

[Bia81] Wayne F. Bialas. Bounds for the error in an approximation to the distribution of continuous random vectors with bounded supports. Communications in Statistics: Simulation and Computation, 10 (1):67–76, 1981. CODEN CSSCDB. ISSN 0361-0918.

Birch:1980:ESV

[Bir80] Jeffrey B. Birch. Effects of the starting value and stopping rule on robust estimates obtained by iterated weighted least squares. Communications in Statistics: Simulation and Computation, 9(2): 141–154, 1980. CODEN CSSCDB. ISSN 0361-0918.

Birch:1983:PRT

[Bir83] Jeffrey B. Birch. On the power of robust tests in analysis of covariance. Communications in Statistics: Simulation and Computation, 12(2):159–182, 1983. CODEN CSSCDB. ISSN 0361-0918.

Binns:1985:SSN

[BJ85] M. R. Binns and P. Y. Jui. A simulation study of the nearest neighbour analysis method of Papadakis. Communications in Statistics: Simulation and Computation, 14(1):159–172, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Brown:1985:EMO

[BK85] Keith C. Brown and K. Rao Kadiyala. The estimation of missing observations in related time series data: Further results. Communications in Statistics: Simulation and Computation, 14(4):973– 981, 1985. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:1988:MVC

[BL88a] N. Balakrishnan and M. Y. Leung. Means, variances and covariances of order statistics BLUE’s for the Type I generalized logistic. Communications in Statistics: Simulation and Computation, 17(1):51–84, 1988. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:1988:OST

[BL88b] N. Balakrishnan and M. Y. Leung. Order statistics from the Type I generalized logistic distribution. Communications in Statistics: Simulation and Computation, 17(1):25–50, 1988. CODEN CSS-CDB. ISSN 0361-0918.

Bowman:1986:SSN

[BLS86] K. O. Bowman, H. K. Lam, and L. R. Shenton. Series for Student’s non-central t under exponential sampling with comments due to H. P. Mulholland. Communications in Statistics: Simulation and Computation, 15(3):697–708, 1986. CODEN CSSCDB. ISSN 0361-0918.

Bates:1987:GRG

[BLWY87] Douglas M. Bates, Mary J. Lindstrom, Grace Wahba, and Brian S. Yandell. GCVPACK — routines for generalized cross validation. Communications in Statistics: Simulation and Computation, 16 (1):263–297, 1987. CODEN CSSCDB. ISSN 0361-0918.

Berry:1983:CFP

[BM83] Kenneth J. Berry and Paul W. Mielke, Jr. Computation of finite population parameters and approximate probability values for multi-response permutation procedures (MRPP). Communications in Statistics: Simulation and Computation, 12(1):83–107, 1983. CODEN CSSCDB. ISSN 0361-0918.

Berry:1984:CEP

[BM84] Kenneth J. Berry and Paul W. Mielke. Computation of exact probability values for multi-response permutation procedures (MRPP). Communications in Statistics: Simulation and Computation, 13 (3):417–432, 1984. CODEN CSSCDB. ISSN 0361-0918.

Berry:1985:CEA

[BM85] Kenneth J. Berry and Paul W. Mielke. Computation of exact and approximate probability values for a matched pairs permutation test. Communications in Statistics: Simulation and Computation, 14(1):229–248, 1985. CODEN CSSCDB. ISSN 0361-0918.

Baczkowski:1987:ALS

[BM87] A. J. Baczkowski and K. V. Mardia. Approximate lognormality of the sample semi-variogram under a Gaussian process. Communications in Statistics: Simulation and Computation, 16(2):571–585, 1987. CODEN CSSCDB. ISSN 0361-0918.

Bradley:1988:GAL

[BO88] Ralph A. Bradley and Robert E. Odeh. A generating algorithm for linear trend-free and nearly linear trend-free block designs. Communications in Statistics: Simulation and Computation, 17 (4):1259–1280, 1988. CODEN CSSCDB. ISSN 0361-0918.

Boik:1985:NAD

[Boi85] Robert J. Boik. A new approximation to the distribution function of the Studentized maximum root. Communications in Statistics: Simulation and Computation, 14(3):759–767, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Bolstad:1986:EAH

[Bol86] William M. Bolstad. An efficient algorithm for Harrison–Stevens forecasting using the multi-process multivariate dynamic linear model. Communications in Statistics: Simulation and Computation, 15(3):819–828, 1986. CODEN CSSCDB. ISSN 0361-0918.

Bolstad:1987:EMS

[Bol87] William M. Bolstad. An estimation method for the seemingly unrelated regression model with contemporaneous covariances based on an efficient recursive algorithm. Communications in Statistics: Simulation and Computation, 16(3):689–698, 1987. CODEN CSS-CDB. ISSN 0361-0918.

Bolfarine:1988:FPP

[Bol88] Heleno Bolfarine. Finite population prediction under dynamic generalized linear models. Communications in Statistics: Simulation and Computation, 17(1):187–207, 1988. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1984:CPF

[Bow84] K. O. Bowman. Computation of the polygamma functions. Communications in Statistics: Simulation and Computation, 13(3): 409–415, 1984. CODEN CSSCDB. ISSN 0361-0918.

Beauchamp:1986:TCD

[BR86] J. J. Beauchamp and D. S. Robson. Transformation considerations in discriminant analysis. Communications in Statistics: Simulation and Computation, 15(1):147–179, 1986. CODEN CSSCDB. ISSN 0361-0918.

Bremer:1989:NSS

[Bre89] Ronald Bremer. Numerical study of small sample variances of estimators of variance components in the two-way factorial model. Communications in Statistics: Simulation and Computation, 18 (3):985–1009, 1989. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1980:EPR

[BS80] K. O. Bowman and L. R. Shenton. Evaluation of the parameters of s_{u} by rational fractions. Communications in Statistics: Simulation and Computation, 9(2):127–132, 1980. CODEN CSSCDB. ISSN 0361-0918.

Baker:1981:EDFa

[BS81a] T. C. Baker, Jr. and R. L. Sielken, Jr. Estimation of a distribution function by extrapolating upper and lower bounds. Communications in Statistics: Simulation and Computation, 10(1):77–93, 1981. CODEN CSSCDB. ISSN 0361-0918.

Baker:1981:EDFb

[BS81b] Thomas C. Baker, Jr. and Robert L. Sielken, Jr. Estimating the distribution function of a transformed random vector. Communications in Statistics: Simulation and Computation, 10(2):109–128, 1981. CODEN CSSCDB. ISSN 0361-0918.

Bownan:1982:PEG

[BS82] K. O. Bownan and L. R. Shenton. Properties of estimators for the gamma distribution. Communications in Statistics: Simulation and Computation, 11(4):377–519, 1982. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1983:MLE

[BS83] K. O. Bowman and L. R. Shenton. Maximum likelihood estimators for the gamma distribution revisited. Communications in Statistics: Simulation and Computation, 12(6):697–710, 1983. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1985:DME

[BS85] K. O. Bowman and L. R. Shenton. The distribution of a moment estimator for a parameter of the generalized Poisson distribution. Communications in Statistics: Simulation and Computation, 14 (4):867–893, 1985. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1986:TCU

[BS86] K. O. Bowman and L. R. Shenton. The two-color urn model: recursions for the moments. Communications in Statistics: Simulation and Computation, 15(3):787–799, 1986. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1987:SPB

[BS87] K. O. Bowman and L. R. Shenton. Sums of powers of binomial coefficients. Communications in Statistics: Simulation and Computation, 16(4):1189–1207, 1987. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1988:SJ

[BS88] K. O. Bowman and L. R. Shenton. Solutions to Johnson’s S_{B} and S_{U}. Communications in Statistics: Simulation and Computation, 17(2):343–348, 1988. CODEN CSSCDB. ISSN 0361-0918.

Blackhurst:1989:LRP

[BS89a] Dawn W. Blackhurst and Mark D. Schluchter. Logistic regression with a partially observed covariate. Communications in Statistics: Simulation and Computation, 18(1):163–177, 1989. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1989:DFP

[BS89b] K. O. Bowman and L. R. Shenton. S_{B} and S_{U} distributions fitted by percentiles: a general criterion. Communications in Statistics: Simulation and Computation, 18(1):1–13, 1989. CODEN CSS-CDB. ISSN 0361-0918.

Blair:1987:LRT

[BSH87] R. Clifford Blair, Shlomo S. Sawilowsky, and James J. Higgins. Limitations of the rank transform statistic in tests for interactions. Communications in Statistics: Simulation and Computation, 16 (4):1133–1145, 1987. CODEN CSSCDB. ISSN 0361-0918.

Basilevsky:1985:MDE

[BSHA85] Alexander Basilevsky, Donald Sabourin, Derek Hum, and Andy Anderson. Missing data estimators in the general linear model: an evaluation of simulated data as an experimental design. Communications in Statistics: Simulation and Computation, 14(2):371–394, 1985. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1987:SEP

[BSL87] K. O. Bowman, L. R. Shenton, and H. K. Lam. Simulation and estimation problems associated with the 3-parameter gamma density. Communications in Statistics: Simulation and Computation, 16(4):1147–1188, 1987. CODEN CSSCDB. ISSN 0361-0918.

Bowman:1981:EAS

[BSS81] K. O. Bowman, C. A. Serbin, and L. R. Shenton. Explicit approximate solutions for S_{B}. Communications in Statistics: Simulation and Computation, 10(1):1–15, 1981. CODEN CSSCDB. ISSN 0361-0918.

Bugaighis:1989:CSB

[Bug89] Mohamed M. Bugaighis. Comparative study of BLU and ML estimators of location and scale parameters of an extreme value distribution for small samples and Type I censorship. Communications in Statistics: Simulation and Computation, 18(2):601–611, 1989. CODEN CSSCDB. ISSN 0361-0918.

Bates:1984:MRG

[BW84a] Douglas M. Bates and Donald G. Watts. A multi-response Gauss– Newton algorithm. Communications in Statistics: Simulation and Computation, 13(5):705–715, 1984. CODEN CSSCDB. ISSN 0361-0918.

Bates:1984:NNR

[BW84b] Douglas M. Bates and Dennis A. Wolf. Non-negative regression by Givens rotations. Communications in Statistics: Simulation and Computation, 13(6):841–850, 1984. CODEN CSSCDB. ISSN 0361-0918.

Betro:1987:MCS

[BZ87] Bruno Betrò and Ryszard Zieliński. A Monte Carlo study of a Bayesian decision rule concerning the number of different values of a discrete random variable. Communications in Statistics: Simulation and Computation, 16(4):925–938, 1987. CODEN CSSCDB. ISSN 0361-0918.

Cameron:1981:ENC

[Cam81] M. A. Cameron. Estimation of noise correlations in transfer function models. Communications in Statistics: Simulation and Computation, 10(4):369–381, 1981. CODEN CSSCDB. ISSN 0361-0918.

Campbell:1988:CPN

[Cam88] Robert A. Campbell. A comparison of parametric and nonparametric tests for detecting queue waiting time differences. Communications in Statistics: Simulation and Computation, 17(2):463– 470, 1988. CODEN CSSCDB. ISSN 0361-0918.

Carnahan:1989:MLE

[Car89] J. V. Carnahan. Maximum likelihood estimation for the 4parameter beta distribution. Communications in Statistics: Simulation and Computation, 18(2):513–536, 1989. CODEN CSSCDB. ISSN 0361-0918.

Causey:1985:ECS

[Cau85] Beverley D. Causey. Exact calculations for sequential tests based on Bernoulli trials. Communications in Statistics: Simulation and Computation, 14(2):491–495, 1985. CODEN CSSCDB. ISSN 0361-0918.

Causey:1986:EAD

[Cau86] Beverley D. Causey. Expected absolute departure of chi-square from its median. Communications in Statistics: Simulation and Computation, 15(1):181–183, 1986. CODEN CSSCDB. ISSN 0361-0918.

Cran:1985:MCM

[CB85] G. W. Cran and A. J. Bertie. Minicommunication computation of moments of generalized ridge estimators. Communications in Statistics: Simulation and Computation, 14(3):771–774, 1985. CODEN CSSCDB. ISSN 0361-0918.

Christensen:1980:SRR

[CBR80] Mark J. Christensen and A. T. Bharucha-Reid. Stability of the roots of random algebraic polynomials. Communications in Statistics: Simulation and Computation, 9(2):179–192, 1980. CODEN CSSCDB. ISSN 0361-0918.

Chatterjee:1983:EMP

[CC83] Samprit Chatterjee and Sangit Chatterjee. Estimation of misclassification probabilities by bootstrap methods by bootstrap methods. Communications in Statistics: Simulation and Computation, 12(6):645–656, 1983. CODEN CSSCDB. ISSN 0361-0918.

Chen:1985:SAT

[CCS85] Hubert J. Chen, Shun-Yi Chen, and Jirawan Sirichote. Selecting all treatments better than a control with a binomial prior distribution. Communications in Statistics: Simulation and Computation, 14(1):187–221, 1985. CODEN CSSCDB. ISSN 0361-0918.

Chan:1983:SIG

[CCW83] Micah Y. Chan, A. Clifford Cohen, and Betty Jones Whitten. The standardized inverse Gaussian distribution tables of the cumulative probability function. Communications in Statistics: Simulation and Computation, 12(4):423–442, 1983. CODEN CSSCDB. ISSN 0361-0918.

Chan:1984:MML

[CCW84] Micah Y. Chan, A. Clifford Cohen, and Betty Jones Whitten. Modified maximum likelihood and modified moment estimators for the three-parameter inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 13(1):47–68, 1984. CODEN CSSCDB. ISSN 0361-0918.

Chaubey:1987:ERM

[CD87] Yogendra P. Chaubey and Tryambakeshwar D. Dwivedi. Efficiency of regression method in estimating hazard rate from grouped survival data. Communications in Statistics: Simulation and Computation, 16(3):761–769, 1987. CODEN CSSCDB. ISSN 0361-0918.

Chakraborti:1988:GMM

[CD88] S. Chakraborti and M. M. Desu. Generalizations of Mathisen’s median test for comparing several treatments with a control. Communications in Statistics: Simulation and Computation, 17(3):947– 967, 1988. CODEN CSSCDB. ISSN 0361-0918.

Chun:1989:TET

[CDY89] Calvin S. Y. Chun, Olive Jean Dunn, and Mimi C. Yu. Tests on the equality of two multiple correlation coefficients. Communications in Statistics: Simulation and Computation, 18(3):909–928, 1989. CODEN CSSCDB. ISSN 0361-0918.

Chhikara:1980:ECV

[CF80] R. S. Chhikara and A. L. Feiveson. Extended critical values of extreme Studentized deviate test statistics for detecting multiple outliers. Communications in Statistics: Simulation and Computation, 9(2):155–166, 1980. CODEN CSSCDB. ISSN 0361-0918.

Chen:1986:UPD

[CG86] James J. Chen and David W. Gaylor. The upper percentiles of the distribution of the logrank statistic for small numbers of tumors. Communications in Statistics: Simulation and Computation, 15 (4):991–1002, 1986. CODEN CSSCDB. ISSN 0361-0918.

Chaubey:1984:CSN

[Cha84] Yogendra P. Chaubey. On the comparison of some non-negative estimators of variance components for two models. Communications in Statistics: Simulation and Computation, 13(5):619–633, 1984. CODEN CSSCDB. ISSN 0361-0918.

Chauhan:1988:CSR

[Cha88] C. K. Chauhan. A class of single replicate orthogonal designs. Communications in Statistics: Simulation and Computation, 17 (2):537–547, 1988. CODEN CSSCDB. ISSN 0361-0918.

Chen:1988:CIS

[Che88] Pinyuen Chen. Closed inverse sampling procedure for selecting the largest multinomial cell probability. Communications in Statistics: Simulation and Computation, 17(3):969–994, 1988. CODEN CSS-CDB. ISSN 0361-0918.

Chou:1981:ATN

[Cho81] Youn-Min Chou. Additions to the table of normal integrals minicommunications. Communications in Statistics: Simulation and Computation, 10(5):537–538, 1981. CODEN CSSCDB. ISSN 0361-0918.

Chuang:1987:AOP

[Chu87] Christy Chuang. Analyzing an ordinal pharmaceutical data set with a categorical covariate using monotone-scores models. Communications in Statistics: Simulation and Computation, 16(1):1– 15, 1987. CODEN CSSCDB. ISSN 0361-0918.

Cox:1985:TET

[CJ85] T. F. Cox and K. Jaber. Testing the equality of two normal percentiles. Communications in Statistics: Simulation and Computation, 14(2):345–356, 1985. CODEN CSSCDB. ISSN 0361-0918.

Calvin:1986:ACI

[CJG86] James A. Calvin, S. Jeyaratnam, and Franklin A. Graybill. Approximate confidence intervals for the three-factor mixed model. Communications in Statistics: Simulation and Computation, 15 (4):893–903, 1986. CODEN CSSCDB. ISSN 0361-0918.

Carmer:1983:TER

[CL83] S. G. Carmer and W. T. Lin. Type I error rates for divisive clustering methods for grouping means in analysis of variance.

Communications in Statistics: Simulation and Computation, 12 (4):451–466, 1983. CODEN CSSCDB. ISSN 0361-0918.

Callfas:1986:NND

[CM86] D. Callfas and S. G. Mohanty. On non-null distribution of ranks. Communications in Statistics: Simulation and Computation, 15 (3):591–608, 1986. CODEN CSSCDB. ISSN 0361-0918.

Charek:1988:CET

[CMC88] Dennis J. Charek, Albert H. Moore, and Joseph W. Coleman. A comparison of estimation techniques for the three parameter Pareto distribution. Communications in Statistics: Simulation and Computation, 17(4):1395–1407, 1988. CODEN CSSCDB. ISSN 0361-0918.

Collings:1987:SDF

[Col87] Bruce Jay Collings. String decomposition of full-period Tausworthe sequences. Communications in Statistics: Simulation and Computation, 16(3):673–678, 1987. CODEN CSSCDB. ISSN 0361-0918.

Consul:1986:DTG

[Con86] P. C. Consul. On the differences of two generalized Poisson variates. Communications in Statistics: Simulation and Computation, 15(3):761–767, 1986. CODEN CSSCDB. ISSN 0361-0918.

Cooke:1983:SGP

[Coo83] William P. Cooke. Surrogate geometric programming estimates of restricted multinomial proportions. Communications in Statistics: Simulation and Computation, 12(3):291–305, 1983. CODEN CSSCDB. ISSN 0361-0918.

Chung:1989:SMC

[CP89] Jain Chung and Terry Obert Pittman. A simple method for computing the ARL of [ILM0001]-charts with zone tests. Communications in Statistics: Simulation and Computation, 18(4):1275–1293, 1989. CODEN CSSCDB. ISSN 0361-0918.

Cui:1988:CRV

[CR88] Rong-Quan Cui and Marion R. Reynolds, Jr. -chart with runs and variable sampling intervals. Communications in Statistics: Simulation and Computation, 17(3):1073–1093, 1988. CODEN CSSCDB. ISSN 0361-0918.

Consul:1985:GPD

[CS85] P. C. Consul and M. M. Shoukri. The generalized Poisson distribution when the sample mean is larger than the sample variance. Communications in Statistics: Simulation and Computation, 14 (3):667–681, 1985. CODEN CSSCDB. ISSN 0361-0918.

Consul:1986:NIM

[CS86] P. Consul and M. M. Shoukri. The negative integer moments of the generalized Poisson distribution. Communications in Statistics: Simulation and Computation, 15(4):1053–1064, 1986. CODEN CSSCDB. ISSN 0361-0918.

Cohen:1982:MMM

[CW82] A. Clifford Cohen and Betty Jones Whitten. Modified moment and maximum likelihood estimators for parameters of the threeparameter gamma distribution. Communications in Statistics: Simulation and Computation, 11(2):197–216, 1982. CODEN CSS-CDB. ISSN 0361-0918.

Costello:1985:NNA

[CW85] Patricia S. Costello and Douglas A. Wolfe. A new nonparametric approach to the problem of agreement between two groups of judges. Communications in Statistics: Simulation and Computation, 14(4):791–805, 1985. CODEN CSSCDB. ISSN 0361-0918.

Cooray-Wijesinha:1987:SGM

[CWK87] Manel Cooray-Wijesinha and Andrew I. Khuri. The sequential generation of multiresponse D-optimal designs when the variancecovariance matrix is not known. Communications in Statistics: Simulation and Computation, 16(1):239–259, 1987. CODEN CSS-CDB. ISSN 0361-0918.

Chhikara:1988:CIF

[CyD88] Raj S. Chhikara and Lih yuan Deng. Conditional inference in finite population sampling under a calibration model. Communications in Statistics: Simulation and Computation, 17(2):663–681, 1988. CODEN CSSCDB. ISSN 0361-0918.

Dagpunar:1988:CGR

[Dag88] J. S. Dagpunar. Computer generation of random variates from the tail of t and normal distributions. Communications in Statistics: Simulation and Computation, 17(2):653–661, 1988. CODEN CSSCDB. ISSN 0361-0918.

Dagpunar:1989:EIG

[Dag89] J. S. Dagpunar. An easily implemented generalised inverse Gaussian generator. Communications in Statistics: Simulation and Computation, 18(2):703–710, 1989. CODEN CSSCDB. ISSN 0361-0918.

Danaher:1989:SME

[Dan89] Peter J. Danaher. Simulating media exposure distributions. Communications in Statistics: Simulation and Computation, 18(4): 1381–1392, 1989. CODEN CSSCDB. ISSN 0361-0918.

Davenport:1980:CSG

[Dav80a] James M. Davenport. The CORFIS subroutine for the generation of selected level percentage points of the Pearson Type VI distribution. Communications in Statistics: Simulation and Computation, 9(4):429–452, 1980. CODEN CSSCDB. ISSN 0361-0918.

Davis:1980:FTH

[Dav80b] A. W. Davis. Further tabulation of Hotelling’s generalized. Communications in Statistics: Simulation and Computation, 9(4):321– 336, 1980. CODEN CSSCDB. ISSN 0361-0918.

Dwivedi:1981:MRT

[DC81] T. D. Dwivedi and Y. P. Chaubey. Moments of a ratio of two positive quadratic forms in normal variables. Communications in Statistics: Simulation and Computation, 10(5):503–516, 1981. CODEN CSSCDB. ISSN 0361-0918.

Doornbos:1983:MST

[DD83] R. Doornbos and J. B. Dijkstra. A multi sample test for the equality of coefficients of variation in normal populations. Communications in Statistics: Simulation and Computation, 12(2):147–158, 1983. CODEN CSSCDB. ISSN 0361-0918.

DeLosReyes:1989:OLF

[De 89] J. P. De Los Reyes. Optimizing linear functions of random variables having a joint multinomial or multivariate normal distribution. Communications in Statistics: Simulation and Computation, 18(3):835–856, 1989. CODEN CSSCDB. ISSN 0361-0918.

Deutler:1984:SEC

[Deu84] T. Deutler. A series expansion for the cumulants of the χdistribution and a Cornish–Fisher-expansion for the noncentrality parameter of the noncentral t-distribution. Communications in Statistics: Simulation and Computation, 13(4):507–513, 1984. CODEN CSSCDB. ISSN 0361-0918.

Dexter:1986:SMI

[Dex86] Franklin Dexter. A statistical measure of internal consistency. Communications in Statistics: Simulation and Computation, 15 (3):879–886, 1986. CODEN CSSCDB. ISSN 0361-0918.

Dion:1982:RBC

[DF82] L. R. Dion and D. Fridshal. A relationship between chi-square and F -critical values. Communications in Statistics: Simulation and Computation, 11(2):233–235, 1982. CODEN CSSCDB. ISSN 0361-0918.

deGooijer:1988:SSO

[dGS88] Jan G. de Gooijer and Pentti Saikkonen. A specification strategy for order determination in ARMA models. Communications in Statistics: Simulation and Computation, 17(3):1037–1054, 1988. CODEN CSSCDB. ISSN 0361-0918.

Draper:1985:FOR

[DH85] Norman R. Draper and Agnes M. Herzberg. Fourth order rotatability. Communications in Statistics: Simulation and Computation, 14(3):515–528, 1985. CODEN CSSCDB. ISSN 0361-0918.

Dudel:1989:PEA

[DHL89] Helmut P. Dudel, Charles E. Hall, Jr., and Siegfried H. Lehnigk. Parameter estimation algorithms for the hyper-gamma distribution class. Communications in Statistics: Simulation and Computation, 18(3):1135–1153, 1989. CODEN CSSCDB. ISSN 0361-0918.

Dietz:1987:CRE

[Die87] E. Jacquelin Dietz. A comparison of robust estimators in simple linear regression. Communications in Statistics: Simulation and Computation, 16(4):1209–1227, 1987. CODEN CSSCDB. ISSN 0361-0918.

Donner:1983:NAF

[DK83] Allan Donner and John J. Koval. A note on the accuracy of Fisher’s approximation to the large sample variance of an intraclass correlationa. Communications in Statistics: Simulation and Computation, 12(4):443–449, 1983. CODEN CSSCDB. ISSN 0361-0918.

DeDominicis:1984:AAN

[DM84] Rodolfo De Dominicis and Raimondo Manca. An algorithmic approach to non-homogeneous semi-Markov processes. Communications in Statistics: Simulation and Computation, 13(6):823–838, 1984. CODEN CSSCDB. ISSN 0361-0918.

Dhakar:1989:DSF

[DM89] Tej Dhakar and T. H. Mattheiss. Determining the size of a finite population of different objects from a finite sample taken at random with replacement. Communications in Statistics: Simulation and Computation, 18(4):1311–1323, 1989. CODEN CSSCDB. ISSN 0361-0918.

Donegani:1989:SPC

[Don89] M. Donegani. Simulated power curves for some location tests. Communications in Statistics: Simulation and Computation, 18 (4):1393–1400, 1989. CODEN CSSCDB. ISSN 0361-0918.

Dielman:1988:BLA

[DP88] Terry E. Dielman and Roger C. Pfaffenberger. Bootstrapping in least absolute value regression: an application to hypothesis testing. Communications in Statistics: Simulation and Computation, 17(3):843–856, 1988. CODEN CSSCDB. ISSN 0361-0918.

Deaton:1983:EHT

[DRM83] Michael L. Deaton, Mation R. Reynolds, Jr., and Raymond H. Myers. Estimation and hypothesis testing in regression in the presence of nonhomogeneous error variances. Communications in Statistics: Simulation and Computation, 12(1):45–66, 1983. CO-DEN CSSCDB. ISSN 0361-0918.

Deutsch:1980:ESD

[DS80] Stuart Jay Deutsch and Bruce Wayne Schmeiser. Estimation of the sum of differences distribution. Communications in Statistics: Simulation and Computation, 9(6):563–587, 1980. CODEN CSSCDB. ISSN 0361-0918.

Dunn:1980:VPD

[DT80] J. E. Dunn and J. D. Tubbs. VARSTAB: a procedure for determining homoscedastic transformations of multivariate normal populations. Communications in Statistics: Simulation and Computation, 9(6):589–598, 1980. CODEN CSSCDB. ISSN 0361-0918.

Dijkstra:1981:TES

[DW81] Jan B. Dijkstra and Paul S. P. J. Werter. Testing the equality of several means when the population variances are unequal. Communications in Statistics: Simulation and Computation, 10 (6):557–569, 1981. CODEN CSSCDB. ISSN 0361-0918.

deWaal:1988:PSM

[dWN88] D. J. de Waal and D. G. Nel. A procedure to select a ML-II prior in a multivariate normal case. Communications in Statistics: Simulation and Computation, 17(3):1021–1035, 1988. CODEN CSSCDB. ISSN 0361-0918.

Emrich:1989:SCP

[EB89] Lawrence J. Emrich and Mark F. Brady. Stochastic curtailment and power calculations for long-term clinical trials via simulation. Communications in Statistics: Simulation and Computation, 18 (1):383–395, 1989. CODEN CSSCDB. ISSN 0361-0918.

Elperin:1987:MLE

[EG87] T. Elperin and I. Gertsbakh. Maximum likelihood estimation in a Weibull regression model with Type-1 censoring: a Monte Carlo study. Communications in Statistics: Simulation and Computation, 16(2):349–371, 1987. CODEN CSSCDB. ISSN 0361-0918.

Embrechts:1986:IAP

[EHA86] Paul Embrechts, Agnes M. Herzberg, and C. K. Allen. An investigation of Andrews’ plots to detect period and outliers in time series data. Communications in Statistics: Simulation and Computation, 15(4):1027–1051, 1986. CODEN CSSCDB. ISSN 0361-0918.

Eastham:1987:SSP

[ELS87] Jerome F. Eastham, Vincent N. LaRiccia, and John H. Schenemeyer. Small sample properties of the maximum likelihood estimators for an alternative parameterization of the three-parameter lognormal distribution. Communications in Statistics: Simulation and Computation, 16(3):871–884, 1987. CODEN CSSCDB. ISSN 0361-0918.

Emerson:1981:ECR

[Eme81] John D. Emerson. Effects of censoring on the robustness of exponential-based confidence intervals for median lifetime. Communications in Statistics: Simulation and Computation, 10(6): 617–627, 1981. CODEN CSSCDB. ISSN 0361-0918.

Eberhardt:1989:CFE

[EMR89] Keith R. Eberhardt, Robert W. Mee, and Charles P. Reeve. Computing factors for exact two-sided tolerance limits for a normal distribution. Communications in Statistics: Simulation and Computation, 18(1):397–413, 1989. CODEN CSSCDB. ISSN 0361-0918.

Engelman:1982:EAC

[Eng82] Laszlo Engelman. An efficient algorithm for computing covariance matrices from data with missing values. Communications in Statistics: Simulation and Computation, 11(1):113–121, 1982. CODEN CSSCDB. ISSN 0361-0918.

Edgeman:1988:MKS

[ESP88] Rick L. Edgeman, Robert C. Scott, and Robert J. Pavur. A modified Kolmogorov–Smirnov test for the inverse Gaussian density with unknown parameters. Communications in Statistics: Simulation and Computation, 17(4):1203–1212, 1988. CODEN CSSCDB. ISSN 0361-0918.

Evans:1987:TOA

[Eva87] Merran A. Evans. The twelfth order analogue of the Durbin– Watson test. Communications in Statistics: Simulation and Computation, 16(3):679–688, 1987. CODEN CSSCDB. ISSN 0361-0918.

Elliott:1986:AUT

[EW86] Alan C. Elliott and Wayne A. Woodward. Analysis of an unbalanced two-way ANOVA on the microcomputer. Communications in Statistics: Simulation and Computation, 15(1):215–225, 1986. CODEN CSSCDB. ISSN 0361-0918.

Farebrother:1989:SEW

[Far89] R. W. Farebrother. Some early work on the duality between points and lines. Communications in Statistics: Simulation and Computation, 18(2):719–727, 1989. CODEN CSSCDB. ISSN 0361-0918.

F:1987:UBT

[FB87] Schuster Eugene F. and Richard C. Barker. Using the bootstrap in testing symmetry versus asymmetry. Communications in Statistics: Simulation and Computation, 16(1):69–84, 1987. CODEN CSSCDB. ISSN 0361-0918.

Federer:1987:FRS

[Fed87] Walter T. Federer. Fractional replication in simulation studies. Communications in Statistics: Simulation and Computation, 16 (1):233–237, 1987. CODEN CSSCDB. ISSN 0361-0918.

Feingold:1982:NWC

[Fei82] Marcia Feingold. A note on weights for combining intra- and interblock estimates in balanced incomplete block designs. Communications in Statistics: Simulation and Computation, 11(5):645–648, 1982. CODEN CSSCDB. ISSN 0361-0918.

Fellner:1987:SME

[Fel87] William H. Fellner. Sparse matrices, and the estimation of variance components by likelihood methods. Communications in Statistics: Simulation and Computation, 16(2):439–463, 1987. CODEN CSS-CDB. ISSN 0361-0918.

Freed:1983:MIP

[FG83] N. Freed and F. Glover. A mixed-integer programming approach to the clustering problem. Communications in Statistics: Simulation and Computation, 12(5):595–607, 1983. CODEN CSSCDB. ISSN 0361-0918.

Firinguetti:1989:SSR

[Fir89] Luis L. Firinguetti. A simulation study of ridge regression estimators with autocorrelated errors. Communications in Statistics: Simulation and Computation, 18(2):673–702, 1989. CODEN CSS-CDB. ISSN 0361-0918.

Fang:1986:ADT

[FK86] C. Fang and P. R. Krishnaiah. On asymptotic distribution of the test statistic for the mean of the non-isotropic principal component. Communications in Statistics: Simulation and Computation, 15(4):1163–1168, 1986. CODEN CSSCDB. ISSN 0361-0918.

Fagerstrom:1983:BAB

[FM83] Richard M. Fagerstrom and Duane A. Meeter. A Bayesian approach to bioassay. Technical report M498. Communications in Statistics: Simulation and Computation, 12(5):541–558, 1983. CODEN CSSCDB. ISSN 0361-0918.

Fomby:1987:SSE

[Fom87] Thomas B. Fomby. Small sample efficiency gains from a first observation correction for Hatanaka’s estimator of the lagged dependent variable-serial correlation regression model. Communications in Statistics: Simulation and Computation, 16(2):551–570, 1987. CODEN CSSCDB. ISSN 0361-0918.

Fung:1985:COD

[FP85] Karen Yuen Fung and S. R. Paul. Comparisons of outlier detection procedures in Weibull or extreme-value distribution. Communications in Statistics: Simulation and Computation, 14(4):895–917, 1985. CODEN CSSCDB. ISSN 0361-0918.

Fung:1980:TSW

[FR80] Karen Yuen Fung and Sheikh M. Rahman. The two-sample Winsorized t. Communications in Statistics: Simulation and Computation, 9(4):337–347, 1980. CODEN CSSCDB. ISSN 0361-0918.

Franklin:1989:NEA

[Fra89] L. A. Franklin. A note on the Edgeworth approximation to the distribution of Spearman’s rho with a correction to Pearson’s approximation. Communications in Statistics: Simulation and Computation, 18(1):245–252, 1989. CODEN CSSCDB. ISSN 0361-0918.

Fridshal:1988:POI

[Fri88] Donald Fridshal. The probability of an ordering of independent uniform random variables. Communications in Statistics: Simulation and Computation, 17(2):685–687, 1988. CODEN CSSCDB. ISSN 0361-0918.

Frosini:1988:CVM

[Fro88] Benito V. Frosini. Characterizations of variability measures. Communications in Statistics: Simulation and Computation, 17(4): 1459–1481, 1988. CODEN CSSCDB. ISSN 0361-0918.

Fawcett:1984:MCS

[FS84] R. F. Fawcett and K. C. Salter. A Monte Carlo study of the F test and three tests based on ranks of treatment effects in randomized block designs. Communications in Statistics: Simulation and Computation, 13(2):213–225, 1984. CODEN CSSCDB. ISSN 0361-0918.

Fujii:1989:TEM

[Fuj89] Yoshinori Fujii. Test for the equality of marginal distributions on positively dependent bivariate survival data. Communications in Statistics: Simulation and Computation, 18(2):633–642, 1989. CODEN CSSCDB. ISSN 0361-0918.

Frome:1980:SEP

[FY80] E. L. Frome and G. J. Yakatan. Statistical estimation of the pharmokinetic parameters in the one compartment open model. Communications in Statistics: Simulation and Computation, 9(3): 201–222, 1980. CODEN CSSCDB. ISSN 0361-0918.

Gans:1981:UPT

[Gan81] D. J. Gans. Use of a preliminary test in comparing two sample means. Communications in Statistics: Simulation and Computation, 10(2):163–174, 1981. CODEN CSSCDB. ISSN 0361-0918.

Gbur:1981:PID

[Gbu81] Edward E. Gbur. On the Poisson index of dispersion. Communications in Statistics: Simulation and Computation, 10(5):531–535, 1981. CODEN CSSCDB. ISSN 0361-0918.

Gbur:1989:EMG

[GC89a] Edward E. Gbur and Robert A. Collins. Estimation of the moment generating function. Communications in Statistics: Simulation and Computation, 18(3):1113–1134, 1989. CODEN CSSCDB. ISSN 0361-0918.

Gunel:1989:EPB

[GC89b] Erdogan Günel and Daniel Chilko. Estimation of parameter n of the binomial distribution. Communications in Statistics: Simulation and Computation, 18(2):537–551, 1989. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:1987:ISR

[GG87] Subir Ghosh and Roberto Gomez. Interpenetrating subsampling regression estimation with or without jackknifing. Communications in Statistics: Simulation and Computation, 16(4):1105–1116, 1987. CODEN CSSCDB. ISSN 0361-0918.

Giani:1987:DSE

[Gia87] Guido Giani. Designing selection experiments with Bernoulli populations. Communications in Statistics: Simulation and Computation, 16(2):535–549, 1987. CODEN CSSCDB. ISSN 0361-0918.

Gillespie:1984:ITL

[Gil84] John A. Gillespie. Inner tolerance limits controlling tail proportions for the normal distribution. Communications in Statistics: Simulation and Computation, 13(2):257–267, 1984. CODEN CSS-CDB. ISSN 0361-0918.

Gupta:1989:PRTa

[GK89a] Ramesh C. Gupta and S. N. U. A. Kirmani. On predicting repair times in a minimal repair process. Communications in Statistics: Simulation and Computation, 18(4):1359–1368, 1989. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1989:PRTb

[GK89b] Ramesh C. Gupta and S. N. U. A. Kirmani. On predicting repair times in a minimal repair process. Communications in Statistics: Simulation and Computation, 18(4):1359–1368, 1989. CODEN CSSCDB. ISSN 0361-0918.

Gould:1988:EEL

[GL88] Ann Gould and J. F. Laeless. Estimation efficiency in lifetime regression models when responses are censored or grouped. Communications in Statistics: Simulation and Computation, 17(3):689– 712, 1988. CODEN CSSCDB. ISSN 0361-0918.

Gnanadesikan:1983:SAU

[Gna83] Mrudulla Gnanadesikan. Statistical analysis of the usage of a computer system. Communications in Statistics: Simulation and Computation, 12(6):657–674, 1983. CODEN CSSCDB. ISSN 0361-0918.

Gonin:1986:NAS

[Gon86] R. Gonin. Numerical algorithms for solving nonlinear l-norm estimation problems: Part I — a first-order gradient algorithm for well-conditioned small residual problems. Communications in Statistics: Simulation and Computation, 15(3):801–813, 1986. CODEN CSSCDB. ISSN 0361-0918.

Gosling:1985:RER

[GP85] Barbara J. Gosling and Martin L. Puterman. Ridge estimation in regression problems with autocorrelated errors: A Monte Carlo study. Communications in Statistics: Simulation and Computation, 14(3):577–613, 1985. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1985:DSM

[GPS85] Shanti S. Gupta, S. Panchapakesan, and Joong K. Sohn. On the distribution of the Studentized maximum of equally correlated normal random variables. Communications in Statistics: Simulation and Computation, 14(1):103–135, 1985. CODEN CSSCDB. ISSN 0361-0918.

Greenwood:1981:PFA

[Gre81] J. Arthur Greenwood. A portable formulation of the alias method for random numbers with discrete distributions. Communications in Statistics: Simulation and Computation, 10(6):649–655, 1981. CODEN CSSCDB. ISSN 0361-0918.

Grove:1986:PAT

[Gro86] Daniel M. Grove. Positive association in a two-way contingency table: a numerical study. Communications in Statistics: Simulation and Computation, 15(3):633–648, 1986. CODEN CSSCDB. ISSN 0361-0918.

Groggel:1987:MCS

[Gro87] D. J. Groggel. A Monte Carlo study of rank tests for block designs. Communications in Statistics: Simulation and Computation, 16 (3):601–620, 1987. CODEN CSSCDB. ISSN 0361-0918.

Gupta:1981:DRT

[GS81] R. P. Gupta and C. Singh. Distributions of the ratio of two ranges of samples from nonnormal populations. Communications in Statistics: Simulation and Computation, 10(6):629–637, 1981. CODEN CSSCDB. ISSN 0361-0918.

George:1987:MFC

[GS87] E. Olusegun George and Meenakshi Sivaram. A modification of the Fisher-Cornish approximation for the Student t percentiles. Communications in Statistics: Simulation and Computation, 16 (4):1123–1132, 1987. CODEN CSSCDB. ISSN 0361-0918.

Guenther:1982:NTS

[Gue82] William C. Guenther. Normal theory sample size formulas for some nonnormal distributions. Communications in Statistics: Simulation and Computation, 11(6):727–732, 1982. CODEN CSS-CDB. ISSN 0361-0918.

Groggel:1988:NEO

[GWR88] David J. Groggel, D. D. Wackerly, and P. V. Rao. Nonparametric estimation in one-way random effects models. Communications in Statistics: Simulation and Computation, 17(3):887–903, 1988. CODEN CSSCDB. ISSN 0361-0918.

Hauck:1989:CCE

[HA89a] Walter W. Hauck and Sharon Anderson. Comment on “Considerations for Effective Simulation Study Analyses”. Communications in Statistics: Simulation and Computation, 18(4):1245–1248, 1989. CODEN CSSCDB. ISSN 0361-0918. See [LF89].

Hung:1989:EMV

[HA89b] Ken Hung and Frank B. Alt. The effect of misspecification in vector autoregressive moving average models on parameter estimation and forecasting. Communications in Statistics: Simulation and Computation, 18(2):467–479, 1989. CODEN CSSCDB. ISSN 0361-0918.

Haber:1987:CSC

[Hab87] Michael Haber. A comparison of some conditional and unconditional exact tests for 2 × 2 contingency tables. Communications in Statistics: Simulation and Computation, 16(4):999–1013, 1987. CODEN CSSCDB. ISSN 0361-0918.

Hamilton:1980:IAE

[Ham80] Martin Alva Hamilton. Inference about the ED50 using the trimmed Spearman–Karber procedure — a Monte Carlo investigation. Communications in Statistics: Simulation and Computation, 9(3):235–254, 1980. CODEN CSSCDB. ISSN 0361-0918.

Hapuarachchi:1988:SEA

[Hap88] K. P. Hapuarachchi. Sampling errors of analytic statistics: an empirical investigation. Communications in Statistics: Simulation and Computation, 17(2):549–567, 1988. CODEN CSSCDB. ISSN 0361-0918.

Harter:1984:AFC

[Har84] H. Leon Harter. Asymptotic formulas for critical values of a modified Kolmogorov test statistic. Communications in Statistics: Simulation and Computation, 13(5):719–721, 1984. CODEN CSSCDB. ISSN 0361-0918.

Hawkins:1989:SPP

[Haw89] D. L. Hawkins. Some practical problems in implementing a certain sieve estimator of the Gaussian mean function. Communications in Statistics: Simulation and Computation, 18(2):481–500, 1989. CODEN CSSCDB. ISSN 0361-0918.

Hemmerle:1983:SPG

[HC83] William J. Hemmerle and Michele Boulanger Carey. Some properties of generalized ridge estimators. Communications in Statistics: Simulation and Computation, 12(3):239–253, 1983. CODEN CSS-CDB. ISSN 0361-0918.

Higazi:1984:CSE

[HD84] Sohair M. F. Higazi and C. Mitchell Dayton. Comparing several experimental groups with a control in the multivariate case. Communications in Statistics: Simulation and Computation, 13 (2):227–241, 1984. CODEN CSSCDB. ISSN 0361-0918.

Higazi:1988:TME

[HD88] Sohair M. F. Higazi and C. Mitchell Dayton. Tables for a multivariate extension of the Dunnett test when the control group and balanced experimental groups have different sample sizes. Communications in Statistics: Simulation and Computation, 17(1):85– 101, 1988. CODEN CSSCDB. ISSN 0361-0918.

Helland:1988:SPL

[Hel88] Inge S. Helland. On the structure of partial least squares regression. Communications in Statistics: Simulation and Computation, 17(2):581–607, 1988. CODEN CSSCDB. ISSN 0361-0918.

Henery:1988:ACD

[Hen88] R. J. Henery. Asymptotic convergence in distribution for Spearman’s rank correlation coefficient. Communications in Statistics: Simulation and Computation, 17(4):1449–1452, 1988. CODEN CSSCDB. ISSN 0361-0918.

Huang:1989:RDA

[HF89] Mei-Ling Huang and Karen Yuen Fung. R-distribution and its applications. Communications in Statistics: Simulation and Computation, 18(1):99–119, 1989. CODEN CSSCDB. ISSN 0361-0918.

Horton:1985:MEC

[HG85] Raymond L. Horton and John B. Guerard. The management of executive compensation in large, dynamic firms: A further look. Communications in Statistics: Simulation and Computation, 14 (2):441–448, 1985. CODEN CSSCDB. ISSN 0361-0918.

Hawkins:1986:SLS

[HGF86] D. L. Hawkins, A. R. Gallant, and W. Fuller. A simple least squares method for estimating a change in mean. Communications in Statistics: Simulation and Computation, 15(3):523–530, 1986. CODEN CSSCDB. ISSN 0361-0918.

Hedayat:1983:AGB

[HH83] A. S. Hedayat and H. L. Hwang. An algorithm for generating a basis of the trades on t-designs. Communications in Statistics: Simulation and Computation, 12(1):109–125, 1983. CODEN CSS-CDB. ISSN 0361-0918.

Hawkins:1986:PAC

[HH86] D. L. Hawkins and Chien-Pai Han. Power of analysis of covariance tests under conditional specification. Communications in Statistics: Simulation and Computation, 15(2):323–343, 1986. CODEN CSSCDB. ISSN 0361-0918.

Hosmer:1983:CML

[HHF83a] Trina Hosmer, David Hosmer, and Lloyd Fisher. A comparison of the maximum likelihood and discriminant function estimators of the coefficients of the logistic regression model for mixed continuous and discrete variables. Communications in Statistics: Simulation and Computation, 12(1):23–43, 1983. CODEN CSSCDB. ISSN 0361-0918.

Hosmer:1983:CTM

[HHF83b] Trina A. Hosmer, David W. Hosmer, and Lloyd Fisher. A comparison of three methods of estimating the logistic regression coefficients. Communications in Statistics: Simulation and Computation, 12(5):577–593, 1983. CODEN CSSCDB. ISSN 0361-0918.

Huang:1988:VEU

[HI88] Elizabeth T. Huang and Cary T. Isaki. Variance estimation using auxiliary information under a one unit per stratum sample design. Communications in Statistics: Simulation and Computation, 17 (4):1431–1447, 1988. CODEN CSSCDB. ISSN 0361-0918.

Hietikko:1981:NBC

[Hie81] Harri Hietikko. On the numerical behaviour of ARMA(P, Q) covariance determinants for various sample sizes. Communications in Statistics: Simulation and Computation, 10(5):451–463, 1981. CODEN CSSCDB. ISSN 0361-0918.

Hoerl:1980:MNL

[HK80] Arthur E. Hoerl and Robert W. Kennard. M30. A note on least squares estimates. Communications in Statistics: Simulation and Computation, 9(3):315–317, 1980. CODEN CSSCDB. ISSN 0361-0918.

Harter:1984:MKS

[HKL84] H. Leon Harter, Harry J. Khamis, and Richard E. Lamb. Modified Kolmogorov–Smirnov tests of goodness of fit. Communications in Statistics: Simulation and Computation, 13(3):293–323, 1984. CODEN CSSCDB. ISSN 0361-0918.

Homan:1986:REE

[HL86] Sharon M. Homan and Peter A. Lachenbruch. Robust estimation of the exponential mean parameter for small samples: Complete and censored data. Communications in Statistics: Simulation and Computation, 15(4):1087–1108, 1986. CODEN CSSCDB. ISSN 0361-0918.

Ha:1989:PAM

[HN89] Cu D. Ha and Subhash C. Narula. Perturbation analysis for the minimum sum of absolute errors regression. Communications in Statistics: Simulation and Computation, 18(3):957–970, 1989. CODEN CSSCDB. ISSN 0361-0918.

Holms:1980:CPM

[Hol80] Arthur G. Holms. “Chain pooling” model selection for two-level fixed effects factorial experiments. Communications in Statistics: Simulation and Computation, 9(1):51–71, 1980. CODEN CSS-CDB. ISSN 0361-0918.

Holmes:1989:RCM

[Hol89] R. B. Holmes. On random correlation matrices IL the Toeplitz case. Communications in Statistics: Simulation and Computation, 18(4):1511–1537, 1989. CODEN CSSCDB. ISSN 0361-0918.

Homan:1989:CPR

[Hom89] Sharon M. Homan. A comparison of plotting rules under L1 and L2 estimation of the Weibull scale and shape parameters in situations of small samples with possible censoring and outliers. Communications in Statistics: Simulation and Computation, 18(1):121–143, 1989. CODEN CSSCDB. ISSN 0361-0918.

Horn:1989:PIM

[Hor89] Paul S. Horn. A prediction interval for m estimators. Communications in Statistics: Simulation and Computation, 18(3):1155–1167, 1989. CODEN CSSCDB. ISSN 0361-0918.

Hua:1984:TCD

[HP84] Tsushung A. Hua and M. Pourahmadi. Tables of cumulative distribution functions and percentiles of the standardized stable random variables. Communications in Statistics: Simulation and Computation, 13(5):571–601, 1984. CODEN CSSCDB. ISSN 0361-0918.

Hand:1980:ULS

[HS80] M. L. Hand and V. A. Sposito. Using the least squares estimator in Chebyshev estimation. Communications in Statistics: Simulation and Computation, 9(1):43–49, 1980. CODEN CSSCDB. ISSN 0361-0918.

Hazilla:1983:CSS

[HS83] Michael Hazilla and V. Kerry Smith. A comparison of the small sample properties of nonlinear two-stage and nonlinear three-stage least squares. Communications in Statistics: Simulation and Computation, 12(3):307–329, 1983. CODEN CSSCDB. ISSN 0361-0918.

Husler:1986:SBE

[HS86] Jürg Hüsler and Michel Schüpbach. On simple block estimators for the parameters of the extreme-value distribution. Communications in Statistics: Simulation and Computation, 15(1):61–76, 1986. CODEN CSSCDB. ISSN 0361-0918.

Huseby:1980:CMV

[HSV80] John R. Huseby, Neil C. Schwertman, and John Van Ryzin. Computation of the mean vector and dispersion matrix for incomplete multivariate data. Communications in Statistics: Simulation and Computation, 9(3):301–309, 1980. CODEN CSSCDB. ISSN 0361-

Hopark:1987:SSS

[HSV87] Dong Hopark, V. Susarla, and J. Van Ryzin. Small sample study on estimators of life distributions and of mean survival times using randomly censored data. Communications in Statistics: Simulation and Computation, 16(1):221–232, 1987. CODEN CSSCDB. ISSN 0361-0918.

Huh:1986:CPP

[Huh86] Moon T. Huh. Computation of percentage points. Communications in Statistics: Simulation and Computation, 15(4):1191–1198, 1986. CODEN CSSCDB. ISSN 0361-0918.

Husler:1987:TSA

[Hüs87] Jürg Hüsler. On the two-sample adaptive distribution-free test. Communications in Statistics: Simulation and Computation, 16 (1):55–68, 1987. CODEN CSSCDB. ISSN 0361-0918.

Hutchinson:1989:SET

[Hut89] M. F. Hutchinson. A stochastic estimator of the trace of the influence matrix for Laplacian smoothing splines. Communications in Statistics: Simulation and Computation, 18(3):1059–1076, 1989. CODEN CSSCDB. ISSN 0361-0918.

Han:1983:CND

[HW83] Chien-Pai Han and John L. G. Wang. Computation of noncentral F distributions with even denominator degrees of freedom. Communications in Statistics: Simulation and Computation, 12 (1):1–9, 1983. CODEN CSSCDB. ISSN 0361-0918.

Huang:1984:TRO

[HW84] Wen-Tao Huang and An-Bang Wang. Tests for randomness of observations that involve two factors. Communications in Statistics: Simulation and Computation, 13(3):325–336, 1984. CODEN CSSCDB. ISSN 0361-0918.

Harter:1985:MCP

[HW85] H. Leon Harter and Rudolf P. Wiegand. A Monte Carlo of plotting positions. Communications in Statistics: Simulation and Computation, 14(2):317–343, 1985. CODEN CSSCDB. ISSN 0361-0918.

Iyer:1983:CFP

[IBM83] H. K. Iyer, K. J. Berry, and P. W. Mielke. Computation of finite population parameters and approximate probability values for multi-response randomized block permutation procedures (MRBP). Communications in Statistics: Simulation and Computation, 12(4):479–499, 1983. CODEN CSSCDB. ISSN 0361-0918.

Iman:1982:DFA

[IC82] Ronald L. Iman and W. J. Conover. A distribution-free approach to inducing rank correlation among input variables. Communications in Statistics: Simulation and Computation, 11(3):311–334, 1982. CODEN CSSCDB. ISSN 0361-0918.

Iman:1982:RCP

[ID82] Ronald L. Iman and James M. Davenport. Rank correlation plots for use with correlated input variables. Communications in Statistics: Simulation and Computation, 11(3):335–360, 1982. CODEN CSSCDB. ISSN 0361-0918.

Irwin:1986:UWL

[IF86] Christopher Irwin and Stephen J. Finch. Use of weighted least squares regression jn simulation studies. Communications in Statistics: Simulation and Computation, 15(2):291–300, 1986. CODEN CSSCDB. ISSN 0361-0918.

Inder:1987:BOL

[Ind87] Brett A. Inder. Bias in the ordinary least squares estimator in the dynamic linear regression model with autocorrelated disturbances. Communications in Statistics: Simulation and Computation, 16 (3):791–815, 1987. CODEN CSSCDB. ISSN 0361-0918.

Indrayan:1981:ATH

[IS81] Abhaya Indrayan and Rustagi Jagdish S. Approximate tests of hypotheses in regression models with grouped data. Communications in Statistics: Simulation and Computation, 10(4):327–341, 1981. CODEN CSSCDB. ISSN 0361-0918.

Jain:1986:GDM

[Jai86] R. K. Jain. On Gini’s diversity measure. Communications in Statistics: Simulation and Computation, 15(4):987–990, 1986. CODEN CSSCDB. ISSN 0361-0918.

Jarjoura:1988:SHR

[Jar88] David Jarjoura. Smoothing hazard rates with cubic splines. Communications in Statistics: Simulation and Computation, 17(2): 377–392, 1988. CODEN CSSCDB. ISSN 0361-0918.

Jeffrey:1980:SCP

[Jef80] Birch B. Jeffrey. Some convergence properties of iterated reweighted least squares in the location model. Communications in Statistics: Simulation and Computation, 9(4):359–369, 1980. CODEN CSSCDB. ISSN 0361-0918.

Johnson:1983:SPT

[JM83] Paulette Johnson and George A. Milliken. A simple procedure for testing linear hypotheses about the parameters of a nonlinear model using weighted least squares. Communications in Statistics: Simulation and Computation, 12(2):135–145, 1983. CODEN CSSCDB. ISSN 0361-0918.

Johnson:1986:CAC

[Joh86] Robert E. Johnson. A convolution algorithm for calculating coefficients of Jensen’s bivariate F distribution. Communications in Statistics: Simulation and Computation, 15(4):1209–1213, 1986. CODEN CSSCDB. ISSN 0361-0918.

Jones:1985:GIW

[Jon85] M. C. Jones. Generating inverse Wishart matrices. Communications in Statistics: Simulation and Computation, 14(2):511–514, 1985. CODEN CSSCDB. ISSN 0361-0918.

Jeon:1986:AMI

[JP86] T. J. Jeon and S. J. Park. Automatic model identification using vector sample autocorrelation function. Communications in Statistics: Simulation and Computation, 15(4):1147–1161, 1986. CODEN CSSCDB. ISSN 0361-0918.

Johnson:1982:JTS

[JRW82] Mark E. Johnson, John S. Ramberg, and Chiang Wang. The Johnson translation system in Monte Carlo studies. Communications in Statistics: Simulation and Computation, 11(5):521–525, 1982. CODEN CSSCDB. ISSN 0361-0918.

Josvanger:1983:NES

[JS83] Lee Ann Josvanger and V. A. Sposito. L_{1} norm estimates for the simple regression problem. Communications in Statistics: Simulation and Computation, 12(2):215–221, 1983. CODEN CSSCDB. ISSN 0361-0918.

Jennings:1988:ECV

[JY88] Linda W. Jennings and Dean M. Young. Extended critical values of the multivariate extreme deviate test for detecting a single spurious observation. Communications in Statistics: Simulation and Computation, 17(4):1359–1373, 1988. CODEN CSSCDB. ISSN 0361-0918.

Kabe:1980:SCP

[Kab80] D. G. Kabe. On some coverage probability maximization problems. Communications in Statistics: Simulation and Computation, 9(1):73–79, 1980. CODEN CSSCDB. ISSN 0361-0918.

Kampke:1988:CVN

[Käm88] Thomas Kämpke. On controlling variates in network simulation. Communications in Statistics: Simulation and Computation, 17 (1):241–249, 1988. CODEN CSSCDB. ISSN 0361-0918.

Kappenman:1987:IDQ

[Kap87] Russell F. Kappenman. Improved distribution quantile estimation. Communications in Statistics: Simulation and Computation, 16 (2):307–320, 1987. CODEN CSSCDB. ISSN 0361-0918.

Krewski:1984:PFP

[KBB84] D. Krewski, J. Brennan, and M. Bickis. The power of the Fisher permutation test in 2 × k tables. Communications in Statistics: Simulation and Computation, 13(4):433–448, 1984. CODEN CSS-CDB. ISSN 0361-0918.

Kleijnen:1985:PWO

[KCV85] J. P. C. Kleijnen, P. Cremers, and F. Van Belle. The power of weighted and ordinary least squares with estimated unequal variances in experimental design. Communications in Statistics: Simulation and Computation, 14(1):85–102, 1985. CODEN CSSCDB. ISSN 0361-0918.

Keramidas:1987:GPC

[KDG87] E. M. Keramidas, S. J. Devlin, and R. Gnanadesikan. A graphical procedure for comparing the principal components of several covariance matrices. Communications in Statistics: Simulation and Computation, 16(1):161–191, 1987. CODEN CSSCDB. ISSN 0361-0918.

Kemp:1988:SAP

[Kem88] Adrienne W. Kemp. Simple algorithms for the Poisson modal cumulative probability. Communications in Statistics: Simulation and Computation, 17(4):1495–1508, 1988. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1985:SMH

[Ken85] Peter E. Kennedy. A suggestion for measuring heteroskedasticity. Communications in Statistics: Simulation and Computation, 14 (4):845–851, 1985. CODEN CSSCDB. ISSN 0361-0918.

Ketellapper:1982:RLS

[Ket82] Ronald H. Ketellapper. The relevance of large sample properties of estimators for the errors-in-variables model: a Monte Carlo study. Communications in Statistics: Simulation and Computation, 11 (5):625–634, 1982. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1980:ASA

[KG80] William J. Kennedy and James E. Gentle. Algorithms section algorithms section. Communications in Statistics: Simulation and Computation, 9(6):673, 1980. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1982:AS

[KG82] William J. Kennedy and James E. Gentle. Algorithms section. Communications in Statistics: Simulation and Computation, 11 (2):237, 1982. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1983:ACS

[KG83a] Wiliam J. Kennedy and James E. Gentle. Algorithms: Communications In Statistics — Simulation and Computation. Communications in Statistics: Simulation and Computation, 12(1):79–82, 1983. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1983:ASCa

[KG83b] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 12(1):77, 1983. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1983:ASCb

[KG83c] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 12(2):213, 1983. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1983:ASCc

[KG83d] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 12(4):467, 1983. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1984:ACS

[KG84a] William J. Kennedy and James E. Gentle. Algorithms communications in statistics,- simulation and computation. Communications in Statistics: Simulation and Computation, 13(2):271–274, 1984. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1984:ASCa

[KG84b] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 13(2):269, 1984. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1984:ASCb

[KG84c] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 13(3):407, 1984. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1984:ASCc

[KG84d] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 13(5):667, 1984. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1984:ASCd

[KG84e] William J. Kennedy and James E. Gentle. Algorithms section coeditors. Communications in Statistics: Simulation and Computation, 13(6):839, 1984. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1985:ASC

[KG85] William J. Kennedy and James E. Gentle. Algorithms section coeditors coeditor. Communications in Statistics: Simulation and Computation, 14(1):223, 1985. CODEN CSSCDB. ISSN 0361-0918.

Kochar:1988:MCS

[KG88] Subhash C. Kochar and R. P. Gupta. A Monte Carlo study of some asymptotically optimal tests of exponentiality against positive aging. Communications in Statistics: Simulation and Computation, 17(3):803–811, 1988. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1989:AS

[KG89a] William J. Kennedy and James E. Gentle. Algorithm section. Communications in Statistics: Simulation and Computation, 18 (1):381, 1989. CODEN CSSCDB. ISSN 0361-0918.

Kennedy:1989:ASC

[KG89b] William J. Kennedy and James E. Gentle. Algorithm section coeditors. Communications in Statistics: Simulation and Computation, 18(2):827, 1989. CODEN CSSCDB. ISSN 0361-0918.

Knapp:1986:ERM

[KGC86] R. G. Knapp, A. J. Gross, and A. B. Cantor. Estimators of the ratio of means of paired survival data. Communications in Statistics: Simulation and Computation, 15(1):85–100, 1986. CODEN CSSCDB. ISSN 0361-0918.

Keating:1983:NCV

[KH83] Jerome P. Keating and Onas L. Hensley. A note on the critical values of the F max-test. Communications in Statistics: Simulation and Computation, 12(3):257–263, 1983. CODEN CSSCDB. ISSN 0361-0918.

Kabib:1989:EST

[KH89] Abdul R. Kabib and Michael R. Harwell. An empirical study of the Type I error rate and power for some selected normal-theory and nonparametric tests of the independence of two sets of variables. Communications in Statistics: Simulation and Computation, 18 (2):793–826, 1989. CODEN CSSCDB. ISSN 0361-0918.

Khattree:1987:SR

[Kha87] Ravindra Khattree. On selection with restriction. Communications in Statistics: Simulation and Computation, 16(4):1093–1103, 1987. CODEN CSSCDB. ISSN 0361-0918.

Kimber:1988:TUL

[Kim88] A. C. Kimber. Testing upper and lower outlier pairs in gamma samples. Communications in Statistics: Simulation and Computation, 17(3):1055–1072, 1988. CODEN CSSCDB. ISSN 0361-0918.

Kitagawa:1984:BAO

[Kit84] Genshiro Kitagawa. Bayesian analysis of outliers via Akaike’s predictive likelihood of a model. Communications in Statistics: Simulation and Computation, 13(1):107–126, 1984. CODEN CSSCDB. ISSN 0361-0918.

Khedr:1982:PTE

[KK82a] Magdy Khedr and S. K. Katti. On a preliminary test estimator for one of the parameters of the log-zero-Poisson distribution. Communications in Statistics: Simulation and Computation, 11 (6):733–751, 1982. CODEN CSSCDB. ISSN 0361-0918.

Krewski:1982:LDE

[KK82b] Daniel Krewski and John Kovar. Low dose extrapolation under single parameter dose response models. Communications in Statistics: Simulation and Computation, 11(1):27–45, 1982. CODEN CSSCDB. ISSN 0361-0918.

Kleijnen:1986:TMA

[KKM86] J. P. C. Kleijnen, G. L. J. Kloppenburg, and F. L. Meeuwsen. Testing the mean of an asymmetric population: Johnson’s modified t test revisited. Communications in Statistics: Simulation and Computation, 15(3):715–732, 1986. CODEN CSSCDB. ISSN 0361-0918.

Lam:1980:RDS

[kL80] Hing kam Lam. Remarks on the distribution of the sample variance in exponential sampling. Communications in Statistics: Simulation and Computation, 9(6):639–647, 1980. CODEN CSSCDB. ISSN 0361-0918.

Koenig:1985:PSS

[KL85] Lloyd W. Koenig and Averill M. Law. A procedure for selecting a subset of size m containing the l best of k independent normal populations, with applications to simulation. Communications in Statistics: Simulation and Computation, 14(3):719–734, 1985. CODEN CSSCDB. ISSN 0361-0918.

Kshirsagar:1989:RRS

[KL89] Anant M. Kshirsagar and Min-Chu Lin. Reinforcement of response surface designs. Communications in Statistics: Simulation and Computation, 18(3):971–984, 1989. CODEN CSSCDB. ISSN 0361-0918.

Kleijnen:1981:SSB

[Kle81] Jack Kleijnen. Small-sample behavior of weighted least squares in experimental design applications. Communications in Statistics: Simulation and Computation, 10(3):303–313, 1981. CODEN CSSCDB. ISSN 0361-0918.

Klefsjo:1986:CST

[Kle86] Bengt Klefsjö. Comments on a simple test for new better than used in expectation1 by Borges, Proschan and Rodrigues. Communications in Statistics: Simulation and Computation, 15(1):285–286, 1986. CODEN CSSCDB. ISSN 0361-0918.

Klimko:1980:NCB

[Kli80] Eugene M. Klimko. Numerical computation of bridge probabilities. Communications in Statistics: Simulation and Computation, 9(2):103–125, 1980. CODEN CSSCDB. ISSN 0361-0918.

Klotz:1988:IBS

[Klo88] Jerome Klotz. An incomplete beta series with argument x(l-x). Communications in Statistics: Simulation and Computation, 17 (4):1333–1337, 1988. CODEN CSSCDB. ISSN 0361-0918.

Khosrovshahi:1987:LAA

[KM87a] G. B. Khosrovshahi and E. S. Mahmoodian. A linear algebraic algorithm for reducing the support size of t-designs and to generate a basis for trades. Communications in Statistics: Simulation and Computation, 16(4):1015–1038, 1987. CODEN CSSCDB. ISSN 0361-0918.

Klein:1987:IDC

[KM87b] John P. Klein and M. L. Moeschberger. Independent or dependent competing risks: does it make a difference. Communications in Statistics: Simulation and Computation, 16(2):507–533, 1987. CODEN CSSCDB. ISSN 0361-0918.

Khosrovshahi:1988:BDV

[KM88] G. B. Khosrovshahi and E. S. Mahmoodian. On BIB designs with various support sizes for v = 9 and k = 3. Communications in Statistics: Simulation and Computation, 17(3):765–770, 1988. CODEN CSSCDB. ISSN 0361-0918.

Klein:1989:RSE

[KM89] John P. Klein and Melvin L. Moeschberger. The robustness of several estimators of the survivorship function with randomly censored data. Communications in Statistics: Simulation and Computation, 18(3):1087–1112, 1989. CODEN CSSCDB. ISSN 0361-

Keller-McNulty:1987:ETW

[KMH87] Allie Keller-McNulty and James J. Higgins. Effect of tail weight and outliers on power and Type-I error of robust permutation tests for location. Communications in Statistics: Simulation and Computation, 16(1):17–35, 1987. CODEN CSSCDB. ISSN 0361-0918.

Keller-McNulty:1986:EFG

[KMK86] Sallie Keller-McNulty and W. J. Kennedy. An error-free generalized matrix inversion and linear least squares method based on bordering. Communications in Statistics: Simulation and Computation, 15(3):769–785, 1986. CODEN CSSCDB. ISSN 0361-0918.

Koehler:1981:IMC

[Koe81] Kenneth J. Koehler. An improvement of a Monte Carlo technique using asymptotic moments with an application to the likelihood ratio statistic. Communications in Statistics: Simulation and Computation, 10(4):343–357, 1981. CODEN CSSCDB. ISSN 0361-0918.

Koutrouvelis:1981:IPE

[Kou81] Ioannis A. Koutrouvelis. An iterative procedure for the estimation of the parameters of stable laws. Communications in Statistics: Simulation and Computation, 10(1):17–28, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Kulkarni:1986:IGP

[KP86] S. R. Kulkarni and S. R. Paranjape. An improved graphical procedure for multivariate-quality control. Communications in Statistics: Simulation and Computation, 15(1):135–146, 1986. CODEN CSSCDB. ISSN 0361-0918.

Khatri:1989:TIC

[KPR89] C. G. Khatri, T. M. Pukkila, and C. Radhakrishna Rao. Testing intraclass correlation coefficients. Communications in Statistics: Simulation and Computation, 18(1):15–30, 1989. CODEN CSS-CDB. ISSN 0361-0918.

Kennard:1981:MDF

[KR81] Robert W. Kennard and John E. Reith. M32. On the distribution of first digits. Communications in Statistics: Simulation and Computation, 10(1):97–98, 1981. CODEN CSSCDB. ISSN 0361-

0918. URL http://www.tandfonline.com/doi/abs/10.1080/
03610918108812195.

Kramer:1984:OLS

[Krä84] W. Krämer. Ordinary least squares estimation of the functional errors-in-variables model with trended data: some Monte Carlo evidence. Communications in Statistics: Simulation and Computation, 13(5):655–665, 1984. CODEN CSSCDB. ISSN 0361-0918.

Krogstad:1989:SMG

[Kro89] Harald E. Krogstad. Simulation of multivariate Gaussian time series. Communications in Statistics: Simulation and Computation, 18(3):929–941, 1989. CODEN CSSCDB. ISSN 0361-0918.

Khatri:1986:TOC

[KRS86] C. G. Khatri, C. Radhakrishna Rao, and Y. N. Sun. Tables for obtaining confidence bounds for realized signal to noise ratio with an estimated discriminant function. Communications in Statistics: Simulation and Computation, 15(1):1–14, 1986. CODEN CSSCDB. ISSN 0361-0918.

Klemm:1980:MLS

[KS80] R. J. Klemm and V. A. Sposito. M31. Least squares solutions over interval restrictions least squares solutions. Communications in Statistics: Simulation and Computation, 9(4):423–425, 1980. CODEN CSSCDB. ISSN 0361-0918.

Khatri:1981:IEC

[KS81] C. G. Khatri and K. R. Shah. Interval estimation of the common mean. Communications in Statistics: Simulation and Computation, 10(2):99–107, 1981. CODEN CSSCDB. ISSN 0361-0918.

Kelton:1986:NPS

[KS86] Christina M. L. Kelton and Marlene A. Smith. Nonlinear programming solutions to the nonstationary Markov model. Communications in Statistics: Simulation and Computation, 15(4):1169–1190, 1986. CODEN CSSCDB. ISSN 0361-0918.

Kurkjian:1987:REE

[KSK87] Badrig M. Kurkjian, George Q. Strong, and Marvin J. Karson. Reliability estimation for the exponential distribution. Communications in Statistics: Simulation and Computation, 16(3):835–853, 1987. CODEN CSSCDB. ISSN 0361-0918.

Kroorienberg:1987:CGF

[KV87] Pieter M. Kroorienberg and Albert Verbeek. Comments on “A generalization of Fisher’s exact test in p × q contingency tables using more concordant relations”. Communications in Statistics: Simulation and Computation, 16(1):301–306, 1987. CODEN CSS-CDB. ISSN 0361-0918. See [Ngu85].

Kraemer:1987:CTH

[KW87] Dale F. Kraemer and Robert F. Woolson. A comparison of tests of homogeneity for sparse contingency tables. Communications in Statistics: Simulation and Computation, 16(2):465–483, 1987. CODEN CSSCDB. ISSN 0361-0918.

Kamps:1989:MLE

[KW89] Udo Kamps and Heinz Weingarten. Maximum likelihood estimation of a Poisson parameter in a model of equioverlapping samples: A simulation study. Communications in Statistics: Simulation and Computation, 18(4):1369–1379, 1989. CODEN CSSCDB. ISSN 0361-0918.

Klotz:1986:SSR

[KYR^{+}86] Jerome Klotz, Rouh-Yun Yu, K. Rai, V. Susarla, and J. Van Ryzin. Small sample relative performance of the spline smooth survival estimator. Communications in Statistics: Simulation and Computation, 15(3):271–298, 1986. CODEN CSSCDB. ISSN 0361-0918.

Lam:1988:ITS

[Lam88] K. Lam. An improved two-stage selection procedure. Communications in Statistics: Simulation and Computation, 17(3):995–1006, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lee:1988:FPC

[LB88] Wonwoo Lee and Jeffrey B. Birch. Fractional principal components regression: a general approach to biased estimators. Communications in Statistics: Simulation and Computation, 17(3): 713–727, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lee:1984:MSM

[Lee84] Sik-Yum Lee. Multidimensional scaling models with inequality and equality constraints. Communications in Statistics: Simulation and Computation, 13(1):127–140, 1984. CODEN CSSCDB. ISSN 0361-0918.

Lee:1986:EIP

[Lee86] Che-Ping Lee. Estimation of the intensity of a Poisson process. Communications in Statistics: Simulation and Computation, 15 (3):747–759, 1986. CODEN CSSCDB. ISSN 0361-0918.

Lee:1988:CCP

[Lee88] Carl M.-S. Lee. On the computation of F -cumulative probabilities. Communications in Statistics: Simulation and Computation, 17 (4):1191–1201, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lee:1989:MTI

[Lee89] Mei-Ling Ting Lee. A moment test to identify uniformity. Communications in Statistics: Simulation and Computation, 18(1):253– 261, 1989. CODEN CSSCDB. ISSN 0361-0918.

Lemmer:1987:TMC

[Lem87] Hermanus H. Lemmer. A test for the median, combining the sign and signed-rank tests. Communications in Statistics: Simulation and Computation, 16(3):621–627, 1987. CODEN CSSCDB. ISSN 0361-0918.

Levine:1981:SDK

[Lev81] David M. Levine. Sampling distributions of Kruskal’s measure of stress for an asymmetric matrix. Communications in Statistics: Simulation and Computation, 10(4):395–404, 1981. CODEN CSSCDB. ISSN 0361-0918.

Levin:1987:CLA

[Lev87] Bruce Levin. Conditional likelihood analysis in stratum-matched retrospective studies with polytomous disease states. Communications in Statistics: Simulation and Computation, 16(3):699–718, 1987. CODEN CSSCDB. ISSN 0361-0918.

Lachenbruch:1989:CES

[LF89] Peter A. Lachenbruch and Virginia F. Flack. Considerations for effective simulation study analysis. Communications in Statistics: Simulation and Computation, 18(4):1223–1240, 1989. CODEN CSSCDB. ISSN 0361-0918. See comment [HA89a].

Lam:1985:BBD

[LfH85] H. K. Lam and Yat fan Ho. On the bounded binomial distribution and its parameter estimation. Communications in Statistics: Simulation and Computation, 14(1):43–53, 1985. CODEN CSSCDB. ISSN 0361-0918.

Leiva:1986:CIV

[LG86] Ricardo A. Leiva and Franklin A. Graybill. Confidence intervals for variance components in the balanced two way model with interaction. Communications in Statistics: Simulation and Computation, 15(2):301–322, 1986. CODEN CSSCDB. ISSN 0361-0918.

Li:1988:AEL

[LH88] W. K. Li and Y. V. Hui. An algorithm for the exact likelihood of periodic autoregressive moving average models. Communications in Statistics: Simulation and Computation, 17(4):1483–1494, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lemeshow:1980:ENN

[LHH80] Stanley Lemeshow, David W. Hosmer, and David Hislop. The effect of non-normality on estimating the variance of the combined ratio estimate in complex surveys. Communications in Statistics: Simulation and Computation, 9(4):371–387, 1980. CODEN CSS-CDB. ISSN 0361-0918.

Lemeshow:1981:CSS

[LHS81] Stanley Lemeshow, David W. Hosmer, and James P. Stewart. A comparison of sample size determination methods in the two group trial where the underlying disease is rare. Communications in Statistics: Simulation and Computation, 10(5):437–449, 1981. CODEN CSSCDB. ISSN 0361-0918.

Lindstrom:1981:AEI

[Lin81] F. Tom Lindstrom. An algorithm for the evaluation of the incomplete gamma function and the first two partial derivatives with respect to the parameter. Communications in Statistics: Simulation and Computation, 10(5):465–478, 1981. CODEN CSSCDB. ISSN 0361-0918.

Lin:1988:UMP

[Lin88] Paul K. H. Lin. The uniformly most powerful unbiased tests involving the F -distribution. Communications in Statistics: Simulation and Computation, 17(4):1339–1358, 1988. CODEN CSSCDB. ISSN 0361-0918.

Liski:1982:TMS

[Lis82] Erkki P. Liski. A test of the mean square error criterion for shrinkage estimators. Communications in Statistics: Simulation and Computation, 11(5):543–562, 1982. CODEN CSSCDB. ISSN 0361-0918.

Liski:1985:EID

[Lis85] Erkki P. Liski. Estimation from incomplete data in growth curves models. Communications in Statistics: Simulation and Computation, 14(1):13–27, 1985. CODEN CSSCDB. ISSN 0361-0918.

Liu:1984:ERT

[Liu84] Lon-Mu Liu. Estimation of rational transfer function models. Communications in Statistics: Simulation and Computation, 13 (6):775–784, 1984. CODEN CSSCDB. ISSN 0361-0918.

Ljung:1989:NEM

[Lju89] Greta M. Ljung. A note on the estimation of missing values in time series. Communications in Statistics: Simulation and Computation, 18(2):459–465, 1989. CODEN CSSCDB. ISSN 0361-0918.

Loukas:1986:CGB

[LK86] S. Loukas and C. D. Kemp. The computer generation of bivariate binomial and negative binomial random variables. Communications in Statistics: Simulation and Computation, 15(1):15–25, 1986. CODEN CSSCDB. ISSN 0361-0918.

Lee:1982:SPQ

[LKH82] K. R. Lee, C. H. Kapadia, and Donald Hutcherson. Statistical properties of quasi-range in small samples from a gamma density. Communications in Statistics: Simulation and Computation, 11 (2):175–195, 1982. CODEN CSSCDB. ISSN 0361-0918.

Law:1981:RWS

[LKK81] Averill M. Law, W. David Kelton, and Lloyd W. Koenig. Relative width sequential confidence intervals for the mean. Communications in Statistics: Simulation and Computation, 10(1):29–39, 1981. CODEN CSSCDB. ISSN 0361-0918.

Longley:1981:MGS

[Lon81a] James W. Longley. Modified Gram–Schmidt process vs. classical Gram–Schmidt. Communications in Statistics: Simulation and Computation, 10(5):517–527, 1981. CODEN CSSCDB. ISSN 0361-

Longley:1981:LSC

[Lon81b] James Wildon Longley. Least squares computations and the condition of the matrix. Communications in Statistics: Simulation and Computation, 10(6):593–615, 1981. CODEN CSSCDB. ISSN 0361-0918.

Lin:1983:TCS

[LRL^{+}83] Paul K. H. Lin, Dale O. Richards, David R. Long, Matthew D. Myers, and Joyce A. Taylor. Tables for computing shortest confidence intervals involving the F -distribution. Communications in Statistics: Simulation and Computation, 12(6):711–725, 1983. CODEN CSSCDB. ISSN 0361-0918.

Lewis:1989:VRU

[LRW89a] Peter A. W. Lewis, Richard L. Ressler, and R. Kevin Wood. Variance reduction using nonlinear controls and transformations. Communications in Statistics: Simulation and Computation, 18 (2):655–672, 1989. CODEN CSSCDB. ISSN 0361-0918.

Lucas:1989:ALP

[LRW89b] Larry A. Lucas, Tim Robertson, and F. T. Wright. Approximating the level probabilities for a unimodal ordering. Communications in Statistics: Simulation and Computation, 18(4):1401–1420, 1989. CODEN CSSCDB. ISSN 0361-0918.

Lawrence:1981:CLS

[LS81] Kenneth D. Lawrence and Douglas R. Shier. A comparison of least squares and least absolute deviation regression models for estimating Weibull parameters. Communications in Statistics: Simulation and Computation, 10(3):315–326, 1981. CODEN CSSCDB. ISSN 0361-0918.

Lin:1983:RLR

[LS83] Lawrence I-Kuei Lin and Roy L. Sanford. The robustness of the likelihood ratio test, the nonparametric sum rank test, and F -ratio tests when the populations are from the negative binomial family. Communications in Statistics: Simulation and Computation, 12 (5):523–539, 1983. CODEN CSSCDB. ISSN 0361-0918.

Lemeshow:1984:CAV

[LS84] Stanley Lemeshow and Anne M. Stoddard. A comparison of alternative variance estimation strategies for estimating the slope of a linear regression in sample surveys. Communications in Statistics: Simulation and Computation, 13(2):153–168, 1984. CODEN CSSCDB. ISSN 0361-0918.

Lauterbach:1988:AME

[LS88a] Jörg Lauterbach and Peter Stahlecker. Approximate minimax estimation in linear regression: a simulation study. Communications in Statistics: Simulation and Computation, 17(1):209–227, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lee:1988:REL

[LS88b] A. H. Lee and M. J. Silvapulle. Ridge estimation in logistic regression. Communications in Statistics: Simulation and Computation, 17(4):1231–1257, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lee:1988:CP

[LS88c] Carl M.-S. Lee and Karan P. Singh. On the t cumulative probabilities. Communications in Statistics: Simulation and Computation, 17(1):129–135, 1988. CODEN CSSCDB. ISSN 0361-0918.

Leemis:1989:EPE

[LS89a] Lawrence M. Leemis and Li-Hsing Shih. Exponential parameter estimation for data sets containing left and right censored observations. Communications in Statistics: Simulation and Computation, 18(3):1077–1085, 1989. CODEN CSSCDB. ISSN 0361-0918.

Lerman:1989:DCC

[LS89b] Zvi Lerman and Edna Schechtman. Detecting a change in the correlation coefficient in a sequence of bivariate normal variables. Communications in Statistics: Simulation and Computation, 18 (2):589–599, 1989. CODEN CSSCDB. ISSN 0361-0918.

Limam:1988:STI

[LT88] Mohamed H. T. Limam and David R. Thomas. Simultaneous tolerance intervals in the random one-way model with covariates. Communications in Statistics: Simulation and Computation, 17 (3):1007–1019, 1988. CODEN CSSCDB. ISSN 0361-0918.

Lyons:1980:MCE

[Lyo80] N. I. Lyons. M29. Closed expressions for noncentral hypergeometric probabilities. Communications in Statistics: Simulation and Computation, 9(3):313–314, 1980. CODEN CSSCDB. ISSN 0361-0918.

Mack:1981:QED

[Mac81] Gregory A. Mack. A quick and easy distribution-free test for main effects in a two-factor ANOVA. Communications in Statistics: Simulation and Computation, 10(6):571–591, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Magel:1988:TEM

[Mag88] Rhonda Magel. A test for the equality of k medians against the simple tree alternative under right censorship. Communications in Statistics: Simulation and Computation, 17(3):917–925, 1988. CODEN CSSCDB. ISSN 0361-0918.

Money:1982:LRM

[MAGHB82] A. H. Money, J. F. Affleck-Graves, M. L. Hart, and G. D. I. Barr. The linear regression model: L_{p} norm estimation and the choice of p. Communications in Statistics: Simulation and Computation, 11(1):89–109, 1982. CODEN CSSCDB. ISSN 0361-0918.

Marasinghe:1986:NMC

[Mar86] Mervyn G. Marasinghe. A note on methods for computing Tukey’s and Mandel’s interaction sums of squares. Communications in Statistics: Simulation and Computation, 15(3):649–654, 1986. CODEN CSSCDB. ISSN 0361-0918.

Markowski:1987:CTJ

[Mar87a] Edward P. Markowski. Comparing tests judged asymptotically equally efficient. Communications in Statistics: Simulation and Computation, 16(3):629–643, 1987. CODEN CSSCDB. ISSN 0361-0918.

Maryak:1987:APP

[Mar87b] John L. Maryak. On the asymptotic properties of parameter estimates in a regression model with non-normally distributed errors. Communications in Statistics: Simulation and Computation, 16 (4):1117–1121, 1987. CODEN CSSCDB. ISSN 0361-0918.

Matloff:1982:JSR

[Mat82] Norman S. Matloff. James–Stein regression estimation in a prediction context. Communications in Statistics: Simulation and Computation, 11(5):589–601, 1982. CODEN CSSCDB. ISSN 0361-

Mazumdar:1983:ERT

[Maz83] M. Mazumdar. An estimator of the ratio of two variances. Communications in Statistics: Simulation and Computation, 12(4):399– 409, 1983. CODEN CSSCDB. ISSN 0361-0918.

MacGillivray:1988:MMK

[MB88a] H. L. MacGillivray and Kevin P. Balanda. Mixtures, myths and kurtosis. Communications in Statistics: Simulation and Computation, 17(3):789–802, 1988. CODEN CSSCDB. ISSN 0361-0918.

Mengersen:1988:CBS

[MB88b] Kerrie Mengersen and Eve Bofinger. Confidence bounds and selection of the t best populations. Communications in Statistics: Simulation and Computation, 17(3):927–945, 1988. CODEN CSS-CDB. ISSN 0361-0918.

McAllister:1983:SCJ

[McA83] Paul R. McAllister. A system for computing joint probabilities from Jensen’s bivariate F distribution. Communications in Statistics: Simulation and Computation, 12(2):223–238, 1983. CODEN CSSCDB. ISSN 0361-0918.

McArthur:1987:ESD

[McA87] Richard D. McArthur. An evaluation of sample designs for estimating a locally concentrated pollutant. Communications in Statistics: Simulation and Computation, 16(3):735–759, 1987. CODEN CSSCDB. ISSN 0361-0918.

McCulloch:1986:SCE

[McC86] J. Huston McCulloch. Simple consistent estimators of stable distribution parameters. Communications in Statistics: Simulation and Computation, 15(4):1109–1136, 1986. CODEN CSSCDB. ISSN 0361-0918.

Mclachlan:1982:BVS

[Mcl82] G. J. Mclachlan. On the bias and variance of some proportion estimators. Communications in Statistics: Simulation and Computation, 11(6):715–726, 1982. CODEN CSSCDB. ISSN 0361-0918.

Mcculloch:1985:EBA

[MD85] Charles E. Mcculloch and Aimee Dechter. An empirical Bayes approach to estimating the probability of correct selection. Communications in Statistics: Simulation and Computation, 14(1):173– 186, 1985. CODEN CSSCDB. ISSN 0361-0918.

Mittal:1987:EPD

[MD87] Mukul M. Mittal and Ram C. Dahiya. Estimating the parameters of a doubly truncated normal distribution. Communications in Statistics: Simulation and Computation, 16(1):141–159, 1987. CODEN CSSCDB. ISSN 0361-0918.

Meester:1986:TPS

[ME86] S. G. Meester and D. Eaves. A table of predictive success probabilities for logistic regression. Communications in Statistics: Simulation and Computation, 15(4):1137–1139, 1986. CODEN CSSCDB. ISSN 0361-0918.

McLachlan:1982:UDF

[MG82] G. J. McLachlan and S. Ganesalingam. Updating a discriminant function on the basis of unclassified data. Communications in Statistics: Simulation and Computation, 11(6):753–767, 1982. CODEN CSSCDB. ISSN 0361-0918.

Martin:1984:SSB

[MG84] R. Douglas Martin and Daniel M. Goodfellow. Small sample behavior of robust stochastic approximation and iterated weighted least squares estimates for location. Communications in Statistics: Simulation and Computation, 13(1):1–46, 1984. CODEN CSSCDB. ISSN 0361-0918.

Magel:1987:CDN

[MH87] Rhonda C. Magel and Doris Hertsgaard. A collinearity diagnostic for nonlinear regression. Communications in Statistics: Simulation and Computation, 16(1):85–97, 1987. CODEN CSSCDB. ISSN 0361-0918.

Marasinghe:1982:DMG

[MK82] Mervyn G. Marasinghe and William J. Kennedy, Jr. Direct methods for generating extreme characteristic roots of certain random matrices. Communications in Statistics: Simulation and Computation, 11(5):527–542, 1982. CODEN CSSCDB. ISSN 0361-0918.

Mardia:1983:NDM

[MK83] K. V. Mardia and M. Kanazawa. The null distribution of multivariate kurtosis. Communications in Statistics: Simulation and Computation, 12(5):569–576, 1983. CODEN CSSCDB. ISSN 0361-0918.

McAllister:1981:CPJ

[MLH81] Paul R. McAllister, Ru-Ying Lee, and Burt S. Holland. Computation of probabilities from Jensen’s bivariate f distribution. Communications in Statistics: Simulation and Computation, 10 (3):249–263, 1981. CODEN CSSCDB. ISSN 0361-0918.

Moschopoulos:1983:LRB

[MM83] Panagis G. Moschopoulos and Govind S. Mudholkar. A likelihoodratio-based normal approximation for the non-null distribution of the multiple correlation coefficient. Communications in Statistics: Simulation and Computation, 12(3):355–371, 1983. CODEN CSS-CDB. ISSN 0361-0918.

Mandel:1988:ISW

[MM88] John Mandel and Frank L. McCrackin. An iterative self-weighting procedure for fitting straight lines to heteroscedastic data. Communications in Statistics: Simulation and Computation, 17(2): 609–635, 1988. CODEN CSSCDB. ISSN 0361-0918.

McCune:1989:MEC

[MM89] S. K. McCune and E. D. McCune. Modified Edgeworth and Cornish-Fisher expansions with unknown cumulants and no singularities. Communications in Statistics: Simulation and Computation, 18(3):1203–1221, 1989. CODEN CSSCDB. ISSN 0361-0918.

Michalek:1985:MCP

[MMW85] Joel E. Michalek, Daniel Mihalko, and Thomas J. White. A Monte Carlo power study of logrank, Wilcoxon and normal scores procedures on matched and censored data. Communications in Statistics: Simulation and Computation, 14(2):449–465, 1985. CODEN CSSCDB. ISSN 0361-0918.

Monahan:1984:IATa

[Mon84a] John F. Monahan. An improved algorithm for the triples test. Communications in Statistics: Simulation and Computation, 13 (4):545–553, 1984. CODEN CSSCDB. ISSN 0361-0918.

Monahan:1984:IATb

[Mon84b] John F. Monahan. An improved algorithm for the triples test. Communications in Statistics: Simulation and Computation, 13 (4):555–569, 1984. CODEN CSSCDB. ISSN 0361-0918.

Moore:1981:DAF

[Moo81] Marva Houston Moore. Derivatives of the annuity function assuming Makeham–Gompertz mortality; and the n-ages method. Communications in Statistics: Simulation and Computation, 10 (4):359–367, 1981. CODEN CSSCDB. ISSN 0361-0918.

Mehta:1980:NAE

[MP80] Cyrus R. Mehta and Nitin R. Patel. A network algorithm for the exact treatment of the 2 × k contingency table. Communications in Statistics: Simulation and Computation, 9(6):649–664, 1980. CODEN CSSCDB. ISSN 0361-0918.

McConalogue:1981:NTC

[MP81] Denis J. McConalogue and Antònio Pacheco. Numerical treatment of convolution integrals involving distributions with densities having singularities at the origin. Communications in Statistics: Simulation and Computation, 10(3):265–280, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Manly:1989:CPL

[MR89] Bryan F. J. Manly and J. C. W. Rayner. Comments on a paper by lachenbruch and flack. Communications in Statistics: Simulation and Computation, 18(4):1241–1244, 1989. CODEN CSSCDB. ISSN 0361-0918.

McKean:1984:CMS

[MS84] Joseph W. McKean and Ronald M. Schrader. A comparison of methods for Studentizing the sample median. Communications in Statistics: Simulation and Computation, 13(6):751–773, 1984. CODEN CSSCDB. ISSN 0361-0918.

Mian:1989:CST

[MST89] I. U. H. Mian, M. M. Shoukri, and D. S. Tracy. A comparison of significance testing procedures for the intraclass correlation from family data. Communications in Statistics: Simulation and Computation, 18(2):613–631, 1989. CODEN CSSCDB. ISSN 0361-0918.

Manoukian:1985:SHN

[MT85] E. B. Manoukian and J. A. M. Theriault. Sampling from halfnormal and exponential distributions and efficiency of tests. Communications in Statistics: Simulation and Computation, 14(1): 137–141, 1985. CODEN CSSCDB. ISSN 0361-0918.

Moskowitz:1989:EBP

[MT89] Herbert Moskowitz and Hsien-Tsang Tsai. An error-bounded polynomial approximation for bivariate normal probabilities. Communications in Statistics: Simulation and Computation, 18(4):1421– 1437, 1989. CODEN CSSCDB. ISSN 0361-0918.

Meilijson:1982:NMM

[MTNY82] Isaac Meilijson, Aaron Tenenbein, Marian R. Newborn, and Uri Yechiali. Number of matches and matched people in the birthday problem. Communications in Statistics: Simulation and Computation, 11(3):361–370, 1982. CODEN CSSCDB. ISSN 0361-0918.

Marco:1987:EDC

[MYT87] Virgil R. Marco, Dean M. Young, and Danny W. Turner. The Euclidean distance classifier: an alternative to the linear discriminant function. Communications in Statistics: Simulation and Computation, 16(2):485–505, 1987. CODEN CSSCDB. ISSN 0361-0918.

Nath:1981:RTT

[ND81] R. Nath and B. S. Duran. The rank transform in the two-sample location problem. Communications in Statistics: Simulation and Computation, 10(4):383–394, 1981. CODEN CSSCDB. ISSN 0361-0918.

Nelson:1981:NEE

[Nel81] Peter R. Nelson. Numerical evaluation of an equicorrelated multivariate non-central t distribution. Communications in Statistics: Simulation and Computation, 10(1):41–50, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Nelson:1982:MND

[Nel82] Peter R. Nelson. Multivariate normal and t distributions with ρ_{jk} = α_{jk}α_{jk}. Communications in Statistics: Simulation and Computation, 11(2):239–248, 1982. CODEN CSSCDB. ISSN 0361-0918.

Nelsen:1987:DBD

[Nel87a] Roger B. Nelsen. Discrete bivariate distributions with given marginals and correlation. Communications in Statistics: Simulation and Computation, 16(1):199–208, 1987. CODEN CSSCDB. ISSN 0361-0918.

Nelson:1987:PVR

[Nel87b] Barry L. Nelson. A perspective on variance reduction in dynamic simulation experiments. Communications in Statistics: Simulation and Computation, 16(2):385–426, 1987. CODEN CSSCDB. ISSN 0361-0918.

Neuts:1982:CVM

[Neu82] Marcel F. Neuts. On the coefficient of variation of mixtures of probability distributions. Communications in Statistics: Simulation and Computation, 11(6):649–657, 1982. CODEN CSSCDB. ISSN 0361-0918.

Neuts:1984:ACL

[Neu84] Marcel F. Neuts. The abscissa of convergence of the Laplace– Stieltjes transform of a PH-distribution. Communications in Statistics: Simulation and Computation, 13(3):367–373, 1984. CODEN CSSCDB. ISSN 0361-0918.

Neudecker:1986:GRT

[Neu86] H. Neudecker. A generalization of a result of Trenkler and Trenkler. Communications in Statistics: Simulation and Computation, 15(1):273–275, 1986. CODEN CSSCDB. ISSN 0361-0918.

Nguyen:1985:GFE

[Ngu85] True T. Nguyen. A generalization of Fisher’s exact test in p × q contingency tables using more concordant relations. Communications in Statistics: Simulation and Computation, 14(3):633–645, 1985. CODEN CSSCDB. ISSN 0361-0918. See comments [KV87].

Nigm:1985:BLT

[NI85] A. M. Nigm and M. A. Ismail. Bayesian life test sampling plans for the two parameter exponential distribution. Communications in Statistics: Simulation and Computation, 14(3):691–707, 1985. CODEN CSSCDB. ISSN 0361-0918.

Neuhaus:1981:RCS

[NK81] Georg Neuhaus and Erhard Kremer. Repeated chi-square testing repeated chi-square testing. Communications in Statistics: Simulation and Computation, 10(2):143–161, 1981. CODEN CSSCDB. ISSN 0361-0918.

Niki:1984:HOA

[NK84] Naoto Niki and Sadanori Konishi. Higher order asymptotic expansions for the distribution of the sample correlation coefficient.

Communications in Statistics: Simulation and Computation, 13 (2):169–182, 1984. CODEN CSSCDB. ISSN 0361-0918.

Naes:1985:CPM

[NM85] Tormod Naes and Harald Martens. Comparison of prediction methods for multicollinear data. Communications in Statistics: Simulation and Computation, 14(3):545–576, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Nelson:1988:RMA

[NM88] Norma A. Nelson and Patricia F. Moodie. Robust multivariate analysis of variability. Communications in Statistics: Simulation and Computation, 17(4):1409–1430, 1988. CODEN CSSCDB. ISSN 0361-0918.

Naik-Nimbalkar:1986:CRT

[NNS86] U. V. Naik-Nimbalkar and A. Subramanyam. Cramer-rao type bounds for abstract parameters with consequences to real parameters. Communications in Statistics: Simulation and Computation, 15(4):973–985, 1986. CODEN CSSCDB. ISSN 0361-0918.

Nomura:1988:ESS

[Nom88a] Masuo Nomura. On exact small sample properties of the minimax generalized ridge regression estimators. Communications in Statistics: Simulation and Computation, 17(4):1213–1229, 1988. CODEN CSSCDB. ISSN 0361-0918.

Nomura:1988:AUR

[Nom88b] Masuo Nomura. On the almost unbiased ridge regression estimator. Communications in Statistics: Simulation and Computation, 17(3):729–743, 1988. CODEN CSSCDB. ISSN 0361-0918.

Nordberg:1982:PDG

[Nor82] Lennart Nordberg. A procedure for determination of a good ridge parameter in linear regression. Communications in Statistics: Simulation and Computation, 11(3):285–309, 1982. CODEN CSSCDB. ISSN 0361-0918.

Norton:1985:RBF

[Nor85] Robert M. Norton. A relationship between a family of waiting time distributions and the asymptotic residual waiting time distribution. Communications in Statistics: Simulation and Computation, 14(3):709–717, 1985. CODEN CSSCDB. ISSN 0361-0918.

Nozari:1984:GOL

[Noz84] Ardavan Nozari. Generalized and ordinary least squares with estimated and unequal variances. Communications in Statistics: Simulation and Computation, 13(4):521–537, 1984. CODEN CSS-CDB. ISSN 0361-0918.

Nakano:1982:VSA

[NT82] Junji Nakano and Shigemi Tagami. On the variance of the sample autocovariance function for a Gaussian once integrated first order moving average process. Communications in Statistics: Simulation and Computation, 11(5):637–639, 1982. CODEN CSSCDB. ISSN 0361-0918.

Narula:1986:CPN

[NW86] Subhash C. Narula and H. R. Weistroffer. Computation of probability and non-centrality parameter of a non-central F - distribution. Communications in Statistics: Simulation and Computation, 15(3):871–878, 1986. CODEN CSSCDB. ISSN 0361-0918.

Okabe:1986:PDF

[OA86] Atsuyuki Okabe and Yasushi Asami. The probability density function of an area covered with disks whose centers are randomly distributed. Communications in Statistics: Simulation and Computation, 15(1):121–134, 1986. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1982:CVS

[Ode82a] Robert E. Odeh. Critical values of the sample product-moment correlation coefficient in the bivariate normal distribution. Communications in Statistics: Simulation and Computation, 11(1):1– 26, 1982. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1982:TPP

[Ode82b] Robert E. Odeh. Tables of percentage points of the distribution of the maximum absolute value of equally correlated normal random variables. Communications in Statistics: Simulation and Computation, 11(1):65–87, 1982. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1989:SOS

[Ode89a] Robert E. Odeh. Simultaneous one-sided prediction intervals to contain all of k future means from a normal distribution. Communications in Statistics: Simulation and Computation, 18(4):1557– 1585, 1989. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1989:STS

[Ode89b] Robert E. Odeh. Simultaneous two-sided prediction intervals to contain at least l out of k future means from a normal distribution. Communications in Statistics: Simulation and Computation, 18 (2):429–457, 1989. CODEN CSSCDB. ISSN 0361-0918.

Ott:1988:EBI

[OH88] K. O. Ott and H.-J Hoffmann. Evaluation of the bias of the isotonic regression. Communications in Statistics: Simulation and Computation, 17(3):745–764, 1988. CODEN CSSCDB. ISSN 0361-0918.

Ohtani:1987:UGD

[Oht87] Kazuhiro Ohtani. On the use of the Graybill–Deal estimator when two normal population means may not be common. Communications in Statistics: Simulation and Computation, 16(3):855–870, 1987. CODEN CSSCDB. ISSN 0361-0918.

Ozturk:1988:NTE

[ÖK88] Aydin Öztürk and Serdar Korukoğlu. A new test for the extreme value distribution. Communications in Statistics: Simulation and Computation, 17(4):1375–1393, 1988. CODEN CSSCDB. ISSN 0361-0918.

Owen:1988:SPE

[OL88] D. B. Owen and Huaixiang Li. The starship for point estimates and confidence intervals on a mean and for percentiles. Communications in Statistics: Simulation and Computation, 17(2):325–341, 1988. CODEN CSSCDB. ISSN 0361-0918.

Oldford:1985:BMC

[Old85] R. Wayne Oldford. Bootstrapping by Monte Carlo versus approximating the estimator and bootstrapping exactly: Cost and performance. Communications in Statistics: Simulation and Computation, 14(2):395–424, 1985. CODEN CSSCDB. ISSN 0361-0918.

OGorman:1987:MEO

[OM87] Mary Ann O’Gorman and Raymond H. Myers. Measures of error with outliers in regression. Communications in Statistics: Simulation and Computation, 16(3):771–789, 1987. CODEN CSSCDB. ISSN 0361-0918.

Oman:1980:EPW

[Oma80] Samuel D. Oman. Estimating a proportion when sampling from a subpopulation. Communications in Statistics: Simulation and Computation, 9(1):81–101, 1980. CODEN CSSCDB. ISSN 0361-0918.

Odhiambo:1985:TSG

[OP85] J. W. Odhiambo and M. S. Patel. Three-stage group screening with errors in observations. Communications in Statistics: Simulation and Computation, 14(3):647–666, 1985. CODEN CSSCDB. ISSN 0361-0918.

Osborne:1989:LMT

[Osb89] M. R. Osborne. Lad for multiway tables. Communications in Statistics: Simulation and Computation, 18(2):829–834, 1989. CODEN CSSCDB. ISSN 0361-0918.

Ohtani:1985:MCS

[OT85] Kazuhiro Ohtani and Toshihisa Toyoda. A Monte Carlo study of the Wald, LM, and LR tests in a heteroscedastic linear model. Communications in Statistics: Simulation and Computation, 14 (3):735–746, 1985. CODEN CSSCDB. ISSN 0361-0918.

Oprian:1980:GCI

[OTVA80] C. Oprian, V. Taneja, D. Voss, and L. A. Aroian. General considerations and interrelationships between MA and AR models, time series in m dimensions, the ARMA model. Communications in Statistics: Simulation and Computation, 9(5):515–532, 1980. CODEN CSSCDB. ISSN 0361-0918.

Owen:1980:TNI

[Owe80] D. B. Owen. A table of normal integrals. Communications in Statistics: Simulation and Computation, 9(4):389–419, 1980. CO-DEN CSSCDB. ISSN 0361-0918.

Owen:1985:E

[Owe85] D. B. Owen. Editorial. Communications in Statistics: Simulation and Computation, 14(1):iii–v, 1985. CODEN CSSCDB. ISSN 0361-0918.

Owen:1988:S

[Owe88] D. B. Owen. The starship. Communications in Statistics: Simulation and Computation, 17(2):315–323, 1988. CODEN CSSCDB. ISSN 0361-0918.

OGorman:1988:MCS

[OWJL88] Thomas W. O’Gorman, Robert F. Woolson, Michael P. Jones, and Jon H. Lemke. A Monte Carlo study of three odds ratio estimators and four tests of association in several 2x2 tables when the data are sparse. Communications in Statistics: Simulation and Computation, 17(3):813–835, 1988. CODEN CSSCDB. ISSN 0361-0918.

Odeh:1987:PCS

[OYO87] Robert E. Odeh, Min Younchou, and D. B. Owen. The precision for coverages and sample size requirements for normal tolerance intervals. Communications in Statistics: Simulation and Computation, 16(4):969–985, 1987. CODEN CSSCDB. ISSN 0361-0918.

Paik:1988:RMA

[Pai88] Myunghee C. Paik. Repeated measurement analysis for nonnormal data in small samples. Communications in Statistics: Simulation and Computation, 17(4):1155–1171, 1988. CODEN CSSCDB. ISSN 0361-0918.

Palta:1983:ESA

[Pal83] Mari Palta. The effects of a stratified analysis of censored data. Communications in Statistics: Simulation and Computation, 12 (3):273–290, 1983. CODEN CSSCDB. ISSN 0361-0918.

Panton:1988:PPS

[Pan88] Don B. Panton. A PASCAL program for simulating stable random variates. Communications in Statistics: Simulation and Computation, 17(3):837–842, 1988. CODEN CSSCDB. ISSN 0361-0918.

Parson:1981:SSD

[Par81] Van L. Parson. Small sample distribution for a nonparametric test for trend. Communications in Statistics: Simulation and Computation, 10(3):289–302, 1981. CODEN CSSCDB. ISSN 0361-0918.

Parsian:1988:AEZ

[Par88] Ahmad Parsian. On the admissibility of estimators of the zero class in Poisson populations. Communications in Statistics: Simulation and Computation, 17(1):177–186, 1988. CODEN CSSCDB. ISSN 0361-0918.

Patel:1989:ENI

[Pat89] Kartik R. Patel. An extension of the normal/independent family for studies of robust estimators. Communications in Statistics: Simulation and Computation, 18(4):1587–1602, 1989. CODEN CSSCDB. ISSN 0361-0918.

Pfeifer:1980:CEP

[PD80a] Phillip E. Pfeifer and Stuart Jay Deutsch. A comparison of estimation procedures for the parameters of the star model. Communications in Statistics: Simulation and Computation, 9(3):255–270, 1980. CODEN CSSCDB. ISSN 0361-0918.

Pfeifer:1980:IST

[PD80b] Phillip E. Pfeifer and Stuart Jay Deutsch. Independence and sphericity tests for the residuals of space-time ARMA models. Communications in Statistics: Simulation and Computation, 9(5): 533–549, 1980. CODEN CSSCDB. ISSN 0361-0918.

Pfeifer:1980:SIR

[PD80c] Phillip E. Pfeifer and Stuart Jay Deutsch. Stationarity and invertibility regions for low order STARMA models. Communications in Statistics: Simulation and Computation, 9(5):551–562, 1980. CODEN CSSCDB. ISSN 0361-0918.

Paulson:1984:SPM

[PD84] A. S. Paulson and T. A. Delehanty. Some properties of modified integrated squared error estimators for the stable laws. Communications in Statistics: Simulation and Computation, 13(3):337–365, 1984. CODEN CSSCDB. ISSN 0361-0918.

Paulson:1985:MWS

[PD85] A. S. Paulson and T. A. Delehanty. Modified weighted squared error estimation procedures with special emphasis on the stable laws. Communications in Statistics: Simulation and Computation, 14(4):927–972, 1985. CODEN CSSCDB. ISSN 0361-0918.

Peddada:1989:TNN

[Ped89] Shyamal Das Peddada. Two non-negative estimators for the model with a common mean. Communications in Statistics: Simulation and Computation, 18(2):501–512, 1989. CODEN CSSCDB. ISSN 0361-0918.

Pepple:1989:AME

[Pep89] Patricia A. Pepple. Approximation methods for estimating gamma means using a hierarchical exchangeable model. Communications in Statistics: Simulation and Computation, 18(1):83–98, 1989. CODEN CSSCDB. ISSN 0361-0918.

Pantula:1986:CAF

[PF86a] Sastry G. Pantula and Wayne A. Fuller. Computational algorithms for the factor model. Communications in Statistics: Simulation and Computation, 15(1):227–259, 1986. CODEN CSSCDB. ISSN 0361-0918.

Paul:1986:CVD

[PF86b] S. R. Paul and Karen Yuen Fung. Critical values for Dixon type test statistics for testing outliers in Weibull or extreme value distributions. Communications in Statistics: Simulation and Computation, 15(1):277–283, 1986. CODEN CSSCDB. ISSN 0361-0918.

Pope:1988:FND

[PG88] S. B. Pope and R. Gadh. Fitting noisy data using cross-validated cubic smoothing splines. Communications in Statistics: Simulation and Computation, 17(2):349–376, 1988. CODEN CSSCDB. ISSN 0361-0918.

Panchang:1989:DTP

[PG89] Vijay G. Panchang and Ramesh C. Gupta. On the determination of three-parameter Weibull MLE’s. Communications in Statistics: Simulation and Computation, 18(3):1037–1057, 1989. CODEN CSSCDB. ISSN 0361-0918.

Phillips:1984:JTS

[PM84] G. D. A. Phillips and B. P. McCabe. Jackknifing the two stage least squares estimator. Communications in Statistics: Simulation and Computation, 13(2):141–152, 1984. CODEN CSSCDB. ISSN 0361-0918.

Patel:1987:SWG

[PM87] M. S. Patel and M. M. Manene. Step-wise group screening with equal prior probabilities and no errors in observations. Communications in Statistics: Simulation and Computation, 16(3):817–833, 1987. CODEN CSSCDB. ISSN 0361-0918.

Pandey:1989:EMR

[PM89] B. N. Pandey and H. J. Malik. Estimation of the mean and the reciprocal of the mean of the inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 18(3):1187– 1201, 1989. CODEN CSSCDB. ISSN 0361-0918.

Peres:1989:ACF

[PN89] Clovis A. Peres and Subhash C. Narula. An approximate closed form solution for a quadratic dose-response model. Communications in Statistics: Simulation and Computation, 18(2):643–653, 1989. CODEN CSSCDB. ISSN 0361-0918.

Penm:1984:MSA

[PT84] J. H. W. Penm and R. D. Terrell. Multivariate subset autoregression. Communications in Statistics: Simulation and Computation, 13(4):449–461, 1984. CODEN CSSCDB. ISSN 0361-0918.

Padgett:1986:SNQ

[PT86] W. J. Padgett and L. A. Thombs. Smooth nonparametric quantile estimation under censoring: simulations and bootstrap methods. Communications in Statistics: Simulation and Computation, 15 (4):1003–1025, 1986. CODEN CSSCDB. ISSN 0361-0918.

Pukkila:1984:DDB

[Puk84] Tarmo Pukkila. On the distributions of the differences between the estimated autocorrelations and partial autocorrelations at the same lag. Communications in Statistics: Simulation and Computation, 13(4):463–488, 1984. CODEN CSSCDB. ISSN 0361-0918.

Quesenberry:1982:CPF

[Que82] C. P. Quesenberry. Clarification of priority of Fisher and Pearson and further remarks on combined tests of significance. Communications in Statistics: Simulation and Computation, 11(2):229–231, 1982. CODEN CSSCDB. ISSN 0361-0918.

Quinn:1980:MLB

[Qui80] B. G. Quinn. M28. Limiting behaviour of autocorrelation function of ARMA process as several roots of characteristic equation approach unit circle. Communications in Statistics: Simulation and Computation, 9(2):195–198, 1980. CODEN CSSCDB. ISSN 0361-0918.

Qwen:1981:EE

[Qwe81] D. B. Qwen. Editorial editorial. Communications in Statistics: Simulation and Computation, 10(1):iii–iv, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Razzaghi:1987:SPM

[Raz87] Mehdi Razzaghi. A superiority problem in misspecified restricted linear models. Communications in Statistics: Simulation and Computation, 16(3):899–902, 1987. CODEN CSSCDB. ISSN 0361-0918.

Rodrigues:1988:BAC

[RBL88] Josemar Rodrigues, Heleno Bolfarine, and José Galvão Leite. A Bayesian analysis in closed animal populations from capture recapture experiments with trap response. Communications in Statistics: Simulation and Computation, 17(2):407–430, 1988. CODEN CSSCDB. ISSN 0361-0918.

Rocke:1981:EVR

[RD81] David M. Rocke and George W. Downs. Estimating the variances of robust estimators of location: influence curve, jackknife and bootstrap robust estimators of location. Communications in Statistics: Simulation and Computation, 10(3):221–248, 1981. CODEN CSSCDB. ISSN 0361-0918.

Reynolds:1982:CST

[RD82] Marion R. Reynolds, Jr. and M. L. Deaton. Comparisons of some tests for validation of stochastic simulation models. Communications in Statistics: Simulation and Computation, 11(6):769–799, 1982. CODEN CSSCDB. ISSN 0361-0918.

Rao:1985:JPP

[RD85] Poduri S. R. S. Rao and Atsu Shelter Dorvlo. The jackknife procedure for the probabilities of misclassification. Communications in Statistics: Simulation and Computation, 14(4):779–790, 1985. CODEN CSSCDB. ISSN 0361-0918.

Roberson:1983:LRT

[RF83] Paula K. Roberson and Lloyd Fisher. Lack of robustness in timespace disease clustering. Communications in Statistics: Simulation and Computation, 12(1):11–22, 1983. CODEN CSSCDB. ISSN 0361-0918.

Rhiel:1989:IRE

[Rhi89] Steven G. Rhiel. An improved range estimator of sigma for determining sample sizes. Communications in Statistics: Simulation and Computation, 18(4):1295–1309, 1989. CODEN CSSCDB. ISSN 0361-0918.

Raveh:1988:STS

[RM88] Adi Raveh and Gur Mosheiov. Smoothing time-series data by nonmetric polytone curves. Communications in Statistics: Simulation and Computation, 17(2):515–536, 1988. CODEN CSSCDB. ISSN 0361-0918.

Ramig:1980:PIB

[RN80] Pauline F. Ramig and Peter R. Nelson. The probability integral for a bivariate generalization of the non-central t. Communications in Statistics: Simulation and Computation, 9(6):621–631, 1980. CODEN CSSCDB. ISSN 0361-0918.

Rech:1989:LAR

[RSE89] Paul Rech, Paul Schmidbauer, and Jamie Eng. Least absolute regression revisited a simple labeling method for finding a lar line. Communications in Statistics: Simulation and Computation, 18 (3):943–955, 1989. CODEN CSSCDB. ISSN 0361-0918.

Ryan:1988:SSW

[RSG88] G. W. Ryan, W. B. Smith, and C. E. Gates. Simulation study of wildlife density estimators. Communications in Statistics: Simulation and Computation, 17(2):431–450, 1988. CODEN CSSCDB. ISSN 0361-0918.

Rai:1980:SEN

[RSV80] Kamta Rai, V. Susarla, and John Van Ryzin. Shrinkage estimation in nonparametric Bayesian survival analysis: a simulation study. Communications in Statistics: Simulation and Computation, 9(3): 271–298, 1980. CODEN CSSCDB. ISSN 0361-0918.

Rubin:1984:GRP

[Rub84] Paul A. Rubin. Generating random points in a polytope. Communications in Statistics: Simulation and Computation, 13(3):375– 396, 1984. CODEN CSSCDB. ISSN 0361-0918.

Rubin:1985:SRC

[Rub85] Paul A. Rubin. Short-run characteristics of samples drawn by random walks. Communications in Statistics: Simulation and Computation, 14(2):473–490, 1985. CODEN CSSCDB. ISSN 0361-0918.

Sklar:1984:ADC

[SA84] Michael G. Sklar and Ronald D. Armstrong. An algorithm for discrete Chebychev curve fitting for the simple model using a dual linear programming approach. Communications in Statistics: Simulation and Computation, 13(4):555–569, 1984. CODEN CSSCDB. ISSN 0361-0918.

Sambamoorthi:1989:ITC

[Sam89] N. Sambamoorthi. Information theoretic criterion approach to dimensionality reduction in multinomial logistic regression models: Part II: Some simulations. Communications in Statistics: Simulation and Computation, 18(3):897–908, 1989. CODEN CSSCDB. ISSN 0361-0918.

Sanchez:1987:SSP

[San87] Susan M. Sanchez. Small-sample performance of a modified leastfailures sampling procedure for Bernoulli subset selection. Communications in Statistics: Simulation and Computation, 16(4): 1051–1065, 1987. CODEN CSSCDB. ISSN 0361-0918.

Sathe:1982:MAT

[Sat82] Y. S. Sathe. M42. Another test for equality of two proportions. Communications in Statistics: Simulation and Computation, 11 (3):373–375, 1982. CODEN CSSCDB. ISSN 0361-0918.

Saleh:1983:SEP

[SAU83] A. K. Ehsanes Saleh, M. Masoom Ali, and Dale Umbach. Simplified estimation of the parameters of exponential distribution from samples censored in the middle. Communications in Statistics: Simulation and Computation, 12(5):609–627, 1983. CODEN CSSCDB. ISSN 0361-0918.

Shapiro:1987:TWD

[SB87] Samuel S. Shapiro and Carlos W. Brain. W -test for the Weibull distribution. Communications in Statistics: Simulation and Computation, 16(1):209–219, 1987. CODEN CSSCDB. ISSN 0361-0918.

Sanchez:1989:USB

[SC89a] J. M. Prada Sánchez and X. L. Otero Cepeda. The use of smooth bootstrap techniques for estimating the error rate of a prediction rule. Communications in Statistics: Simulation and Computation, 18(3):1169–1186, 1989. CODEN CSSCDB. ISSN 0361-0918.

Shoukri:1989:BAG

[SC89b] M. M. Shoukri and P. C. Consul. Bayesian analysis of a generalized Poisson distribution. Communications in Statistics: Simulation and Computation, 18(4):1465–1480, 1989. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:1989:CBM

[SC89c] M. S. Srivastava and Y. M. Chan. A comparison of bootstrap method and Edgeworth expansion in approximating the distribution of sample variance — one sample and two sample cases. Communications in Statistics: Simulation and Computation, 18 (1):339–361, 1989. CODEN CSSCDB. ISSN 0361-0918.

Scariano:1985:EPI

[Sca85] Stephen M. Scariano. Estimating the point of intersection of two lines under correlation. Communications in Statistics: Simulation and Computation, 14(1):255–261, 1985. CODEN CSSCDB. ISSN 0361-0918.

Schafer:1982:BED

[Sch82a] Ray E. Schafer. Bounds on expected differences of order statistics. Communications in Statistics: Simulation and Computation, 11 (5):643–644, 1982. CODEN CSSCDB. ISSN 0361-0918.

Schwertman:1982:MCS

[Sch82b] Neil C. Schwertman. A Monte Carlo study of the L_{N} statistic for the multivariate nonparametric median and rank sum tests for two populations. Communications in Statistics: Simulation and Computation, 11(6):667–676, 1982. CODEN CSSCDB. ISSN 0361-0918.

Schwertman:1987:APD

[Sch87] Neil C. Schwertman. An alternative procedure for determining analysis of variance sample size. Communications in Statistics: Simulation and Computation, 16(4):957–967, 1987. CODEN CSS-CDB. ISSN 0361-0918.

Schaalje:1986:CED

[SCJ86] G. B. Schaalje, W. A. Charnetski, and D. L. Johnson. A comparison of estimators of the degree of insect control. Communications in Statistics: Simulation and Computation, 15(4):1065–1086, 1986. CODEN CSSCDB. ISSN 0361-0918.

Schwemer:1980:PPE

[SD80] Gregory T. Schwemer and Olive Jean Dunn. Posterior probability estimators in classification simulations. Communications in Statistics: Simulation and Computation, 9(2):133–140, 1980. CODEN CSSCDB. ISSN 0361-0918.

Schall:1986:TRB

[SD86] R. Schall and T. T. Dunne. Transformation and reparametrization for best linear unbiased estimation of a linear model with arbitrary known variance. Communications in Statistics: Simulation and Computation, 15(4):905–916, 1986. CODEN CSSCDB. ISSN 0361-0918.

Strijbosch:1988:CBR

[SD88] Leo W. G. Strijbosch and Ronald J. M. M. Does. Comparison of bias-reducing methods for estimating the parameter in dilution series. Communications in Statistics: Simulation and Computation, 17(4):1173–1190, 1988. CODEN CSSCDB. ISSN 0361-0918.

Salter:1985:RPR

[SF85] K. C. Salter and R. F. Fawcett. A robust and powerful rank test of treatment effects in balanced incomplete block designs. Communications in Statistics: Simulation and Computation, 14(4):807–828, 1985. CODEN CSSCDB. ISSN 0361-0918.

Schwertman:1981:AIG

[SFM81] N. C. Schwertman, D. Fridshal, and J. M. Magrey. On the analysis of incomplete growth curve data, a Monte Carlo study of two nonparametric procedures. Communications in Statistics: Simulation and Computation, 10(1):51–66, 1981. CODEN CSSCDB. ISSN 0361-0918.

Surekha:1984:MCC

[SG84] K. Surekha and W. E. Griffiths. A Monte Carlo comparison of some Bayesian and sampling theory estimators in two heteroscedastic error models. Communications in Statistics: Simulation and Computation, 13(1):85–105, 1984. CODEN CSSCDB. ISSN 0361-0918.

Sechhofer:1989:TBK

[SG89] Robert E. Sechhofer and David M. Goldsman. Truncation of the Bechhofer–Kiefer–Sobel sequential procedure for selecting the normal population which has the largest mean (111): supplementary truncation numbers and resulting performance characteristics. Communications in Statistics: Simulation and Computation, 18(1):63–81, 1989. CODEN CSSCDB. ISSN 0361-0918.

Smith:1981:ERT

[SH81] Wendell C. Smith and Chien-Pai Han. Error rate for testing a contrast after a significant f test. Communications in Statistics: Simulation and Computation, 10(6):545–556, 1981. CODEN CSS-CDB. ISSN 0361-0918.

Searle:1983:FCA

[SH83] S. R. Searle and Harold V. Henderson. Faults in a computing algorithm for reparameterizing linear models. Communications in Statistics: Simulation and Computation, 12(1):67–76, 1983. CO-DEN CSSCDB. ISSN 0361-0918.

Sehwertman:1980:CEP

[SHA80] Neil C. Sehwertman, John R. Huseby, and David M. Allen. Computation of the estimated parameters and Wald statistic for the generalized growth curve model. Communications in Statistics: Simulation and Computation, 9(6):675–693, 1980. CODEN CSS-CDB. ISSN 0361-0918.

Shaarawy:1989:BIF

[Sha89a] Samir Moustafa Shaarawy. Bayesian inferences and forecasts with multiple autoregressive moving average models. Communications in Statistics: Simulation and Computation, 18(4):1481–1509, 1989. CODEN CSSCDB. ISSN 0361-0918.

Shayib:1989:PST

[Sha89b] Mohammed A. Shayib. The procedure for selection of transformations from the Johnson system. Communications in Statistics: Simulation and Computation, 18(4):1457–1464, 1989. CODEN CSSCDB. ISSN 0361-0918.

Schmitz:1983:CPF

[SHHR83] P. I. M. Schmitz, J. D. F. Habbema, J. Hermans, and J. W. Raatgever. Comparative performance of four discriminant analysis methods for mixtures of continuous and discrete variables. Communications in Statistics: Simulation and Computation, 12 (6):727–751, 1983. CODEN CSSCDB. ISSN 0361-0918.

Shoemaker:1986:NMA

[Sho86] Lewis H. Shoemaker. A nonparametric method for analysis of variance. Communications in Statistics: Simulation and Computation, 15(3):609–632, 1986. CODEN CSSCDB. ISSN 0361-0918.

Sposito:1983:EUS

[SHS83] V. A. Sposito, M. L. Hand, and Bradley Skarpness. On the efficiency of using the sample kurtosis in selecting optimal l_{p} estimators. Communications in Statistics: Simulation and Computation, 12(3):265–272, 1983. CODEN CSSCDB. ISSN 0361-0918.

Silvapulle:1985:ECR

[Sil85] M. J. Silvapulle. An examination of criticisms of ridge regression simulation methodology. Communications in Statistics: Simulation and Computation, 14(4):829–835, 1985. CODEN CSSCDB. ISSN 0361-0918.

Simonoff:1984:COD

[Sim84] Jeffrey S. Simonoff. The calculation of outlier detection statistics. Communications in Statistics: Simulation and Computation, 13 (2):275–285, 1984. CODEN CSSCDB. ISSN 0361-0918.

Sim:1986:SWG

[Sim86] Chiaw-Hock Sim. Simulation of Weibull and gamma autoregressive stationary process. Communications in Statistics: Simulation and Computation, 15(4):1141–1146, 1986. CODEN CSSCDB. ISSN 0361-0918.

Singh:1980:EDL

[Sin80] Anita Singh. On the exact distribution of the likelihood ratio criterion for testing. Communications in Statistics: Simulation and Computation, 9(6):611–619, 1980. CODEN CSSCDB. ISSN 0361-0918.

Singh:1982:EDW

[Sin82] Anita Singh. Exact distribution of Wilks’ l_{V c} criterion and its percentage points in the complex case. Communications in Statistics: Simulation and Computation, 11(2):217–225, 1982. CODEN CSSCDB. ISSN 0361-0918.

Stephenson:1988:CNA

[SJ88] W. Robert Stephenson and David Jacobson. A comparison of nonparametric analysis of covariance techniques. Communications in Statistics: Simulation and Computation, 17(2):451–461, 1988. CODEN CSSCDB. ISSN 0361-0918.

Snapinn:1988:BSC

[SK88] Steven M. Snapinn and James D. Knoke. Bootstrapped and smoothed classification error rate estimators. Communications in Statistics: Simulation and Computation, 17(4):1135–1153, 1988. CODEN CSSCDB. ISSN 0361-0918.

Soundararajan:1989:MSI

[SK89] V. Soundararajan and V. Kuralmani. Multiple sampling inspection plans for attributes. Communications in Statistics: Simulation and Computation, 18(4):1251–1274, 1989. CODEN CSSCDB. ISSN 0361-0918.

Skillings:1983:NAT

[Ski83] John H. Skillings. Nonparametric approaches to testing and multiple comparisons in a one-way ANOVA. Communications in Statistics: Simulation and Computation, 12(4):373–387, 1983. CODEN CSSCDB. ISSN 0361-0918.

Shier:1984:CRR

[SL84] Douglas R. Shier and Kenneth D. Lawrence. A comparison of robust regression techniques for the estimation of Weibull parameters. Communications in Statistics: Simulation and Computation, 13(6):743–750, 1984. CODEN CSSCDB. ISSN 0361-0918.

Slud:1983:TSF

[Slu83] Eric V. Slud. Testing separate families of hypotheses using rightcensored data. Communications in Statistics: Simulation and Computation, 12(4):507–509, 1983. CODEN CSSCDB. ISSN 0361-0918.

Schwemer:1980:NLD

[SM80] Gregory T. Schwemer and M. Ray Mickey. A note on the linear discriminant function when group means are equal. Communications in Statistics: Simulation and Computation, 9(6):633–638, 1980. CODEN CSSCDB. ISSN 0361-0918.

Strenzwilk:1989:TDA

[SMF89] Denis F. Strenzwilk, Michael P. Meredith, and Walter T. Federer. A two dimensional ARMA model for the simulation of IR backgrounds. Communications in Statistics: Simulation and Computation, 18(4):1539–1555, 1989. CODEN CSSCDB. ISSN 0361-0918.

Smith:1982:HTB

[Smi82] Patricia L. Smith. Hypothesis testing in B-spline regression. Communications in Statistics: Simulation and Computation, 11(2): 143–157, 1982. CODEN CSSCDB. ISSN 0361-0918.

Smith:1984:BAI

[Smi84] Philip J. Smith. A Bayesian analysis of interrelated Bernoulli processes with an application for clinical trials. Communications in Statistics: Simulation and Computation, 13(1):69–84, 1984. CO-DEN CSSCDB. ISSN 0361-0918.

Shyu:1986:OST

[SO86] Jyh-Cherng Shyu and Donald B. Owen. One-sided tolerance intervals for the two-parameter double exponential distribution. Communications in Statistics: Simulation and Computation, 15(1): 101–119, 1986. CODEN CSSCDB. ISSN 0361-0918.

Shyu:1987:ETI

[SO87] Jyh-Cherng Shyu and D. B. Owen. B-expectation tolerance intervals for the double exponential distribution. Communications in Statistics: Simulation and Computation, 16(1):129–139, 1987. CODEN CSSCDB. ISSN 0361-0918.

Soms:1983:BTA

[Som83] Andrew P. Soms. Bounds for the t-tail area. Communications in Statistics: Simulation and Computation, 12(5):559–568, 1983. CODEN CSSCDB. ISSN 0361-0918.

Soms:1989:ECI

[Som89] Andrew P. Soms. Exact confidence intervals, based on the z statistic, for the difference between two proportions. Communications in Statistics: Simulation and Computation, 18(4):1325–1341, 1989. CODEN CSSCDB. ISSN 0361-0918.

Sparks:1987:SEV

[Spa87] Ross S. Sparks. Selecting estimators and variables in the seemingly unrelated regression model. Communications in Statistics: Simulation and Computation, 16(1):99–127, 1987. CODEN CSSCDB. ISSN 0361-0918.

Shen:1983:RTT

[SQ83a] C. David Shen and Dana Quade. A randomization test for a three-period three-treatment crossover experiment. Communications in Statistics: Simulation and Computation, 12(2):183–199, 1983. CODEN CSSCDB. ISSN 0361-0918.

Silva:1983:EAR

[SQ83b] Claudio Silva and Dana Quade. Estimating the asymptotic relative efficiency of weighted rankings. Communications in Statistics: Simulation and Computation, 12(5):511–521, 1983. CODEN CSS-CDB. ISSN 0361-0918.

Shoemaker:1983:MEM

[SR83] Lewis H. Shoemaker and James L. Rosenberger. Moments and efficiency of the median and trimmed mean for finite populations. Communications in Statistics: Simulation and Computation, 12 (4):411–422, 1983. CODEN CSSCDB. ISSN 0361-0918.

Solomon:1983:NST

[SS83] Herbert Solomon and Michael A. Stephens. On Neyman’s statistic for testing uniformity. Communications in Statistics: Simulation and Computation, 12(2):127–134, 1983. CODEN CSSCDB. ISSN 0361-0918.

Singer:1985:ARE

[SS85a] Julio M. Singer and Pranab K. Sen. Asymptotic relative efficiency of multivariate m-estimators. Communications in Statistics: Simulation and Computation, 14(1):29–41, 1985. CODEN CSSCDB. ISSN 0361-0918.

Singh:1985:RNR

[SS85b] Sukhminder Singh and Ravindra Singh. On random non-response in double sampling with difference estimator. Communications in Statistics: Simulation and Computation, 14(3):747–757, 1985. CODEN CSSCDB. ISSN 0361-0918.

Shafer:1986:SST

[SS86] Nancy J. Shafer and James A. Sullivan. A simulation study of a test for the equality of the coefficients of variation. Communications in Statistics: Simulation and Computation, 15(3):681–695, 1986. CODEN CSSCDB. ISSN 0361-0918.

Schader:1988:SSP

[SS88a] Martin Schader and Friedrich Schmid. Small sample properties of the maximum likelihood estimators of the parameters μ and σ from a grouped sample of a normal population. Communications in Statistics: Simulation and Computation, 17(1):229–239, 1988. CODEN CSSCDB. ISSN 0361-0918.

Smith:1988:BBS

[SS88b] L. A. Smith and R. L. Sielken, Jr. Bootstrap bounds for “safe” doses in the multistage cancer dose-response model. Communications in Statistics: Simulation and Computation, 17(1):153–175, 1988. CODEN CSSCDB. ISSN 0361-0918.

Sugiura:1989:LBI

[SS89a] Nariaki Sugiura and Hiromi Sasamoto. Locally best invariant test for outliers in a gamma type distribution. Communications in Statistics: Simulation and Computation, 18(2):415–427, 1989. CODEN CSSCDB. ISSN 0361-0918.

Swain:1989:CVM

[SS89b] James J. Swain and Bruce W. Schmeiser. Control variates for Monte Carlo analysis of nonlinear statistical models, i: Overview. Communications in Statistics: Simulation and Computation, 18 (3):1011–1036, 1989. CODEN CSSCDB. ISSN 0361-0918.

Starks:1982:MCE

[Sta82] T. H. Starks. A Monte Carlo evaluation of response surface analysis based on paired-comparison data. Communications in Statistics: Simulation and Computation, 11(5):603–617, 1982. CODEN CSSCDB. ISSN 0361-0918.

Schipp:1988:MEA

[STS88] B. Schipp, G. Trenkler, and P. Stahlecker. Minimax estimation with additional linear restrictions — a simulation study. Communications in Statistics: Simulation and Computation, 17(2):393– 406, 1988. CODEN CSSCDB. ISSN 0361-0918.

Sun:1988:FSC

[Sun88] Hong-Jie Sun. A Fortran subroutine for computing normal orthant probabilities of dimensions up to nine. Communications in Statistics: Simulation and Computation, 17(3):1097–1111, 1988. CODEN CSSCDB. ISSN 0361-0918.

Schork:1980:NOR

[SW80] M. Anthony Schork, Ph. D. and George W. Williams, Ph. D. Number of observations required for the comparison of two correlated proportions. Communications in Statistics: Simulation and Computation, 9(4):349–357, 1980. CODEN CSSCDB. ISSN 0361-0918.

Schechtman:1985:MCP

[SW85] Edna Schechtman and Douglas A. Wolfe. Multiple changepoints problem-nonparmetric procedures for estimation of the points of change. Communications in Statistics: Simulation and Computation, 14(3):615–631, 1985. CODEN CSSCDB. ISSN 0361-0918.

Scariano:1988:ACN

[SW88a] Stephen M. Scariano and Terry A. Watkins. An algorithm for computing nonparametric estimators of change-points. Communications in Statistics: Simulation and Computation, 17(4):1511– 1532, 1988. CODEN CSSCDB. ISSN 0361-0918.

Shiau:1988:RCS

[SW88b] Jyh-Jen Horng Shiau and Grace Wahba. Rates of convergence of some estimators for a semiparametric model. Communications in Statistics: Simulation and Computation, 17(4):1117–1133, 1988. CODEN CSSCDB. ISSN 0361-0918.

Swain:1988:CVM

[Swa88] James J. Swain. Control variates for Monte Carlo analysis of nonlinear statistical models. IV. Nonnormal error distributions. Communications in Statistics: Simulation and Computation, 17 (1):251–274, 1988. CODEN CSSCDB. ISSN 0361-0918.

Spurrier:1989:GDM

[SWP89] John D. Spurrier, Josephine Wilson, and James W. Park. A generalized distance multi-sample test of normality with applications to process control. Communications in Statistics: Simulation and Computation, 18(2):553–569, 1989. CODEN CSSCDB. ISSN 0361-0918.

Szatrowski:1985:MDP

[Sza85] Ted H. Szatrowski. Missing data in the k-population multivariate normal patterned mean and covariance matrix testing and estimation problem. Communications in Statistics: Simulation and Computation, 14(2):357–370, 1985. CODEN CSSCDB. ISSN 0361-0918.

Taneja:1980:TSD

[TA80] Vidya S. Taneja and Leo A. Aroian. Time series in m dimensions: Autoregressive models. Communications in Statistics: Simulation and Computation, 9(5):491–513, 1980. CODEN CSSCDB. ISSN 0361-0918.

Tamarkin:1982:SSS

[Tam82] Maurry Tamarkin. A simulation study of the stochastic ridge k. Communications in Statistics: Simulation and Computation, 11 (2):159–173, 1982. CODEN CSSCDB. ISSN 0361-0918.

Tan:1980:CSE

[Tan80] W. Y. Tan. Comparative studies on the estimation of linkage by Bayesian method at maximum likelihood method. Communications in Statistics: Simulation and Computation, 9(1):19–41, 1980. CODEN CSSCDB. ISSN 0361-0918.

Tan:1986:IRC

[Tan86] W. Y. Tan. Inferences on regression coefficients in a regression model under heteroscedasticity and robustness with respect to departure from normality. Communications in Statistics: Simulation and Computation, 15(1):35–60, 1986. CODEN CSSCDB. ISSN 0361-0918.

Thornton:1985:NAE

[TB85] Daniel L. Thornton and Dallas S. Batten. A note on Almon’s endpoint constraints. Communications in Statistics: Simulation and Computation, 14(3):683–690, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tiku:1986:RTT

[TB86] M. L. Tiku and N. Balakrishnan. A robust test for testing the correlation coefficient. Communications in Statistics: Simulation and Computation, 15(4):945–971, 1986. CODEN CSSCDB. ISSN 0361-0918.

Tuku:1989:ERR

[TBA89] M. L. Tuku, N. Balakrishnan, and R. S. Ambagaspittya. Error rates of a robust classification procedure based on dichotomous and continuous random variables. Communications in Statistics: Simulation and Computation, 18(2):571–588, 1989. CODEN CSS-CDB. ISSN 0361-0918.

Taylor:1988:NCI

[TC88] Greg Taylor and W. J. Conover. A nonparametric confidence interval for slope based on Spearman’s rho. Communications in Statistics: Simulation and Computation, 17(3):905–916, 1988. CODEN CSSCDB. ISSN 0361-0918.

Tong:1989:SES

[TC89] Lee-Ing Tong and P. L. Cornelius. Studies on the estimation of the slope parameter in the simple linear regression model with one-fold nested error structure. Communications in Statistics: Simulation and Computation, 18(1):201–225, 1989. CODEN CSSCDB. ISSN 0361-0918.

Tarter:1987:GPD

[TF87] Michael E. Tarter and William R. Freeman. A graphical procedure for distinguishing between two data analysis pitfalls. Communications in Statistics: Simulation and Computation, 16(2):427–437, 1987. CODEN CSSCDB. ISSN 0361-0918.

Tarter:1986:FIU

[TFH86] Michael E. Tarter, William Freeman, and Alan Hopkins. A Fortran implementation of univariate Fourier series density estimation. Communications in Statistics: Simulation and Computation, 15(3):855–870, 1986. CODEN CSSCDB. ISSN 0361-0918.

Tan:1982:ANC

[TG82] W. Y. Tan and R. P. Gupta. On approximating the non-central Wishart distribution by central Wishart distribution a Monte Carlo study. Communications in Statistics: Simulation and Computation, 11(1):47–64, 1982. CODEN CSSCDB. ISSN 0361-0918.

Trader:1985:PPI

[TH85] Ramona L. Trader and H. Fenwick Huss. Prediction in the presence of imperfect inspections. Communications in Statistics: Simulation and Computation, 14(2):425–440, 1985. CODEN CSS-CDB. ISSN 0361-0918.

Turiel:1982:NSR

[THT82] Thomas P. Turiel, Gerald J. Hahn, and William T. Tucker. New simulation results for the calibration and inverse median estimation problems. Communications in Statistics: Simulation and Computation, 11(6):677–713, 1982. CODEN CSSCDB. ISSN 0361-0918.

Thursby:1985:NSM

[Thu85] Jerry G. Thursby. A note on Scheffé’s method of deriving confidence sets for directions and ratios of normals. Communications in Statistics: Simulation and Computation, 14(4):919–925, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tiku:1987:RPT

[Tik87] M. L. Tiku. A robust procedure for testing an assumed value of the population correlation coefficient. Communications in Statistics: Simulation and Computation, 16(4):907–924, 1987. CODEN CSSCDB. ISSN 0361-0918.

Tadikamalla:1982:TFF

[TJ82] Pandu R. Tadikamalla and Norman L. Johnson. Tables to facilitate fitting l_{u} distributions. Communications in Statistics: Simulation and Computation, 11(3):249–271, 1982. CODEN CSSCDB. ISSN 0361-0918.

Taylor:1985:IDM

[TJ85] John A. Taylor and Anthony J. Jakeman. Identification of a distributional model. Communications in Statistics: Simulation and Computation, 14(2):497–508, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tsai:1988:CTP

[TK88] Kao-Tai Tsai and James A. Koziol. A correlation type procedure for assessing multivariate normality. Communications in Statistics: Simulation and Computation, 17(2):637–651, 1988. CODEN CSSCDB. ISSN 0361-0918.

Tracy:1989:CSM

[TK89] Derrick S. Tracy and Khushnood A. Khan. Comparison of some MRPP and standard rank tests for two unequal samples. Communications in Statistics: Simulation and Computation, 18(2):729– 756, 1989. CODEN CSSCDB. ISSN 0361-0918.

Thode:1988:LSP

[TLF88] Henry C. Thode, Jr., Hung Kung Liu, and Stephen J. Finch. Large sample power of absolute moment tests. Communications in Statistics: Simulation and Computation, 17(4):1453–1458, 1988. CODEN CSSCDB. ISSN 0361-0918.

Tandon:1989:CNP

[TM89] P. K. Tandon and M. L. Moeschberger. Comparison of nonparametric and parametric methods in repeated measures designs — a simulation study. Communications in Statistics: Simulation and Computation, 18(2):777–792, 1989. CODEN CSSCDB. ISSN 0361-

Terza:1985:AEP

[TO85] Joseph V. Terza and A. Ason Okoruwa. An algorithm for the estimation of Poisson regressions involving structural change. Communications in Statistics: Simulation and Computation, 14(4): 853–866, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tse:1984:EAR

[Tse84] Y. K. Tse. Edgeworth approximations for t-ratios of 2SLS estimates of a dynamic model. Communications in Statistics: Simulation and Computation, 13(5):603–618, 1984. CODEN CSSCDB. ISSN 0361-0918.

Thode:1983:PTN

[TSF83] Henry C. Thode, Jr., Laurel A. Smith, and Stephen J. Finch. Power of tests of normality for detecting scale contaminated normal samples. Communications in Statistics: Simulation and Computation, 12(6):675–695, 1983. CODEN CSSCDB. ISSN 0361-0918.

Tabatabai:1985:SCS

[TT85a] M. A. Tabatabai and W. Y. Tan. Some comparative studies on testing parallelism of several straight lines under heteroscedastic variances. Communications in Statistics: Simulation and Computation, 14(4):837–844, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tan:1985:SRA

[TT85b] W. Y. Tan and M. A. Tabatabai. Some robust ANOVA procedures under heteroscedasticity and nonnormality. Communications in Statistics: Simulation and Computation, 14(4):1007–1026, 1985. CODEN CSSCDB. ISSN 0361-0918.

Tan:1986:RPC

[TT86] W. Y. Tan and M. A. Tabatabai. A robust procedure for comparing several means under heteroscedasticity and nonnormality. Communications in Statistics: Simulation and Computation, 15 (3):733–745, 1986. CODEN CSSCDB. ISSN 0361-0918.

Umphrey:1983:CMA

[Ump83] G. J. Umphrey. A comment on McKay’s approximation for the coefficient of variation. Communications in Statistics: Simulation and Computation, 12(5):629–635, 1983. CODEN CSSCDB. ISSN 0361-0918.

Uppuluri:1985:DFD

[UP85] V. R. R. Uppuluri and S. A. Patil. The distribution of the first j digits of beta related random variables. Communications in Statistics: Simulation and Computation, 14(2):467–472, 1985. CODEN CSSCDB. ISSN 0361-0918.

Ury:1980:FTS

[USM80] Hans K. Ury, Michael R. Stoline, and Brian T. Mitchell. Further tables of the Studentized maximum modulus distribution. Communications in Statistics: Simulation and Computation, 9(2):167– 178, 1980. CODEN CSSCDB. ISSN 0361-0918.

Vang:1987:AMN

[Van87] Yuchung J. Vang. An approximation of multivariate normal orthant probabilities of dimension 4 o: a contingency-table approach. Communications in Statistics: Simulation and Computation, 16(4):1039–1049, 1987. CODEN CSSCDB. ISSN 0361-0918.

Vuchkov:1989:GOD

[VDD89] I. N. Vuchkov, D. L. Damgaliev, and A. N. Donev. Generation of D-optimal designs in a finite design space. Communications in Statistics: Simulation and Computation, 18(1):319–337, 1989. CODEN CSSCDB. ISSN 0361-0918.

vanderLaan:1987:ETE

[vdLvP87] Paul van der Laan and Brand van Putten. Extensive tables of exact critical values of the nomparametric two-sample Halperin test for censored samples, and comparison with the normal approximation. Communications in Statistics: Simulation and Computation, 16(1):37–53, 1987. CODEN CSSCDB. ISSN 0361-0918.

Voss:1980:MAM

[VOA80] D. A. Voss, C. A. Oprian, and L. A. Aroian. Moving average models-time series in m-dimensions. Communications in Statistics: Simulation and Computation, 9(5):467–489, 1980. CODEN CSSCDB. ISSN 0361-0918.

Walker:1984:IAI

[Wal84] Joseph J. Walker. Improved approximations for the inverse moments of the positive binomial distribution. Communications in Statistics: Simulation and Computation, 13(4):515–519, 1984. CODEN CSSCDB. ISSN 0361-0918.

Wang:1982:CRV

[Wan82a] Paul C. C. Wang. On the computation of a robusl version of the optimal c(α) test. Communications in Statistics: Simulation and Computation, 11(3):273–284, 1982. CODEN CSSCDB. ISSN 0361-0918.

Wansbeek:1982:AAI

[Wan82b] Tom Wansbeek. Another approach to inverting a covariance matrix when data are unbalanced. Communications in Statistics: Simulation and Computation, 11(5):583–588, 1982. CODEN CSS-CDB. ISSN 0361-0918.

Wang:1988:OSC

[Wan88] C. Ming Wang. One-sided confidence intervals for the positive linear combination of two variances. Communications in Statistics: Simulation and Computation, 17(1):283–292, 1988. CODEN CSSCDB. ISSN 0361-0918.

Warren:1982:ACS

[War82] W. G. Warren. On the adequacy of the chi-squared approximation for the coefficient of variation. Communications in Statistics: Simulation and Computation, 11(6):659–666, 1982. CODEN CSS-CDB. ISSN 0361-0918.

Warren:1984:PST

[War84] W. G. Warren. The power of some tests of normality against Weibull alternatives. Communications in Statistics: Simulation and Computation, 13(2):243–255, 1984. CODEN CSSCDB. ISSN 0361-0918.

Whitten:1981:POC

[WC81] Betty Jones Whitten and A. Clifford Cohen. Percentiles and other characteristics of the four-parameter generalized gamma distribution. Communications in Statistics: Simulation and Computation, 10(2):175–218, 1981. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1986:NMC

[WCT86] Rand R. Wilcox, Ventura L. Charlin, and Karen L. Thompson. New Monte Carlo results on the robustness of the ANOVA F, W and F statistics. Communications in Statistics: Simulation and Computation, 15(4):933–943, 1986. CODEN CSSCDB. ISSN 0361-0918.

Wellek:1989:UAE

[Wel89a] Stefan Wellek. On the use of asymptotic expansion in computing the null distribution of Page’s L-statistic. Communications in Statistics: Simulation and Computation, 18(1):227–235, 1989. CODEN CSSCDB. ISSN 0361-0918.

Wells:1989:BLG

[Wel89b] Martin T. Wells. Basu’s lemma and goodness-of-fit tests. Communications in Statistics: Simulation and Computation, 18(2):711– 717, 1989. CODEN CSSCDB. ISSN 0361-0918.

Wesolowsky:1981:NDA

[Wes81] G. O. Wesolowsky. A new descent algorithm for the least absolute value regression problem. Communications in Statistics: Simulation and Computation, 10(5):479–491, 1981. CODEN CSSCDB. ISSN 0361-0918.

Wang:1982:ESE

[WF82] George H. K. Wang and Wayne A. Fuller. Estimators for a simultaneous equation model with lagged endogenous variables and autocorrelated errors. Communications in Statistics: Simulation and Computation, 11(2):123–142, 1982. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1982:CAX

[Wil82] Rand R. Wilcox. A comment on approximating the x^{2} distribution in the equiprobable case. Communications in Statistics: Simulation and Computation, 11(5):619–623, 1982. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1985:ESI

[Wil85a] Rand R. Wilcox. An extended and slightly improved table of critical values for testing q linear contrasts in a repeated measures design. Communications in Statistics: Simulation and Computation, 14(1):55–69, 1985. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1985:PPP

[Wil85b] Rand R. Wilcox. Percentage points of the product of two correlated t variates. Communications in Statistics: Simulation and Computation, 14(1):143–157, 1985. CODEN CSSCDB. ISSN 0361-

Wilcox:1986:CVC

[Wil86a] Rand R. Wilcox. Critical values for the correlated t-test when there are missing observations. Communications in Statistics: Simulation and Computation, 15(3):709–714, 1986. CODEN CSS-CDB. ISSN 0361-0918.

Wilcox:1986:ISC

[Wil86b] Rand R. Wilcox. Improved simultaneous confidence intervals for linear contrasts and regression parameters. Communications in Statistics: Simulation and Computation, 15(4):917–932, 1986. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:1989:PPW

[Wil89] Rand R. Wilcox. Percentage points of a weighted Kolmogorov– Smirnov statistic. Communications in Statistics: Simulation and Computation, 18(1):237–244, 1989. CODEN CSSCDB. ISSN 0361-0918.

Winterbottom:1980:EBN

[Win80] Alan Winterbottom. Estimation for the bivariate normal correlation coefficient using asymptotic expansions. Communications in Statistics: Simulation and Computation, 9(6):599–609, 1980. CODEN CSSCDB. ISSN 0361-0918.

Wingo:1983:ELC

[Win83] Dallas R. Wingo. Estimating the location of the Cauchy distribution by numerical global optimization. Communications in Statistics: Simulation and Computation, 12(2):201–212, 1983. CODEN CSSCDB. ISSN 0361-0918.

Withers:1989:DCS

[Wit89] C. S. Withers. The distribution and cumulants of a Studentised statistic. Communications in Statistics: Simulation and Computation, 18(1):295–318, 1989. CODEN CSSCDB. ISSN 0361-0918.

Wellington:1981:VSM

[WN81] John F. Wellington and Subhash C. Narula. Variable selection in multiple linear regression using the minimum sum of weighted absolute errors criterion. Communications in Statistics: Simulation and Computation, 10(6):641–648, 1981. CODEN CSSCDB. ISSN 0361-0918.

Wellington:1984:ARQ

[WN84] John F. Wellington and Subhash C. Narula. An algorithm for regression quantiles. Communications in Statistics: Simulation and Computation, 13(5):683–704, 1984. CODEN CSSCDB. ISSN 0361-0918.

Wood:1987:SSD

[Woo87] Andrew T. A. Wood. The simulation of spherical distributions in the Fisher-Bingham family. Communications in Statistics: Simulation and Computation, 16(3):885–898, 1987. CODEN CSSCDB. ISSN 0361-0918.

Wood:1989:FAD

[Woo89] Andrew T. A. Wood. An f approximation to the distribution of a linear combination of chi-squared variables. Communications in Statistics: Simulation and Computation, 18(4):1439–1456, 1989. CODEN CSSCDB. ISSN 0361-0918.

Whaley:1985:OPT

[WQ85] Fredrick S. Whaley and Dana Quade. Optimizing the power of the two-sample multidimensional runs statistic: guidelines based on computer simulation. Communications in Statistics: Simulation and Computation, 14(1):1–11, 1985. CODEN CSSCDB. ISSN 0361-0918.

Wang:1987:MCS

[WR87] Chinying J. Wang and Jagdish S. Rustagi. A Monte Carlo study of the effect of grouping on Cox’s regression model. Communications in Statistics: Simulation and Computation, 16(2):321–331, 1987. CODEN CSSCDB. ISSN 0361-0918.

Walter:1988:ETM

[WS88] S. D. Walter and L. W. Stitt. Extended tables for moments of gamma distribution order statistics. Communications in Statistics: Simulation and Computation, 17(2):471–487, 1988. CODEN CSSCDB. ISSN 0361-0918.

Woodruff:1986:MGF

[WYMD86] Brian W. Woodruff, John D. Yoder, Albert H. Moore, and Edward J. Dunne. Modified goodness-of-fit tests for the logistic distribution with unknown location and scale parameters. Communications in Statistics: Simulation and Computation, 15(1):77–83, 1986. CODEN CSSCDB. ISSN 0361-0918.

Xie:1989:SRT

[Xie89] Min Xie. On the solution of renewal-type integral equations. Communications in Statistics: Simulation and Computation, 18(1): 281–293, 1989. CODEN CSSCDB. ISSN 0361-0918.

Yandell:1988:AMS

[Yan88] Brian S. Yandell. Algorithms for multidimensional semiparametric GLM’s. Communications in Statistics: Simulation and Computation, 17(1):295–312, 1988. CODEN CSSCDB. ISSN 0361-0918.

Young:1987:TER

[YB87] David H. Young and Saad T. Bakir. Testing exponential regression against a gamma alternative. Communications in Statistics: Simulation and Computation, 16(3):719–734, 1987. CODEN CSS-CDB. ISSN 0361-0918.

Yohai:1987:TBW

[YF87] Víctor J. Yohai and Nélida E. Ferretti. Tests based on weighted rankings in complete blocks: exact distribution and Monte Carlo simulation. Communications in Statistics: Simulation and Computation, 16(2):333–347, 1987. CODEN CSSCDB. ISSN 0361-0918.

Yen:1988:MGF

[YM88] Vincent C. Yen and Albert H. Moore. Modified goodness-of-fit test for the Laplace distribution. Communications in Statistics: Simulation and Computation, 17(1):275–281, 1988. CODEN CSSCDB. ISSN 0361-0918.

Zacks:1980:NDD

[Zac80] S. Zacks. Numerical determination of the distributions of stopping variables associated with sequential procedures for detecting epochs of shift in distributions of discrete random variables. Communications in Statistics: Simulation and Computation, 9(1): 1–18, 1980. CODEN CSSCDB. ISSN 0361-0918.

Zaher:1984:CSP

[ZH84] Adel M. Zaher and Robert L. Heiny. A comparison of selection procedures for selecting the best normal population under heterogeneity of variance. Communications in Statistics: Simulation and Computation, 13(5):635–654, 1984. CODEN CSSCDB. ISSN 0361-0918.